### v2

# Spectraformer: A Unified Random Feature Framework for Transformer

Abstract
Linearization of attention using various kernel approximation and kernel learning techniques has shown promise. Past methods use a subset of combinations of component functions and weight matrices within the random features paradigm. We identify the need for a systematic comparison of different combinations of weight matrix and component functions for attention learning in Transformer. In this work, we introduce Spectraformer, a unified framework for approximating and learning the kernel function in linearized attention of the Transformer. We experiment with broad classes of component functions and weight matrices for three textual tasks in the LRA benchmark. Our experimentation with multiple combinations of component functions and weight matrices leads us to a novel combination with 23.4% faster training time and 25.2% lower memory consumption over the previous SOTA random feature Transformer, while maintaining the performance, as compared to the Original Transformer. Our code is available at: https://github.com/dukeraphaelng/spectraformer.
1 Introduction
Transformer [1] has revolutionized the landscape of natural language processing (NLP) and forms the basis of almost all state-of-the-art language models. Its influence has reached beyond NLP into computer vision [2], speech processing [3] and other fields. Compared to its precursors like the LSTMs, Transformer leverages parallelism since it is fully based on the attention mechanism. Therefore, the performance of Transformer is tied to the effective use of attention. Attention in the Original Transformer (referred to as ‘OT’ hereafter) uses the softmax function which is quadratic in time complexity. However, softmax is not fundamental to attention. The softmax serves the role of a compatibility function, which captures the degree of compatibility between two tokens (see Equations 1 and 2). Extensive experiments with other compatibility functions show that the softmax is not the only possibility, and that the ‘best’ compatibility function depends on the task [4].
Replacing the softmax formulation with alternatives may allow to reduce the time complexity of attention computation. In this paper, we focus on the subset of alternatives that employs kernel functions. Kernel functions have been historically used in machine learning algorithms to simplify the estimation of parameters [5]. [4] show that attention can be formulated using kernels and then be linearized [6], given the feature map of the respective attention kernel. Random feature-based algorithms [7] are a family of algorithms, inspired by spectral analysis, which provides the feature map associated with a kernel. A random feature consists of three components: component function, weight matrix, and component weight (see Section 3). Due to their robustness and flexibility, random features give rise to ‘random feature Transformers’, the family of Transformers with linearized attention via random features (hereby abbreviated as RF), first with the work by [8]. This is seen to lead to three strands of research: (A) Optimization of weight matrices either by making the matrices have lower time or space complexity, or by incorporating beneficial properties like orthogonality to reduce variance [9, 8, 10, 11]; (B) Enhancement of component functions by either: (i) finding a function with a tighter variance or approximation bound, or (ii) be engineered to output with numerical stability and be bounded [8, 12, 13]; (C) Parameterization of weight matrices to perform kernel learning instead of kernel approximation [14]. (A) and (B) seek to improve approximation quality via approximation error and variance reduction. (C) dispenses with the approximation of the attention kernel (typically the softmax), and makes the kernel itself learnable, to attain a better performance. However, the three have been explored separately, creating a situation where there are different ways to construct the random features. Additionally, past works only explore certain combinations of component functions and weight matrices, creating disjoint overlaps and gaps in the literature. Finally, the weight matrices are often those explicitly designed in the context of kernelized attention in the Transformer [9, 8, 10, 11], ignoring those which are popular in the kernel methods literature (e.g., [7]). These limitations necessitate a unified framework that allows the comparison of different combinations in search of the best-performing one.
Figure 1: Peak memory consumption (x-axis), average accuracy (y-axis) on three textual LRA [15] tasks: ListOps, Text, Retrieval. Color indicates training time, the darker the faster, the brighter the slower (see legend). Models compared are the best Spectraformer variant (PosRF-MM) (boldened), the OT, and other efficient Transformers (RF and non-RF) in the literature. Values for non-RF Transformers are taken from [16].
We introduce Spectraformer, a unified framework of random feature-based attention in Transformer. It utilizes spectral analysis-inspired random features to construct linearized attention in Transformers. Spectraformer allows for the experimentation of various combinations of weight matrices and component functions. We will open-source our codebase upon acceptance. Through our benchmarking experiments on three textual Long Range Arena [15] tasks and 18 combinations of weight matrices and component functions, Spectraformer is able to find novel combinations that outperform existing random feature-based Transformers in terms of performance, training time, and peak memory consumption. Figure 1 compares our best-known Spectraformer combination (PosRF-MM) against the previous SOTA models in random feature Transformer, efficient Transformers, and the OT. The x and y axes denote peak memory and average accuracy, while the colour of the circle indicates the memory consumption. Spectraformer (specifically, PosRF-MM) is more accurate than the previous SOTA models, SADERF-ORF (a.k.a., FAVOR# [13]), OPRF-ORF (a.k.a., FAVOR++ [12]), PosRF-ORF (a.k.a., FAVOR+ [8]) and maintains a comparable accuracy against the OT, whilst significantly reducing training time and memory consumption over both. However, it underperforms in accuracy and training time against Skyformer [16] and Nyströmformer [17]. We hope that Spectraformer will enable the mechanism to benchmark more random feature-based Transformers for these and other tasks. Our contributions are:
  1. 1.
    Our novel framework, Spectraformer, which generalizes over all past works in linearized attention with kernel approximation and kernel learning.
  2. 2.
    We demonstrate how Spectraformer can discover novel SOTA models in the random feature Transformer family for the LRA benchmark. Our best model has comparable accuracy to the OT whilst significantly reducing training time and peak memory consumption.
  3. 3.
    This is the first work to compare random feature Transformers with other efficient Transformers.
2 Preliminaries
2.1 Kernelized Attention
OT attention [1] can be defined using the unified attention model [18] as follows. Given a collection of inputs and a learning objective, attention mechanism learns the attention matrix A which captures the associative information between pairs of tokens in a sequence of length $N$, with each row $\textbf{A}_{i}$ being the representation of a token $i$ with every other token $j\in[1,N]$. This is done by creating three learnable representations of each token $i$ called query ($\boldsymbol{q}_{i}$), key ($\boldsymbol{k}_{i}$), and value ($\boldsymbol{v}_{i}$). The attention representation $\textbf{A}_{i}$ is calculated as a weighted sum of value $\boldsymbol{v}_{j}$ given attention weights $a_{j}$. The general equation for $\textbf{A}_{i}$ is given as:
$\begin{array}[]{cccc}\boldsymbol{A}_{i}=\sum_{j=1}^{N}\frac{f(\boldsymbol{q}_{i},\boldsymbol{k}_{j})}{\sum_{l}f(\boldsymbol{q}_{i},\boldsymbol{k}_{l})}\boldsymbol{v}_{j}=\sum_{j=1}^{N}a_{j}\boldsymbol{v}_{j};&a=g(e);&e=f(\boldsymbol{q}_{i},\boldsymbol{k}_{j})\end{array}$ (1)
Attention weights are generated using a distribution function $g$ applied on an energy score $e$. The energy score is captured by applying a compatibility function on a query $\boldsymbol{q}_{i}$ and a set of keys K, in order to compute a score of a specific relationship between the pairs. The attention in the OT is called the ‘scaled dot-product’ which defines the compatibility function $f$ and the distribution function $g$ as follows:
$\begin{array}[]{cc}f(\boldsymbol{q}_{i},\textbf{K})=\frac{\boldsymbol{q}_{i}^{T}\textbf{K}}{\sqrt{d_{k}}}&g(\boldsymbol{x}_{j})=\frac{\exp(\boldsymbol{x}_{j})}{\sum_{l}\exp(\boldsymbol{x}_{l})}\end{array}$ (2)
We can then have the attention matrix (with $\exp(.)$ being applied element-wise) as:
$\begin{array}[]{cc}\boldsymbol{A}_{i}=\sum_{j=1}^{N}\frac{\exp(\boldsymbol{q}_{i}\boldsymbol{k}_{j}^{T}/\sqrt{d_{k}})}{\sum_{l}\exp(\boldsymbol{q}_{i}\boldsymbol{k}_{l}^{T}/\sqrt{d_{k}})}\boldsymbol{v}_{j}&A=softmax(\frac{\boldsymbol{QK}^{T}}{\sqrt{d_{k}}})\boldsymbol{V}\end{array}$ (3)
with $\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V}$ being the matrix of queries, keys, and values respectively with each row being $\boldsymbol{q}_{i},\boldsymbol{k}_{i},\boldsymbol{v}_{i}$.
We now discuss the intrinsic connection between kernel and attention. With the compatibility function $f$ as a kernel $f(\boldsymbol{q}_{i},\boldsymbol{k}_{j})=\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{j})$, Equation 1 can be rewritten as:
$\boldsymbol{A}_{i}=\sum_{j=1}^{N}\frac{\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{j})}{\sum_{l}\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{l})}\boldsymbol{v}_{j}$ (4)
As shown in [4], this is indeed the Nadaraya-Watson Kernel Estimator, with $\mathbb{E}_{p(\boldsymbol{k}_{j}|\boldsymbol{q}_{i})}[\boldsymbol{v}_{j}|X=\boldsymbol{k}_{j}]=\boldsymbol{A}_{i}$, $l_{i}\left(\boldsymbol{k}_{j})=p(\boldsymbol{k}_{j}|\boldsymbol{q}_{i}\right)$, $Y_{i}=\boldsymbol{v}_{j}$:
$\begin{array}[]{ccc}\mathbb{E}_{l_{i}}[Y_{i}|X=\boldsymbol{u}]=\sum_{i=1}^{N}l_{i}(\boldsymbol{u})Y_{i};&l_{i}(\boldsymbol{u})=\frac{\mathcal{K}(\frac{\boldsymbol{u}-\boldsymbol{x}_{i}}{h})}{\sum_{j=1}^{n}\mathcal{K}(\frac{\boldsymbol{u}-\boldsymbol{x}_{j}}{h})}\end{array}$ (5)
Specifically, the scaled dot-product attention in the OT in Equation 3 corresponds to the softmax kernel (see Equation 6).
$\mathcal{K}_{softmax}(\boldsymbol{q}_{i},\boldsymbol{k}_{j})=\exp(\boldsymbol{q}_{i}^{T}\boldsymbol{k}_{j}/\sqrt{d_{k}})$ (6)
2.2 Linearizing Kernelized Attention via Random Features
Since a kernel is the inner product of two vectors in some space with respect to some feature map $\phi$, i.e., $\mathcal{K}(\boldsymbol{x},\boldsymbol{y})=\phi(\boldsymbol{x})\phi(\boldsymbol{y})^{T}$, then we can rewrite Equation 4 as:
$\begin{array}[]{c}\boldsymbol{A}_{i}=\frac{\sum_{j}\phi(\boldsymbol{q}_{i})\phi(\boldsymbol{k}_{j})^{T}\boldsymbol{v}_{j}}{\sum_{l}\phi(\boldsymbol{q}_{i})\phi(\boldsymbol{k}_{l})^{T}}=\frac{\phi(\boldsymbol{q}_{i})\sum_{j}\phi(\boldsymbol{k}_{j})\otimes\boldsymbol{v}_{j}}{\phi(\boldsymbol{q}_{i})\sum_{l}\phi(\boldsymbol{k}_{l})^{T}}\end{array}$ (7)
If we pre-compute $\sum_{j}\phi(\boldsymbol{k}_{j})\otimes\boldsymbol{v}_{j}$ and $\sum_{l}\phi(\boldsymbol{k}_{l})^{T}$, the entire term becomes $O(1)$ and we only need to compute Equation 7 $N$ times for each $i$, resulting in an attention matrix calculated in linear time $O(N)$. We now refer to Equation 7 as linearized attention [6].
All that is left is to retrieve the feature map $\phi$, the process of which is called kernel approximation (see Section 2.3), which corresponds to the kernels we specify. There are many popular kernels from which we can choose, but typically, we want to approximate the softmax kernel in Equation 6. However, since the softmax relates to the $\mathcal{K}_{RBF}(\boldsymbol{x},\boldsymbol{y})=\exp(-\frac{||\boldsymbol{x}-\boldsymbol{y}||^{2}}{2})$ via the following equation:
$\mathcal{K}_{softmax}(\boldsymbol{x},\boldsymbol{y})=\exp(\frac{||\boldsymbol{x}||^{2}}{2})\mathcal{K}_{RBF}(\boldsymbol{x},\boldsymbol{y})\exp(\frac{||\boldsymbol{y}||^{2}}{2})$ (8)
Approximating the RBF is easier than the softmax directly in the method we introduce in Section 2.3, hence we approximate the softmax indirectly via the RBF in Equation 8.
Oftentimes, however, we might not want to specify a kernel, since kernel choice is indeed a hyperparameter and the softmax is by no means essential or necessary as shown in [4]. Then kernel learning would be a more suitable option, the detail of which is discussed in Section 2.4. Notwithstanding, we first introduce the main kernel approximation technique: random features.
2.3 Random Features for Kernel Approximation
Of the kernel approximation family of methods, the most successful and appropriate for linearized attention in Transformer has been the random features approach [19]. Estimating kernels using random features relies on a fundamental insight from harmonic analysis called Bochner’s theorem [20] stated as follows:
A continuous shift-invariant kernel $\mathcal{K}(\boldsymbol{x},\boldsymbol{y})=\mathcal{K}(\boldsymbol{x}-\boldsymbol{y})$ on $\mathbb{R}^{d}$ is positive definite if and only if $\mathcal{K}(\boldsymbol{\delta})$ is the Fourier transform of a non-negative measure $p(\boldsymbol{\omega})$. If $\mathcal{K}(\boldsymbol{\delta})$ is properly scaled, that is $\mathcal{K}(0)=1$ then the measure $p(\boldsymbol{\omega})$ is a proper probability distribution.
inline]define unambiguously the dimension of x and phi here, also the dimension before and after approximation of the features map
$\mathcal{K}(\boldsymbol{\delta})=\int p(\boldsymbol{\omega})\exp(i\boldsymbol{\omega\delta})\mathrm{d}\boldsymbol{\omega}$ (9)
When approximating the integration in this equation via Monte Carlo sampling with $s$ samples, $\boldsymbol{\omega}\sim p(\boldsymbol{\omega})$, we can obtain the feature map:
$\begin{array}[]{ll}\mathcal{K}(\boldsymbol{\delta})&=\mathbb{E}[\phi_{\boldsymbol{\omega}}(\boldsymbol{x})^{T}\phi_{\boldsymbol{\omega}}(\boldsymbol{y})]\approx<\phi_{\boldsymbol{\omega}}(\boldsymbol{x}),\phi_{\boldsymbol{\omega}}(\boldsymbol{y})>\\ \phi_{\boldsymbol{\omega}}(\boldsymbol{x})\!&=\!\frac{1}{\sqrt{s}}\big{[}\exp(-\mbox{i}{\boldsymbol{\omega}}^{\!\top}_{1}\boldsymbol{x}),\ldots,\exp(-\mbox{i}{\boldsymbol{\omega}}^{\!\top}_{s}\boldsymbol{x})]^{\!\top}\end{array}$ (10)
2.4 Random Features for Kernel Learning
Previous random feature techniques are robust in approximating kernels for specific learning problems. However, kernels may be chosen using heuristics and by convention from a popular subset. Kernel is very much a hyperparameter. This could pose a problem since there is no free lunch in learning as much as in kernel choice [21]. [4] has shown experimental results that this is also true in the context of attention in Transformer. Hence, instead of picking a kernel for the attention, we could learn the kernel via kernel learning an established kernel methods technique. Kernel learning has shown to be effective in general [22, 23, 24] and in the case of attention via Gaussian Mixture Model (GMM), learnable weight matrices (FastFood), deep generative model (DGM) [14]. Kernel learning is achieved by making the weight matrix $\boldsymbol{W}$ learnable where $\boldsymbol{\omega}_{i}$ is the output of a function parameterizing $p(\boldsymbol{\omega})$.
The kernelized attention via random features discussed here is a popular approach, however it is by no means the only one. A more detailed discussion is offered in related work in Section A.2. Specifically, Section A.2.1 provides alternative kernelized attention formulations and Section A.2.2 provides alternative kernel approximations to random features.
3 Spectraformer
Spectraformer is a unifed random feature framework that allows for the combination of any weight matrix with any component function to be used as kernelized attention in Transformer. The general equation for random feature is shown in Equation 11, generalized from Equation 10, based on Liu et al. [7], Choromanski et al. [8].
$\begin{array}[]{rl}\phi_{\boldsymbol{\omega}}(\boldsymbol{x})=&\frac{1}{\sqrt{s}}\big{[}a_{1}f_{1}(\boldsymbol{\omega}_{1},\boldsymbol{x}),...,a_{s}f_{1}(\boldsymbol{\omega}_{s},\boldsymbol{x}),...,\\ &a_{1}f_{l}(\boldsymbol{\omega}_{1},\boldsymbol{x}),...,a_{s}f_{l}(\boldsymbol{\omega}_{s},\boldsymbol{x})]^{\!\top}\\ \boldsymbol{W}=&[\boldsymbol{\omega}_{1},...,\boldsymbol{\omega}_{s}]^{T}\in\mathbb{R}^{s*d}\end{array}$ (11)
Spectraformer is shown in Figure 2. The attention of the OT (on the left-hand side) is replaced by the attention of the Spectraformer which pre-computes the Hadamard (element-wise) product of values $\boldsymbol{v}$ and transformed keys $\phi(\boldsymbol{k})$ scaled by $\phi(\boldsymbol{k})$, the product of which is multiplied with the transformed queries $\phi(\boldsymbol{q})$. The linear map $\phi$ is composed from weight matrices and component functions. The linear map has $s$ components with $a=1$ (see Equation 11).
The figure and the equation together highlight the three strands of research as mentioned previously in Section 1: (A) Weight matrix approximator: constructing $\boldsymbol{W}$ more effectively; (B) Weight matrix learner: parameterizing $\boldsymbol{W}$ instead of sampling; (C) Component function: constructing $f_{j}$ more effectively.
3.1 Weight Matrices
In Spectraformer, the weight matrix $\boldsymbol{\boldsymbol{W}}$ which is either an approximator (from the families of (e.g., ORF [9], SORF [9]), Unit-cube (e.g., QMC [25], MM [26]), Quadrature (e.g., SGQ [27]). or a learner (FastFoodL [14]).
Figure 2: Spectraformer framework based on Equation 11
We term the matrix $\boldsymbol{W}$ as weight matrix instead of random matrix since although $\boldsymbol{W}$ acts like a random matrix ($\boldsymbol{\omega}_{i}\sim p(.)$, see Equation 10), it is not guaranteed to be one. In the case of weight matrix approximator, they approximate a random matrix and in the case of weight matrix learner, they parameterize a weight matrix, thereby implicitly learning a distribution $p(.)$ associated with such $\boldsymbol{W}$. The weight matrix is given in Definition 3.1.
Definition 3.1.
Provided that a matrix can substitute for a random matrix in Equation 11, and the solution for this equation given such matrix and $f$ as TrigRF being the solution for Equation 9, i.e., as an unbiased estimator of $\mathcal{K}$, and TrigRF being $f_{1}=f_{l}=\exp(-i\boldsymbol{\omega}^{T}\boldsymbol{x})$ (see Equation 10) then such a matrix is a weight matrix.
It follows from Definition 3.1 that for any weight matrix $\boldsymbol{W}$, $f=TrigRF$ is a valid component function. Spectraformer enables weight matrix approximation in terms of three families based on how they solve the intractable integral in Equation 9:
  • Monte Carlo sampling-based weight matrix involves approximating the integral of Equation 9 using Monte Carlo sampling with the solution given by Equation 10 [19] with $\boldsymbol{W}$ being the ‘Base’ random matrix of $p(.)$, $\boldsymbol{\omega}_{i}\sim p(.)$. The random matrix $\boldsymbol{W}$ can be constructed more efficiently (either reducing time or space complexity) with FastFood [28]. The random matrix $\boldsymbol{W}$ can also be enforced to form geometrical couplings such as orthogonality to reduce approximation error with ORF, SORF [9]. Both these methods produce valid random matrices.
  • Unit-cube sampling-based weight matrix transforms Equation 9 to an integral on the unit cube $[0,1]^{d}$, then performs the approximation with uniform and independent points, thus reducing variance. The first approach is QMC [25], where $\boldsymbol{W}:\boldsymbol{\omega}_{i}=\Phi^{-1}(\boldsymbol{t}_{i})$, $\Phi$ being the cumulative distribution function (CDF) associated with $p(.)$ and $\boldsymbol{t}_{i}$ being a low discrepancy sequence. $\Phi^{-1}(\boldsymbol{t}_{i})\sim p(.)$ whilst having a lower variance than direct sampling from $p(.)$. MM [26, 7] improves over QMC by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$ on $\Phi^{-1}$.
  • Quadrature-based weight matrix uses quadrature rules with non-uniform deterministic weights. The main approach explored is SGQ [27] which uses one-dimensional Gaussian quadrature rule by assuming $\mathcal{K}$ factorizes with respect to the dimensions. Smolyak rule is further implemented to alleviate the curse of dimensionality.
[7] show that in various kernel benchmark datasets (among combination with TrigRF), QMC and ORF have consistently high performance. Due to the significant number of weight matrices in random features literature, we decide to cover only the most prominent ones. Appendix A.3 provides the technical details of the approximators discussed here and those that we left out.
Spectraformer also allows to learn weight matrices by learning $p(.)$ via parameterizing $\boldsymbol{W}$. [29] introduce several of these parametrization schemes, which are then adapted into the attention setting by [14]. The best performing method is via making the constituent matrices of FastFood learnable, which we term FastFoodL. In general, a weight matrix learner is valid as long as it can implicitly model any distribution. In addition the TrigRF, PosRF has also been applied on FastFoodL. Technical details and other learners we have not covered are explored in Appendix A.4. We now explore existing works which have developed from Rahimi and Recht [19] (see Equation 9, 10) and how they lead to the development of Spectraformer.
3.2 Component functions
Spectraformer incorporates component functions $f$ such as PosRF, OPRF, SADERF. Component functions in Spectraformer combine each weight matrix row $\boldsymbol{\omega}_{i}$ with the input $\boldsymbol{x}$. The base case of the component function is TrigRF. Component function is given in Definition 3.2
Definition 3.2.
A function of $\boldsymbol{x}$ with respect to $\boldsymbol{\omega}$ is a valid component function if and only if $\mathcal{K}\approx\phi_{\boldsymbol{\omega}}(\boldsymbol{x})^{T}\phi_{\boldsymbol{\omega}}(\boldsymbol{y})$ given that $\boldsymbol{\omega}_{i}\sim p(.)$ in Equation 10.
It follows that all component functions have provided theoretical guarantees given that $\boldsymbol{W}$ is a random matrix. In addition, PosRF, OPRF, and SADERF have also been proven theoretically to approximate $\mathcal{K}_{RBF}$ given $\boldsymbol{W}$ being the ORF.
While TrigRF[19] is the base component function and guarantees $\mathcal{K}(\boldsymbol{\delta})$ to be close to $\phi_{TrigRF}(\boldsymbol{x})\cdot\phi_{TrigRF}(\boldsymbol{y})$ due to uniform convergence, it is typically evaluated using only the real component of $\exp(-i\boldsymbol{\omega}_{j}^{T}\boldsymbol{x})$. This leads to $\cos$ and $\sin$ functions which can produce unstable behavior, especially when $K(\boldsymbol{\delta})$ is close to 0, where variance is infinite [8, 12]. Therefore, TrigRF is avoided in Spectraformer.
  • PosRF [8] is introduced in Spectraformer as a simple modification over TrigRF to counter this instability. PosRF guarantees positive function output and its variance approaching 0 as $K(\boldsymbol{\delta})$ approaches 0. [12] generalize TrigRF and PosRF to GERF.
  • OPRF [12] is one of the solution to the reduction of the variance equation of GERF. OPRF is both positive and bounded (whereas PosRF is positive and unbounded). [13] generalize GERF to dense matrix instead of scalar parameters leading to DERF.
  • SADERF [13] is one of the solutions to the reduction of the variance equation of DERF. Whilst DERF has many solutions, they all rely on operations which are incompatible with many deep learning libraries on GPU and TPU like SVD and eigen decompositions. SADERF is an extension of GERF with a tighter variance bound on DERF.
The alternatives for component functions allow Spectraformer to approximate the RBF with $p(.)$ corresponding to the Gaussian. [13] and [12] show that SADERF-ORF is the SOTA combination followed by OPRF-ORF and PosRF-ORF. Technical details and other component function(s) not covered are explored in Appendix A.5.
Table 1: Summary of research gap in kernelized attention. Rows are the component functions, columns are the weight matrices. For each cell, the left item is the cited work for this combination for typical kernel tasks, while the right item is the cited work for Transformer-based tasks. $\times$ indicates no prior work for the combination. Spectraformer explores all combinations listed here, excluding ‘Base’ due to exhaustive work as seen in the table.
f Monte Carlo Unit-cube Quadrature Learner
Base ORF SORF QMC MM SGQ FastFoodL
TrigRF [19], [8] [9], [8] [9], $\times$ [25], $\times$ [26], $\times$ [27], $\times$ [29], [14]
PosRF [8], [8] [8], [8] $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, [14]
OPRF [12], [12] [12], [12] $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$
SADERF [13], [13] [13], [13] $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$
3.3 Utility
There has been an unsystematic attempt at combining component functions and weight matrices. On one hand, most component functions have had theoretical and experimental results in combining with Base and ORF. On the other hand, most weight matrices have had results in combining with TrigRF or PosRF. This means that a lot of potential combinations have not been studied previously. Table 1 shows the wide research gap across 14 combinations which have never been studied in either the kernel or Transformer setting. Spectraformer highlights this research gap. Excluding TrigRF due to its instability and Base due to its underperformance, we conduct experiments on 18 combinations.
Spectraformer is a generic random feature framework that allows for the combination of any weight matrix with any component function provided that the weight matrix satisfies Definition 3.1 and the component function satisfies Definition 3.2. We note that Spectraformer is an experimental rather than a theoretical framework. Due to the large number of weight matrices and component functions, it is infeasible for the time being to derive a general mathematical proof for the bounds and convergence of Spectraformer. Therefore, we seek to validate its feasibility and show its ability to discover novel and performant combination via experimental results.
4 Results
4.1 Experiment Details
The LRA covers tasks of different sequence lengths, difficulty, and objective and it is designed to evaluate efficient Transformers, of which Spectraformer belongs to. Specifically, we evaluate the models on three LRA tasks: ListOps [30], Byte-Level Text Classification (using the IMDb review dataset) [31], and Byte-Level Document Retrieval (using the ACL Anthology Network dataset) [32]. All our models are experimented on Google Cloud Platform with n1-standard-4 (4 vCPUs, 15 GB RAM) and 1 NVIDIA T4 GPU. Our codebase is based on and includes the adapted or original implementation from [16], [14], [7], [8], [12], [13], [33] (Apache License, Version 2.0), [34] (Apache License, Version 2.0), [35], [36] (Apache License Version 2.0), [17].
Due to the significant number of models that are trained, we do not run the experiments on different seeds nor perform hyper-parameter tuning. The parameters are identical to [16] and chosen to limit the parameter count to account for the marked training time of multiple models. All the parameters for Spectraformer variants are kept identical for fair comparison, with the only difference being their random feature component. Our code is implemented in Python 3.12 and Pytorch, we use the Transformer base from [16]. Our hyperparameters are available in Table 5. The model name convention is [component function]-[weight matrix] (e.g., OPRF-FastFoodL).
Table 2: Experimental results of Spectraformer variants on the LRA benchmark. We report accuracy (on the test set), time (training time (hour)), and memory (peak memory consumption (GB)). $L$: ListOps, $T$: Text, $R$: Retrieval, and $\mu$ for the average of the three. The best variant (PosRF-MM) with highest average accuracy, and overall low training time/memory is higlighted in red. Existing RF techniques are cited in the model column. Novel combinations are marked by $\checkmark$ on the Novel column.
Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
Novel Model L T R $\mu$ L T R $\mu$ L T R $\mu$
PosRF-ORF [8] 37.35 61.60 80.53 59.83 1.30 2.84 2.89 2.34 1.17 2.31 2.10 1.86
$\checkmark$ PosRF-SORF 22.98 63.31 65.52 50.60 1.29 2.81 2.83 2.31 1.17 2.31 2.10 1.86
$\checkmark$ PosRF-QMC 37.50 60.41 80.71 59.54 1.30 2.84 2.89 2.34 1.17 2.31 2.10 1.86
$\checkmark$ PosRF-SGQ 37.45 62.68 78.37 59.50 1.30 2.83 2.89 2.34 1.17 2.31 2.10 1.86
$\checkmark$ PosRF-MM 38.05 61.85 80.67 60.19 1.31 2.84 2.89 2.35 1.17 2.31 2.10 1.86
PosRF-FastFoodL [14] 24.95 64.76 76.37 55.36 2.69 5.59 5.61 4.63 0.78 1.56 1.53 1.29
OPRF-ORF [12] 37.50 59.35 80.90 59.25 1.61 3.45 3.50 2.86 1.36 2.71 2.56 2.21
$\checkmark$ OPRF-SORF 32.11 64.34 77.47 57.97 1.57 3.37 3.41 2.78 1.36 2.71 2.56 2.21
$\checkmark$ OPRF-QMC 38.41 60.32 80.80 59.84 1.61 3.46 3.51 2.86 1.36 2.71 2.56 2.21
$\checkmark$ OPRF-MM 38.71 60.39 80.45 59.85 1.61 3.46 3.51 2.86 1.36 2.71 2.56 2.21
$\checkmark$ OPRF-SGQ 22.53 61.34 79.29 54.39 1.60 3.45 3.47 2.84 1.36 2.71 2.56 2.21
$\checkmark$ OPRF-FastFoodL 37.40 64.04 78.32 59.92 2.81 5.84 5.86 4.84 0.85 1.69 1.67 1.40
SADERF-ORF [13] 37.35 61.37 80.87 59.86 1.61 3.52 3.58 2.90 1.44 2.86 2.69 2.33
$\checkmark$ SADERF-SORF 32.31 64.49 76.43 57.74 1.58 3.44 3.49 2.84 1.44 2.86 2.69 2.33
$\checkmark$ SADERF-QMC 37.70 59.89 80.65 59.41 1.61 3.52 3.58 2.90 1.44 2.86 2.69 2.33
$\checkmark$ SADERF-MM 38.00 60.80 80.48 59.76 1.61 3.52 3.58 2.90 1.44 2.86 2.69 2.33
$\checkmark$ SADERF-SGQ 36.79 63.55 77.31 59.22 1.61 3.52 3.57 2.90 1.44 2.86 2.69 2.33
$\checkmark$ SADERF-FastFoodL 28.63 64.64 77.61 56.96 2.81 5.89 5.92 4.87 0.92 1.83 1.79 1.51
4.2 Comparison of Weight Matrix-Component Function Alternatives using Spectraformer
Our Spectraformer comparison results are available in Table 2. We now analyze the performance of component functions and weight matrices. We perform our analysis on both general trends of component functions across weight matrices and vice versa. We also make use of mean statistics (i.e., mean accuracy, mean training time, and mean peak memory consumption) which we give in brackets in the subsequent discussion.
Figure 3: Peak memory consumption (x-axis), accuracy (y-axis) on three textual LRA [15] tasks: ListOps, Text, Retrieval. Color indicates training time, the darker the faster, the brighter the slower (see legend). Models compared are the best Spectraformer variant (PosRF-MM) (boldened), the OT, and other efficient Transformers (RF and non-RF) in the literature. Values for non-RF Transformers are taken from [16].
Component functions. In terms of accuracy, SADERF (58.82%) typically performs better than OPRF (58.53%) and OPRF performs better than PosRF (57.50%) by small margin, agreeing with [12, 13]. However, SADERF (3.21 hours, 2.19 GB) and OPRF (3.17 hours, 2.24 GB) consistently take longer to train and have higher memory consumption than PosRF (2.71%, 1.76 GB) across weight matrices. This is likely due to OPRF and SADERF having additional computation of the L2 norm of the weight matrix which PosRF does not calculate.
Weight matrices. In terms of accuracy, SGQ (59.70%), ORF (59.64%), and QMC (59.59%) have the best performance. This is followed by MM (57.93%), FastFoodL (57.41%), and the worst being SORF (55.43%). In terms of training time, most weight matrices are rather uniform: SORF (2.64 hours) and MM, SGQ, QMC, ORF all at 2.69 hours. FastFoodL takes significant amount of time (4.78 hours). For peak memory consumption, FastFoodL is the most efficient at 1.73 GB whilst all other methods maintain a uniform value of 2.13 GB. FastFoodL reduces the space complexity at the cost of the time complexity. SORF reduces training time whilst compromising to a small extent its accuracy. This is interesting as it violates our initial assumptions that learners should perform better than approximators.
Best model. Overall, we find that PosRF-MM has the best performance, with the top accuracy (60.19%), top-3 training time (2.35 hours) and top-3 peak memory consumption (1.86 GB), outperforming the SOTA SADERF-ORF [13] (59.86%, 2.90 hours, and 2.33 GB) and OPRF-ORF [12] (59.25%, 2.86 hours, 2.21 GB) in the existing RF Transformer family. PosRF-MM achieves this performance due to the combination of both an efficient component function (PosRF) and performative weight matrix (MM). PosRF-MM has higher average accuracy (60.19) than the previous SOTA, SADERF-ORF (59.86), whilst having 23.4% faster training time, and 25.2% lower memory consumption. Our result demonstrates the viability of Spectraformer as a unifying random feature framework of Transformer.
Table 3: Comparison of the best Spectraformer variant (PosRF-MM), highlighted in red, against the OT and other efficient Transformers (RF and non-RF) in the literature. We report three metrics: accuracy (on the test set), time (training time (hour)), and memory (peak memory consumption (GB)). $L$ stands for ListOps, $T$ for Text, $R$ for Retrieval, and $\mu$ for the average of the three. Best results for each section and task are boldened.
Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
Model L T R $\mu$ L T R $\mu$ L T R $\mu$
Original Transformer [1] 38.37 61.95 80.69 60.34 2.24 4.30 8.33 4.96 5.37 10.37 10.77 8.84
Nyströmformer [17] 38.51 64.83 80.52 61.29 0.71 0.71 1.29 0.90 1.37 1.21 2.39 1.66
Linformer [35] 37.45 58.93 78.19 58.19 0.60 0.65 1.13 0.79 0.99 0.99 1.89 1.29
Informer [36] 32.53 62.64 77.57 57.58 1.19 1.60 2.91 1.90 4.85 5.12 5.77 5.25
Reformer [33] 37.68 62.93 78.99 59.87 0.85 0.94 1.73 1.17 1.61 1.61 2.98 2.07
BigBird [37] 39.25 63.86 80.28 61.13 1.88 2.00 3.81 2.56 2.71 2.83 4.97 3.50
Skyformer [16] 38.69 64.70 82.06 61.82 1.29 1.02 1.86 1.39 1.75 1.59 3.15 2.16
PosRF-ORF [8] 37.35 61.60 80.53 59.83 1.30 2.84 2.89 2.34 1.17 2.31 2.10 1.86
OPRF-ORF [12] 37.50 59.35 80.90 59.25 1.61 3.45 3.50 2.86 1.36 2.71 2.56 2.21
SADERF-ORF [13] 37.35 61.37 80.87 59.86 0.00 3.52 3.58 2.36 0.00 2.86 2.69 1.85
PosRF-FastFoodL [14] 24.95 64.76 76.37 55.36 2.69 5.59 5.61 4.63 0.78 1.56 1.53 1.29
Spectraformer (PosRF-MM) 38.05 61.85 80.67 60.19 1.31 2.84 2.89 2.35 1.17 2.31 2.10 1.86
4.3 Comparison with other efficient Transformers
When comparing with other non-random feature Transformers, we observe that PosRF-MM performs comparatively against the OT whilst reducing memory consumption and training time significantly (see Figure 3 and Table 3). On ListOps and Retrieval, PosRF-MM and other Spectraformer variants perform on par with other non-RF efficient Transformers. Results for non-RF efficient Transformers (above the horizontal divider) in Table 3 are taken from [16]. Spectraformer does not outperform more performant Nyström-based transformers which are not based on random features. Having said that, we are the first to have conducted a comprehensive analysis of random feature methods in the Transformer via the Spectraformer framework, comparing it against other past works and the OT. Spectraformer can be extended in the future to include a larger range of component function and weight matrix options. We discuss limitations, and broader impact of our research in Appendix A.1.
5 Conclusion
We present Spectraformer, a framework for approximating and learning the attention kernel in the Transformer. Our paper generalizes past works and presents empirical findings on different component function and weight matrix combinations. We experiment with 18 combinations, of which 14 are novel. Novel combinations outperform the previous SOTA in the family for multiple tasks. Our best model, PosRF-MM, has higher average accuracy (60.19) than the previous SOTA, SADERF-ORF (59.86), whilst having 23.4% faster training time, and 25.2% lower memory consumption. This shows the viability of Spectraformer as a unifying random feature framework of Transformer.
Appendix A Appendix
A.1 Future work, limitations, and broader impact
Spectraformer provides an experimental framework for random features in the Transformer. However, we have not provided formal mathematical proofs for the bounds the these combinations due to the number of combinations discussed. We hope Spectraformer inspires future work to generalizes the proof in such a framework.
Spectraformer has covered a significant number of combinations, however, we have not covered all weight matrices and component functions in the literature. Spectraformer can serve as a foundation for further random feature Transformer research using the techiques not yet experimented with but covered in Appendices A.3, A.4, A.5.
Due to the significant number of experimented models, we are unable to perform hyper-parameter fine-tuning. Future work can benefit from extending the scope of the current research to other tasks, i.e., other non-textual LRA tasks (image classification, path-finder, and path-x) and short-context datasets/ tasks such as GLUE [38]. Transformer-based architecture can also benefit from Spectraformer. Past random feature Transformer works [12, 13] have experimented with conformer and vision Transformer with ViT. [39] provides a good foundation to extend our work. Spectraformer, although currently experimented in the Transformer setting, can also extend to non-parametric kernel classification. It could also benefit from ablation studies like the comparison variance in [14] or in [40]
Given the complex nature of language models in the ethical, energy, and societal domains, our work should be used responsibly  [41]
A.2 Related Work
Spectraformer is based on several theoretical assumptions and choice of mathematical framework and approach. This section discusses alternatives to the choices of Spectraformer. Specifically, Appendix A.2.1 discusses alternative kernelized formulations to Section 2.1; Appendix A.2.2 discusses alternative kernel approximation techniques to random features in Section A.2.2. There are also some component functions and weight matrices which are not covered in the experiments. We introduce them in the related work section in Appendices A.3, A.4, A.5.
A.2.1 Alternative Kernelized Attentions
In addition to the kernelized formulation of attention as shown in Section 2.1, based on our literature survey, we have identified the following three alternative kernel-based formulations of attention.
Nadaraya-Watson kernel estimator in integration form [42]: Given the Gaussian kernel, where the probability of a single variable and the joint probability of two variables are estimated using kernel density estimation with the Gaussian kernel, then we have the following equation:
[…]
Equation 12 lays the foundation for FourierFormer - a Transformer with attention based on this formulation [42].
Gaussian decomposition [43]: the attention mechanism can also be decomposed into the Gaussian kernel directly from Equation 4 without appealing to the use of kernel estimator [43].
$A_{i}=\sum_{j=1}^{N}\frac{1}{Z_{l}(\boldsymbol{q}_{i},\boldsymbol{K})}exp(\frac{-||\boldsymbol{q}_{i}-\boldsymbol{k}_{j}||^{2}_{2}}{2\sqrt{d_{k}}})exp(\frac{||\boldsymbol{q}_{i}||^{2}_{p=2}+||\boldsymbol{k}_{j}||^{2}_{p=2}}{2\sqrt{d_{k}}})\boldsymbol{v}_{j}$ (13)
Equation 13 leads to Implicit Kernel Attention (IKA) - a Transformer with attention based on this formulation [43].
Non-local operation [44] is popular decomposition of attention in computer vision. This is the basis for Vision Transformer and related work. Non-local operation itself a generalisation of the non-local means used in image denoising [45]. Non-local operation is essentially an operation that operates on the entire input range, instead of a specific window which is characteristic of convolutional operations (see Equation 14). Indeed, the attention vector output $\boldsymbol{y}$ can be defined using $f$ which is the softmax (see Equation 15) in the non-local operation.
$\mathbf{\boldsymbol{y}}_{i}=\frac{1}{C(\boldsymbol{x})}\sum_{\forall j}f(\boldsymbol{x}_{i},\boldsymbol{x}_{j})g(\boldsymbol{x}_{j})$ (14)
$\mathbf{\boldsymbol{y}}=softmax(\boldsymbol{x}^{T}\boldsymbol{W}_{\theta}^{T}\boldsymbol{W}_{\theta}\boldsymbol{x})g(\boldsymbol{x})$ (15)
We will show that the non-local operation in Equation 14 is a generalisation of the Nadaraya-Watson kernel estimator. Specifically, when defining $C(\boldsymbol{x})=\sum_{l}\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{l}),f=\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{l}),g=\boldsymbol{v}_{j}$, then we obtain the original Nadaraya-Watson kernel estimator in Equation 4. The variety of kernel decompositions of attention demonstrate the robustness of the kernel interpretation of attention and hence opening up an alley in attention improvement through the lens of kernel.
A.2.2 Kernel Approximation Beyond Random Features
Kernel approximation is presented in Section 2.3 through the use of random features. However, there has been other kernel approximation techniques outside of random features. These include greedy basis selection [46], divide and conquer [47, 48], and nyström methods [49] (which led to the development of Skyformer [16] and Nyströmformer [17]).
Figure 4: Spectraformer weight matrices. Red tick denotes techniques which have been explored in the literature. Green tick denotes techniques which are explored in this work.
A.3 Technical Details of Weight Approximators
Spectraformer weight matrices are shown in Figure 4. Weight matrices are either approximators or learners. There are three approaches to approximators: Monte Carlo sampling, Unit-cube sampling, and quadrature. We have discussed these approaches in detail in Section 3.1. Here we provide a more specific classification schemes for each method.
  • Monte Carlo sampling-based weight matrices are improvements over Base and subdivided into two types: structural and geometric methods. Structural methods decompose the Base random matrices $P$ into smaller matrices which reduce time and/ or space complexity. Structural matrices includes FastFoodF, SCRF and P-Model. Among these, we have explored the learner variant FastFoodL of FastFood. Geometric methods enforce certain geometrical couplings along some dimensions in the construction of random matrices to reduce the approximation error. Geometric methods include: ROM (ORF, SORF) and SimRF. Among these, we have explored the main ROM variants - ORF and SORF.
  • Unit-cube sampling-based weight matrices reformulate the Bochner integral in Equation 9 as Equation A.3 with uniform and independent points. This approach includes QMC, MM, and SSF. We have not covered SSF in our experiments.
  • Quadrature-based weight matrices estimates the intractable integral using non-uniform weights with deterministic rules. There are two approaches: SGQ and SSR. We have not covered SSR in our experiments.
We now explore these weight matrices in the order specified above.
Base [19] is the direct sampling from a distribution $\boldsymbol{P}_{ij}\sim p$, corresponding to the kernel $\mathcal{K}$ without any additional adjustment, and in approximating the RBF, $P=G$ where $G$ is the normal distribution.
$W_{Base}=\frac{1}{\sigma}\boldsymbol{P}$ (16)
FastFoodF [28] makes use of Hadamard and diagonal matrices to speed up Gaussian matrices $G$ constructions for RFF in $O(n\log{d})$ (100 times faster) with $O(n)$ space (1000 times less space) given $n\geq d$ against Base. However, FastFoodF increases variance and approximation error and decreases the concentration bound.
$W_{FastFood_{L}}=\frac{1}{\sigma}\boldsymbol{SHG\Pi HB}$ (17)
[…]
SCRF (Signed Circulant Matrix Projection) [50] maintains a high matrix construction speed and $O(s)$ space ($s=t*d$) whilst maintaining the variance of Base, an improvement over FastFoodF. It does this by making use of circulant matrices.
$\boldsymbol{W}_{SCRF}=[\boldsymbol{P}^{(1)},\boldsymbol{P}^{(2)},...,\boldsymbol{P}^{(t)}],t=d|D$ (18)
where $\boldsymbol{P}^{(i)}$ is a signed circulant Gaussian matrix, with the first column of circulant random matrix $\boldsymbol{C}^{(i)_{[d]}}$ of $\boldsymbol{P}^{(i)}$ drawn randomly from a Gaussian distribution $\mathcal{N}(\boldsymbol{0},\boldsymbol{I}/\sigma^{2})$. The circulant matrix via Discrete Fourier Transform can be defined as:
$\begin{array}[]{cc}\boldsymbol{C}=\frac{1}{m}\boldsymbol{F}^{*}diag(\boldsymbol{Fc})\boldsymbol{F}\\ \boldsymbol{F}=[e^{i\frac{2\pi}{m}kn}]^{m-1}_{k,n=0}\end{array}$ (19)
P-Model [51] generalizes over both circulant matrices (including SCRF), FastFoodF, along with other Gaussian and semi-Gaussian structured matrix methods.
$\boldsymbol{W}_{P_{model}}=[\boldsymbol{g^{T}P_{1}},\boldsymbol{g^{T}P_{2}},...,\boldsymbol{g^{T}P_{s}}]^{T}$ (20)
where $\boldsymbol{g}\in\mathbb{R}^{a}$ is a Gaussian vector and a sequence of $m$ $\boldsymbol{P}_{i}\in\mathbb{R}^{a*d}$ matrices with L2 norm columns, where $P_{i}$ represents a P-model. Semi-Gaussian structured matrices including Toeplitz, Hankel, and other Toeplitz-like matrices are an open area of research for random matrices.
ORF (Orthogonal RF) [9] decreases the approximation error significantly compared to Base. This is done via replacing $G$ with a properly scaled random orthogonal matrix. However, generating orthogonal matrices become costly quickly as the number of dimensions increases.
$\boldsymbol{W}_{ORF}=\frac{1}{\sigma}\boldsymbol{SQ}$ (21)
where $\boldsymbol{Q}$ is a uniformly distributed random orthogonal matrix (on the Stiefel manifold) obtained from the QR decomposition of $\boldsymbol{G}$, the set of rows of $\boldsymbol{Q}$ forming a basis in $\mathbb{R}^{d}$, and $\boldsymbol{S}$ is a diagonal matrix with entries sampled i.i.d from the $\chi$-distribution with $d$ degrees of freedom, thus making the rows of $\boldsymbol{SQ}$ and $\boldsymbol{G}$ identically distributed.
SORF (Structured ORF) [9] decreases time and space complexity of ORF (from $O(d^{2})$ to $O(d\log{d})$ with almost no extra memory cost) by imposing structure on the orthogonal matrices, inspired by structural methods. SORF is unbiased with large $d$. We replace $\boldsymbol{S}$ with $\sqrt{d}$ and $\boldsymbol{Q}$ with structured matrix $\boldsymbol{HD}_{1}\boldsymbol{HD}_{2}\boldsymbol{HD}_{3}$
$\boldsymbol{W}_{SORF}=\frac{\sqrt{d}}{\sigma}\boldsymbol{HD}_{1}\boldsymbol{HD}_{2}\boldsymbol{HD}_{3}$ (22)
where $\boldsymbol{D}_{i}\in\mathbb{R}^{d*d},i=1,2,3$ is diagonal ’sign-flipping‘ matrices with diagonal entries sampled from the Rademacher distribution, $\boldsymbol{H}$ is the normalized Walsh-Hadamard matrix.
ROM (Random Orthogonal Embeddings) [52] generalizes SORF ($t=3$) to:
$\boldsymbol{W}_{ROM}=\frac{\sqrt{d}}{\sigma}\prod_{i=1}^{t}\boldsymbol{SD}_{i}$ (23)
where $\boldsymbol{S}$ is a class of $L_{2}$-normalized versions of Kronecker product matrices, of which the Hadamard $\boldsymbol{H}$ is a representative of, and $\boldsymbol{D}$ being independent diagonal matrices. When $\boldsymbol{S}=\boldsymbol{H}$ and $t=3$, we obtain SORF used in our experiments. Another ROM variant is S-Hybrid, which we do not experiment with, given as
$\boldsymbol{W}_{S_{Hybrid}}=\boldsymbol{SD}_{k}^{(\mathcal{U})}\prod_{i=1}^{k-1}\boldsymbol{SD}_{i}^{(\mathcal{R})}$ (24)
where $\boldsymbol{D}_{k}^{(\mathcal{U})}$ is a diagonal matrix with i.i.d. $Unif(\mathcal{S}^{1})$ on the unit circle of $\mathbb{C}$. When being used in the component function, instead of calculating $\phi(\boldsymbol{x})^{T}\phi(\boldsymbol{y})$ in Equation 10. We model the equation as $Real(\overline{\phi(\boldsymbol{x})}^{T}\phi(\boldsymbol{y}))$.
SimRF (Simplex RF) [11] is developed to be a positive RF with the most optimal solution, the MSE of which is the lowest in the geometric methods. This is an alternative to the ROM-based geometric methods. Two variants of SimRF are: weight-independent and weight-dependent. SimRF is defined as:
$\boldsymbol{W}_{simp}=\boldsymbol{DSR}$ (25)
where $\boldsymbol{D}\in\mathbb{R}^{d*d}=diag(\boldsymbol{\omega}_{i}),\boldsymbol{R}\in\mathbb{R}^{d*d}$ being a random orthogonal matrix drawn from Haar measure on $O(d)$ and $\boldsymbol{S}\in\mathbb{R}^{d*d}$ with row $\boldsymbol{s}_{i}$:
$\boldsymbol{s}_{i}=\begin{cases}\sqrt{\frac{d}{d-1}}\boldsymbol{e}_{i}-\frac{\sqrt{d}+1}{(d-1)^{3/2}}(1,...,1,0)^{T}&1\leq i<d\\ \frac{1}{\sqrt{d-1}}(1,1,...,1,0)^{T}&i=d\end{cases}$ (26)
being unit vectors which are manifestly normalized and subtend obtuse angles. SimRF variants are defined based on how $\boldsymbol{\omega}_{i}$ is constructed. For SimRFindep, $\boldsymbol{\omega}_{i}\sim\chi_{d}$. For Simdep, we permit the random vector direction $\{\hat{\boldsymbol{\omega}}_{i}\}$ to be correlated with norms $\{\boldsymbol{\omega}_{i}\}$:
$\boldsymbol{\omega}_{i}=-\frac{\sum_{j\neq i}\omega_{j}}{||\sum_{j\neq i}\omega_{j}||_{2}}\boldsymbol{\omega}_{i},i=1,...,d$ (27)
QMC (Quasi-Monte Carlo) [25] evaluates on a low discrepancy sequence (e.g., Halton, Sobol’ Faure, and Niederreiter) of points instead of random points in Monte Carlo. Although the approximation error is only reduced minimally, QMC has been shown to perform better than MC in high dimensions and does not have undesirable clustering effect. To calculate QMC, we first assume that $p(\boldsymbol{x})=\prod^{d}_{j=1}p_{j}(\boldsymbol{x}_{j})$ factorizes with respect to the dimensions with $p_{j}(.)$ being a univariate density function. Then we define:
$\Phi^{-1}(\boldsymbol{t})=\left(\Phi_{1}^{-1}\left(\boldsymbol{t}_{1}\right),\ldots,\Phi_{d}^{-1}\left(\boldsymbol{t}_{d}\right)\right)\in\mathbb{R}^{d}$ (28)
where $\Phi_{j}$ being the cumulative distribution function of $p_{j}$, $\boldsymbol{t}_{1},\boldsymbol{t}_{2},...,\boldsymbol{t}_{s}\in[0,1]^{d}$ being a low discrepancy sequence, and $\boldsymbol{\omega}_{i}=\Phi^{-1}(\boldsymbol{t}_{i})$. We can thus transform the integral on $\mathbb{R}^{d}$ in Equation 9 to an integral on the unit cube $[0,1]^{d}$ as
$\mathcal{K}(\boldsymbol{x}-\boldsymbol{x}^{\prime})=\int_{[0,1]^{d}}\exp\big{(}\mbox{i}(\boldsymbol{x}-\boldsymbol{x}^{\prime})^{\!\top}\Phi^{-1}(\boldsymbol{t})\big{)}\mbox{d}\boldsymbol{t}$
The weight matrix then can be defined as:
$\boldsymbol{W}_{\text{QMC}}=[\Phi^{-1}(\boldsymbol{t}_{1}),\Phi^{-1}(\boldsymbol{t}_{2}),\ldots,\Phi^{-1}(\boldsymbol{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}\;.$
QMC can be further improved by a sub-grouped based rank-one lattice construction which improved complexity [53].
MM (Moment Matching) [26] [7] improves over QMC by removing the undesirable clustering effect and having the same approximation error with less features. This is done by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$:
$\boldsymbol{W}_{\text{MM}}=[\widetilde{\Phi}^{-1}(\boldsymbol{t}_{1}),\widetilde{\Phi}^{-1}(\boldsymbol{t}_{2}),\ldots,\widetilde{\Phi}^{-1}(\boldsymbol{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}$
where $\widetilde{\Phi}^{-1}(\boldsymbol{t}_{i})=\tilde{\boldsymbol{A}}^{-1}(\Phi^{-1}(\boldsymbol{t}_{i})-\tilde{\boldsymbol{\mu}})$ can be constructed using moment matching with sample mean $\tilde{\boldsymbol{\mu}}=\frac{1}{s}\sum_{i=1}^{s}\Phi^{-1}(\boldsymbol{t}_{i})$ and the square root of the sample covariance matrix $\tilde{\boldsymbol{A}}$ satisfying $\tilde{\boldsymbol{A}}\tilde{\boldsymbol{A}}^{\!\top}=\text{Cov}(\Phi^{-1}(\boldsymbol{t}_{i})-\tilde{\boldsymbol{\mu}})$.
SSF (Spherical Structured Features) [54] improves over QMC by including rotation-invariant kernel as well as in terms of complexity. Rotation-invariant property suggests the construction of feature maps using spherical equal weight approximation using Riesz s-energy on a d-dimensional sphere $\mathbb{S}^{d}:=\{\boldsymbol{x}\in\mathbb{R}^{d+1}|||\boldsymbol{x}||_{2}=1\}$. Specifically, we construct $\{\boldsymbol{v}_{i}\}^{s}_{i=1}$ asymptotically uniformly distributed on the sphere and obtain $\boldsymbol{V}:=[\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{s}]\in\mathbb{S}^{(d\times s}$, that $\Phi^{-1}(t)$ uses the one-dimensional QMC point, we have the weight matrix:
$\boldsymbol{W}_{\text{SSF}}=[\Phi^{-1}(t)\boldsymbol{v}_{1},\Phi^{-1}(t)\boldsymbol{v}_{2},\ldots,\Phi^{-1}(t)\boldsymbol{v}_{s}]^{\!\top}\in\mathbb{R}^{s\times(d+1)},$
SGQ (Sparse Grid Quadrature) [27] evolves from GQ. GQ (Gaussian Quadrature) [27] assumes that the kernel $k$ factorizes with respect to the dimensions and thus can be approximated by a one-dimensional Gaussian quadrature rule. However the total number of points $s$ scale exponentially with the dimensions. However GQ suffers form the curse of dimensionality. This is alleviated using Smolyak rule resulting in SGQ. Assuming the third-degree SGQ using symmetric univariate quadrature points $\left\{-\hat{p}_{1},0,\hat{p}_{1}\right\}$ with weights $(\hat{a}_{1},\hat{a}_{0},\hat{a}_{1})$, then we have:
$\boldsymbol{W}_{\text{SGQ}}\!=\![\boldsymbol{0}_{d},\hat{p}_{1}\boldsymbol{e}_{1},\ldots,\hat{p}_{1}\boldsymbol{e}_{d},-\hat{p}_{1}\boldsymbol{e}_{1},\ldots,-\hat{p}_{1}\boldsymbol{e}_{d}]^{\!\top}\!\in\!\mathbb{R}^{(2d+1)\times d}\!$
given that $\mathbb{e}_{i}$ is the d-dimensional standard basis vector with the ith element being 1.
SSR (Spherical-Radial Rules) [55] transforms Equation 10 into a double integral over a hyper-sphere and the real line. This leads to the following approximation:
$\boldsymbol{W}_{\text{SSR}}=\boldsymbol{\vartheta}\otimes\begin{bmatrix}(\boldsymbol{QV})^{\top}\\ -(\boldsymbol{Q}\boldsymbol{V})^{\top}\end{bmatrix}\in\mathbb{R}^{(2d+1)\times d},$
where $\boldsymbol{\vartheta}=[\vartheta_{1},\vartheta_{2},\ldots,\vartheta_{s}]$, $\boldsymbol{V}=[\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{d+1}]$, $\vartheta\sim\chi(d+2)$ and $\{\boldsymbol{v}_{i}\}_{i=1}^{d+1}$ being vertices of a unit regular d-simplex randomly rotated by $\boldsymbol{Q}$ (a random orthogonal matrix).
A.4 Technical Details of Weight Learners
We give a non-mathematical definition of weight matrix learner in Section 3.1: weight matrix learner is any function which parameterize distributions. This allows for the discovery and future experimentation with other potential weight matrix learners in Spectraformer and random feature method in general. Weight matrix learner comes from a long line of work in kernel methods, namely, kernel learning via random features. Typically, a component function $f$ is chosen and a weight matrix $\boldsymbol{W}$ is produced via kernel approximation, providing us with the feature map $\phi$. $\phi$ can then be combined with input $\boldsymbol{x}$ and weight $\boldsymbol{v}$ to fit the learning objective. Kernel learning, however, imposes an additional objective: learning the weight matrix $\boldsymbol{W}$.
The approach to the dual objectives separates the kernel learning methods into two-stage and one-stage methods [7]. Two-stage methods (e.g., [56], [23], [57]) solves the dual objective separately: $\boldsymbol{W}$ is typically first learned via a kernel alignment scheme [58], then we solve for $\boldsymbol{v}$. One-stage methods (e.g., [14], [22]), on the other hand, solves the dual objective in parallel: $p(.)$ corresponding to $\boldsymbol{W}$ is parameterized then $\boldsymbol{v}$ is solved typically. It is not plausible to apply kernel alignment [58] on Transformer architectures. Therefore, only one-stage methods can be considered in the context of weight matrix learner. Two kernel learning methods, GMM and FastFoodL, have been experimented in the non-Transformer setting [29], then adapted into the Transformer setting [14]. DGM (Deep Generative Model) has also been effective in modeling distributions [59]. It too has been adapted by [14] for the Transformer setting. Due to the superior performance of FastFoodL in [14], we have only experimented with this technique. We now explore them.
FastFoodL (see Equation 17) is proposed to be learnable by [29] with two variations: FSARD (scaling matrix $\boldsymbol{S}$ previously sampled from chi-squared is made learnable) and FSGBARD (optimizes the marginal likelihood with respect to the diagonal matrices $\boldsymbol{G}$ and $\boldsymbol{B}$). It is further adapted by [14] to make $\boldsymbol{S}$ or $\boldsymbol{S},\boldsymbol{G},\boldsymbol{B}$ learnable parameters. In our experiments, we maintain the set up by [14].
GMM (Gaussian Mixture Model) [23, 24] is a known universal approximator of probability distribution. Specifically, we can consider $\boldsymbol{W}_{GMM}$ being composed of $s$ components, with each component $c$ (with $C$ total components) being a Gaussian with mean vector $\mu_{c}$ and covariance matrix $\Sigma_{c}$ and component weight $\pi_{c}$. $p(\boldsymbol{\omega})$ can then be approximated as $p(\boldsymbol{\omega})=\sum_{c=1}^{C}\pi_{c}\mathcal{N}(\mu_{c},\Sigma_{c})$ [14]. Assuming $\pi_{c}=\frac{1}{C}$, the weight matrix with learnable weights $\Sigma$ and $\mu$ can be parameterized as [14]:
$\boldsymbol{W}_{GMM}=[\Sigma_{1}n+\mu_{1},...,\Sigma_{C}n+\mu_{C}]^{T},n\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ (29)
Another formulation of GMM is proposed in [29] in combination with FastFood, albeit it extends beyond constructing the weight matrix.
DGM (Deep Generative Model) [59] allows for a generic distribution modeling using neural network. DGM is initially proposed to make differentiating sampling processes possible via the reparametrization trick, where the generator $g$ creates samples from the target distribution $\boldsymbol{\omega}_{m}=g(\boldsymbol{n}_{m})$ via transforming from a noise distribution $\boldsymbol{n}_{m}\sim p_{0}(.)$.
In addition to the techniques discussed above, there are also some other prominent one-stage methods (e.g., INN, BBQ). INN (Invertible Neural Network) [60] is another universal approximator of probability distribution, consisting of invertible operations transforming samples from a known distribution to a different more complex distribution. BBQ (Black Box Quantile) [22] follows the QMC scheme and models $\Phi^{-1}$ using parameterized quantile function.
A.5 Technical Details of Component Functions
Figure 5: Spectraformer component functions. Red tick denotes techniques which have been explored in the literature. Green tick denotes techniques which are explored in this work.
Component functions discussed in Section 3.2 all belong to CDERF. However, there are other component functions including DIRF and HRF (see Figure 5) which we do not explore in our experiments. We now investigate the theoretical details of these component functions, the CDERF family is introduced in their order of discovery.
CDERF (Complex dense exponential random feature) is given in Equation 30, where $\boldsymbol{\omega},\boldsymbol{x}\in\mathbb{R}^{d}$, $p=\mathcal{N}(0,1)^{d}$, $k\in\{1,2\}$ and $\mathbb{S}_{d_{c}}$ being a set of $d*d$ complex symmetric matrices. The parameters and constraints of different CDERF functions are specified in Table 4:
$f_{DE}^{(k)}(\boldsymbol{\omega},\boldsymbol{x})=Real(D\exp(\boldsymbol{\omega}^{T}\boldsymbol{A}\boldsymbol{\omega}+\boldsymbol{\omega}^{T}\boldsymbol{B}^{(k)}\boldsymbol{x}+\boldsymbol{x}^{T}\boldsymbol{C}^{(k)}\boldsymbol{x}))$ (30)
Table 4: CDERF component functions
$\boldsymbol{A}$ $\boldsymbol{B}^{(1)}$ $\boldsymbol{B}^{(2)}$ $\boldsymbol{C}^{(1)}$ $\boldsymbol{C}^{(2)}$ $D$ $\boldsymbol{x}$ $\boldsymbol{y}$
[…]
SADERF [13] $A_{GE}\boldsymbol{I}_{d}$ $B_{GE}\boldsymbol{I}_{d}$ $C_{GE}\boldsymbol{I}_{d}$ $D_{GE}$ $\Psi\boldsymbol{x}$ $\Psi^{-1}\boldsymbol{y}$
GERF [12] $A_{GE}\boldsymbol{I}_{d}$ $B_{GE}^{(1)}\boldsymbol{I}_{d}$ $B_{GE}^{(2)}\boldsymbol{I}_{d}$ $C_{GE}\boldsymbol{I}_{d}$ $D_{GE}$ $\boldsymbol{x}$ $\boldsymbol{y}$
TrigRF [19] $0\boldsymbol{I}_{d}$ $i\boldsymbol{I}_{d}$ $-i\boldsymbol{I}_{d}$ $0\boldsymbol{I}_{d}$ $1$ $\boldsymbol{x}$ $\boldsymbol{y}$
PosRF [8] $0\boldsymbol{I}_{d}$ $1\boldsymbol{I}_{d}$ $-1\boldsymbol{I}_{d}$ $1$ $\boldsymbol{x}$ $\boldsymbol{y}$
OPRF [12] $0\boldsymbol{I}_{d}$ $1\boldsymbol{I}_{d}$ $-1\boldsymbol{I}_{d}$ $1$ $\boldsymbol{x}$ $\boldsymbol{y}$
TrigRF (Trigonometric RF) [19] is the base RFF implementation.
$\begin{array}[]{c}f_{TrigRF}^{(1)}(\boldsymbol{\omega},\boldsymbol{x})=\sqrt{2}\cos(\boldsymbol{\omega}^{T}\boldsymbol{x}+\boldsymbol{b})\\ f_{TrigRF}^{(2)}(\boldsymbol{\omega},\boldsymbol{y})=\sqrt{2}\sin(\boldsymbol{\omega}^{T}\boldsymbol{y}+\boldsymbol{b})\\ \boldsymbol{b}\sim Uniform(0,2\pi)\end{array}$ (31)
The use of the trigonometric sine and cosine functions leads to unstable behavior when the inputs have negative dimension-values. This can be further exacerbated when the values TrigRF try to approximate are close to 0 (since most values are of low significance). This causes the variance to approach infinity [8, 12]. Therefore, we do not want to use TrigRF to perform kernel approximation in attention.
PosRF (Positive RF) [8] fixes the problem of TrigRF by enforcing positive component function output in the softmax. The variance of PosRF (in contrast to the variance of TrigRF approaching infinity) approaches 0 as the approximated value of the Softmax kernel approaches 0. PosRF has two forms: Positive RF Base (PosRF-B) (see Equation 32) and Positive RF Hyperbolic (PosRF-Hyp), which is multi-component, i.e., Equation 11: $l=2$, (see Equation 33).
$\begin{array}[]{c}f_{PosRF_{B}}(\boldsymbol{\omega},\boldsymbol{x})=\exp(\boldsymbol{\omega}^{T}\boldsymbol{x}-\frac{||\boldsymbol{x}||^{2}}{2})\end{array}$ (32)
$\begin{array}[]{ll}f_{1,PosRF_{Hyp}}(\boldsymbol{\omega},\boldsymbol{x})=\exp(\boldsymbol{\omega}^{T}\boldsymbol{x}-||\boldsymbol{x}||^{2})\\ f_{2,PosRF_{Hyp}}(\boldsymbol{\omega},\boldsymbol{x})=\exp(-\boldsymbol{\omega}^{T}\boldsymbol{x}-||\boldsymbol{x}||^{2})\end{array}$ (33)
We only use PosRF-B in our experiments.
GERF (Generalized exponential RF) [12] generalizes both TrigRF and PosRF with Equation 34.
$\begin{array}[]{c}f_{GERF}^{(1)}(\boldsymbol{\omega},\boldsymbol{x})=D\exp(A||\boldsymbol{\omega}||^{2}+B\boldsymbol{\omega}^{T}\boldsymbol{x}+\boldsymbol{C}||\boldsymbol{x}||^{2})\\ f_{GERF}^{(2)}(\boldsymbol{\omega},\boldsymbol{y})=D\exp(\boldsymbol{A}||\boldsymbol{\omega}||^{2}+sB\boldsymbol{\omega}^{T}\boldsymbol{y}+\boldsymbol{C}||\boldsymbol{y}||^{2})\\ \\ Re(1-4A)>0,\boldsymbol{B}=\sqrt{s(1-4\boldsymbol{A})}\\ \boldsymbol{C}=-(s+1)/2,D=(\sqrt[4]{1-4\boldsymbol{A}})^{d}\\ \boldsymbol{A}\in\mathbb{C},s\in\{-1,+1\}\end{array}$ (34)
where $\sqrt{.}$ and .$n]{.}$ denoting a principal root with a complex argument
OPRF (Optimized positive RF) [12] is the solution to the minimization of the variance of GERF. Specifically it is defined as Equation 34 with $s=+1$, $||\boldsymbol{x}+\boldsymbol{y}||^{2}>0$, and $\boldsymbol{A}\in\mathbb{R}$ defined in terms of $p^{*}$ as:
$\begin{array}[]{cc}\boldsymbol{A}=(1-1/p^{*})/8\\ p^{*}=(\sqrt{(2||\boldsymbol{x}+\boldsymbol{y}||^{2}+d)^{2}+8d||\boldsymbol{x}+\boldsymbol{y}||^{2}}-2||\boldsymbol{x}+\boldsymbol{y}||^{2}-d)\\ /(4||\boldsymbol{x}+\boldsymbol{y}||^{2})\end{array}$ (35)
Whilst PosRF is not bounded, OPRF is. OPRF can provide $e^{60}$$\times$ variance reduction in estimating the Softmax compared to TrigRF [12].
DERF (Dense-exponential random features) [13] extends GERF and replace $\boldsymbol{A},\boldsymbol{B},\boldsymbol{C}$ with dense matrices. DERF is CDERF when $\boldsymbol{B},C$ are in the real instead of the complex plane. Minimizing the variance of DERF leads to two approaches: ADERF and SDERF. However both these approaches rely on SVD and eigen decompositions which are not extensively supported on GPU and deep learning libraries. Therefore SADERF is proposed.
SADERF (simplified ADERF) [13] is a special case of ADERF and extends GERF, requiring only basic unary operations in addition.
$\begin{array}[]{c}f_{SADE}^{(1)}(\boldsymbol{\omega},\boldsymbol{x})=f_{GE}^{(1)}(\boldsymbol{\omega},\Psi\boldsymbol{x}),f_{SADE}^{(2)}(\boldsymbol{\omega},\boldsymbol{y})=f_{GE}^{(2)}(\boldsymbol{\omega},\Psi^{-1}\boldsymbol{y})\\ \Psi_{l,l}^{*}=(\sum_{j}(\boldsymbol{y}_{l}^{(i)})^{2}/\sum_{i}(\boldsymbol{x}_{l}^{(i)})^{2})^{1/4}\end{array}$ (36)
DIRF (Discretely-induced random features) [12] are based on the assumption of $p(\boldsymbol{\omega})$ being a discrete distribution with $\boldsymbol{\omega}_{1},...,\boldsymbol{\omega}_{d}$ being i.i.d., $\mathbb{P}(\boldsymbol{\omega}_{l}=k)=p_{k}$, $\sum_{k=0}^{\infty}p_{k}=1$, $p_{k}>0,k\in\{0\}\cup\mathbb{N}$. By making use of the Taylor series, we can define DIRF as:
$f_{DI}(\boldsymbol{\omega},\boldsymbol{x})=\exp(-\frac{||\boldsymbol{x}||^{2}}{2})\prod_{l=1}^{d}\boldsymbol{x}_{i}^{\boldsymbol{\omega}_{l}}(\boldsymbol{\omega}_{l}!)^{\frac{-1}{2}}p_{\boldsymbol{\omega}_{l}}^{\frac{-1}{2}}$ (37)
DIRF has two variations: PoisRF (Poisson RF) which requires $p(.)$ to be the Poisson distribution and GeomRF (Geometric RF) which requires $p(.)$ to be the geometric distribution. Since the many weight matrices introduced in Section 3.1 do not readily approximate these distributions, we decide to leave DIRF out of the paper.
HRF (Hybrid RF) [10] combines multiple base estimators, similar to multiple kernel learning methods [61], in order to provide the most accurate approximation in regions of interest. Choromanski et al. on solving softmax approximation using HRF. HRF is constructed using a weighted combination $\epsilon$ of $p+1$ base estimators $\hat{SM}^{k}(\boldsymbol{x},\boldsymbol{y})$ and $\lambda$-coefficients, $\Lambda$ $p$ binary estimators $\lambda^{k}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow[0,1]$, both constructed independently:
$\hat{SM}^{\epsilon,\Lambda}=\sum_{k=1}^{p}\hat{\lambda}^{k}(\boldsymbol{x},\boldsymbol{y})\hat{SM}^{k}(\boldsymbol{x},\boldsymbol{y})+(1-\sum_{k=1}^{p}\hat{\lambda}^{k}(\boldsymbol{x},\boldsymbol{y}))\hat{SM}^{p+1}(\boldsymbol{x},\boldsymbol{y})$ (38)
Since HRF requires the base estimators $SM(\boldsymbol{x},\boldsymbol{y})$ to be unbiased, the easiest solution is the complex exponential $\hat{SM}^{cexp}_{m}$ which can be derived directly from the expectation of the softmax as follows:
$\begin{array}[]{cc}\hat{SM}^{cexp}_{m}(\boldsymbol{x},\boldsymbol{y})=\psi^{m}_{A}(\boldsymbol{x})^{T}\psi^{m}_{(\boldsymbol{A}^{T})^{-1}}(\boldsymbol{y})\\ \psi^{m}_{\boldsymbol{M}}(u)=\frac{1}{\sqrt{m}}\exp(-\frac{(\boldsymbol{Mu})^{2}}{2})\\ (\exp(\boldsymbol{\omega}^{T}_{i}\boldsymbol{Mu}),...,\exp(\boldsymbol{\omega}_{m}^{T}\boldsymbol{Mu}))^{T},\boldsymbol{\omega}_{i}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d})\end{array}$ (39)
To construct $\lambda$-coefficients, three methods are proposed: simultaneous accurate approximation of both small and large values, Gaussian Lambda Coefficients, and adaption to data admitting clustering structure.
A.6 Experimental Details in Spectraformer
Table 5: Hyper-parameters for Spectraformer, following the setting from [16] due to limited computational power. $L$ stands for ListOps, $T$ for Text, and $R$ for Retrieval. The seed that can be used to reproduce our experiments is 42.
L T R
Embedding dim. 64 64 64
Transformer dim 64 64 64
Hidden dim 128 128 128
Head dim 32 32 32
Num. heads 2 2 2
Num. layers 2 2 2
Vocabulary size 32 512 512
Sequence length 2000 4000 4000
Dropout rate 0.1 0.1 0.1
Att. dropout rate 0.1 0.1 0.1
Pooling mode mean
Batch size 32 32 16
Learning rate 0.0001 0.0001 0.0002
Warmup steps 1000 80 800
Learning rate decay linear
Weight decay 0 0 0
Evaluation freq. 500 500 1000
Num. epochs 50k 50K 50k
Num. init steps 1k 3K 3k
Num. eval steps 62 200 300
Patience 10 10 10
Num. features 128 128 128

### v3

@@
 Past methods use a subset of combinations of component functions and weight matrices within the random features paradigm.
-We identify the need for a systematic comparison of different combinations of weight matrix and component functions for attention learning in Transformer.
+We identify the need for a systematic comparison of different combinations of weight matrices and component functions for attention learning in Transformer.
 In this work, we introduce Spectraformer, a unified framework for approximating and learning the kernel function in linearized attention of the Transformer.
 We experiment with broad classes of component functions and weight matrices for three textual tasks in the LRA benchmark.
-Our experimentation with multiple combinations of component functions and weight matrices leads us to a novel combination with 23.4% faster training time and 25.2% lower memory consumption over the previous SOTA random feature Transformer, while maintaining the performance, as compared to the Original Transformer.
-Our code is available at: https://github.com/dukeraphaelng/spectraformer.
+Our empirical findings indicate that different kernels are good at different tasks and that kernel choice is fundamental to performant models.
+Our code is available at: https://anonymous.4open.science/r/spectraformer-8A97.
 1 Introduction
-Transformer [1] has revolutionized the landscape of natural language processing (NLP) and forms the basis of almost all state-of-the-art language models.
-Its influence has reached beyond NLP into computer vision [2], speech processing [3] and other fields.
+Transformer (Vaswani et al., 2017) has revolutionized the landscape of natural language processing (NLP) and forms the basis of almost all state-of-the-art language models.
+Its influence has reached beyond NLP into computer vision (Han et al., 2023), speech processing (Lin et al., 2022) and other fields.
 Compared to its precursors like the LSTMs, Transformer leverages parallelism since it is fully based on the attention mechanism.
@@
 The softmax serves the role of a compatibility function, which captures the degree of compatibility between two tokens (see Equations 1 and 2).
-Extensive experiments with other compatibility functions show that the softmax is not the only possibility, and that the ‘best’ compatibility function depends on the task [4].
-Replacing the softmax formulation with alternatives may allow to reduce the time complexity of attention computation.
+Extensive experiments with other compatibility functions show that the softmax is not the only possibility, and that the ‘best’ compatibility function depends on the task (Tsai et al., 2019).
+Replacing the softmax formulation with alternatives may allow us to reduce the time complexity of attention computation.
 In this paper, we focus on the subset of alternatives that employs kernel functions.
-Kernel functions have been historically used in machine learning algorithms to simplify the estimation of parameters [5].
-[4] show that attention can be formulated using kernels and then be linearized [6], given the feature map of the respective attention kernel.
-Random feature-based algorithms [7] are a family of algorithms, inspired by spectral analysis, which provides the feature map associated with a kernel.
+Kernel functions have been historically used in machine learning algorithms to simplify the estimation of parameters (Hofmann et al., 2008).
+Tsai et al.
+(2019) show that attention can be formulated using kernels and then be linearized (Katharopoulos et al., 2020), given the feature map of the respective attention kernel.
+Random feature-based algorithms (Liu et al., 2022) are a family of algorithms, inspired by spectral analysis, which provides the feature map associated with a kernel.
 A random feature consists of three components: component function, weight matrix, and component weight (see Section 3).
-Due to their robustness and flexibility, random features give rise to ‘random feature Transformers’, the family of Transformers with linearized attention via random features (hereby abbreviated as RF), first with the work by [8].
-This is seen to lead to three strands of research: (A) Optimization of weight matrices either by making the matrices have lower time or space complexity, or by incorporating beneficial properties like orthogonality to reduce variance [9, 8, 10, 11]; (B) Enhancement of component functions by either: (i) finding a function with a tighter variance or approximation bound, or (ii) be engineered to output with numerical stability and be bounded [8, 12, 13]; (C) Parameterization of weight matrices to perform kernel learning instead of kernel approximation [14].
+Due to their robustness and flexibility, random features give rise to ‘random feature Transformers’, the family of Transformers with linearized attention via random features (hereby abbreviated as RF), first with the work by Choromanski et al.
+(2021).
+This is seen to lead to three strands of research: (A) Optimization of weight matrices either by making the matrices have lower time or space complexity, or by incorporating beneficial properties like orthogonality to reduce variance (Yu et al., 2016; Choromanski et al., 2021; 2022; Reid et al., 2023); (B) Enhancement of component functions by either: (i) finding a function with a tighter variance or approximation bound, or (ii) be engineered to output with numerical stability and be bounded (Choromanski et al., 2021; Likhosherstov et al., 2022; 2023); (C) Parameterization of weight matrices to perform kernel learning instead of kernel approximation (Chowdhury et al., 2022).
 (A) and (B) seek to improve approximation quality via approximation error and variance reduction.
@@
 Additionally, past works only explore certain combinations of component functions and weight matrices, creating disjoint overlaps and gaps in the literature.
-Finally, the weight matrices are often those explicitly designed in the context of kernelized attention in the Transformer [9, 8, 10, 11], ignoring those which are popular in the kernel methods literature (e.g., [7]).
+Finally, the weight matrices are often those explicitly designed in the context of kernelized attention in the Transformer (Yu et al., 2016; Choromanski et al., 2021; 2022; Reid et al., 2023), ignoring those which are popular in the kernel methods literature (e.g., Liu et al.
+(2022)).
 These limitations necessitate a unified framework that allows the comparison of different combinations in search of the best-performing one.
-Figure 1: Peak memory consumption (x-axis), average accuracy (y-axis) on three textual LRA [15] tasks: ListOps, Text, Retrieval.
-Color indicates training time, the darker the faster, the brighter the slower (see legend).
-Models compared are the best Spectraformer variant (PosRF-MM) (boldened), the OT, and other efficient Transformers (RF and non-RF) in the literature.
-Values for non-RF Transformers are taken from [16].
 We introduce Spectraformer, a unified framework of random feature-based attention in Transformer.
@@
 We will open-source our codebase upon acceptance.
-Through our benchmarking experiments on three textual Long Range Arena [15] tasks and 18 combinations of weight matrices and component functions, Spectraformer is able to find novel combinations that outperform existing random feature-based Transformers in terms of performance, training time, and peak memory consumption.
-Figure 1 compares our best-known Spectraformer combination (PosRF-MM) against the previous SOTA models in random feature Transformer, efficient Transformers, and the OT.
-The x and y axes denote peak memory and average accuracy, while the colour of the circle indicates the memory consumption.
-Spectraformer (specifically, PosRF-MM) is more accurate than the previous SOTA models, SADERF-ORF (a.k.a., FAVOR# [13]), OPRF-ORF (a.k.a., FAVOR++ [12]), PosRF-ORF (a.k.a., FAVOR+ [8]) and maintains a comparable accuracy against the OT, whilst significantly reducing training time and memory consumption over both.
-However, it underperforms in accuracy and training time against Skyformer [16] and Nyströmformer [17].
+Through our benchmarking experiments on three textual Long Range Arena (Tay et al., 2021) tasks and 18 combinations of weight matrices and component functions, Spectraformer enables novel combinations of component functions and weight matrices from the kernel literature to be experimented easily in the Transformer setting.
+This process makes possible the discovery of combinations that outperform existing random feature-based Transformers in terms of performance, training time, and peak memory consumption.
+Various novel Spectraformer combinations are more accurate than the previous SOTA models, SADERF-ORF (a.k.a., FAVOR# (Likhosherstov et al., 2023)), OPRF-ORF (a.k.a., FAVOR++ (Likhosherstov et al., 2022)), PosRF-ORF (a.k.a., FAVOR+ (Choromanski et al., 2021)), whilst significantly reducing training time and memory consumption over both.
 We hope that Spectraformer will enable the mechanism to benchmark more random feature-based Transformers for these and other tasks.
@@
   2. 2.
-    We demonstrate how Spectraformer can discover novel SOTA models in the random feature Transformer family for the LRA benchmark.
-Our best model has comparable accuracy to the OT whilst significantly reducing training time and peak memory consumption.
+    We demonstrate how Spectraformer can discover novel SOTA models in the random feature Transformer family for the LRA benchmark and that these models are more performant than existing random feature Transformers.
   3. 3.
-    This is the first work to compare random feature Transformers with other efficient Transformers.
+    We empirically identify the relationship between performant combinations and task characteristics, aiding the right kernel choice when tackling a task.
+  4. 4.
+    Our code is currently available for review, but they will be publicly released upon acceptance.
+We detail guide on how to incorporate future kernel works into Spectraformer.
+With this work, every future kernel work that is introduced can easily be incorporated into the transformer.
 2 Preliminaries
+We first cover the theoretical foundations to this work. A list of commonly used acronyms is available in Table 1.
+Table 1: Table of acronyms
+Acronym Description
+OT (Vaswani et al., 2017) Original Transformer
+RF (Rahimi & Recht, 2007) Random Feature method
+RBF Radial Basis Function kernel (also called the Gaussian kernel)
+TrigRF (Rahimi & Recht, 2007) Trigonometric RF (component function)
+PosRF (Choromanski et al., 2021) Positive RF (component function)
+OPRF (Likhosherstov et al., 2022) Optimized positive RF (component function)
+SADERF (Likhosherstov et al., 2023) Simplified Asymmetric Dense-Exponential RF (component function)
+ORF (Yu et al., 2016) Orthogonal RF (weight matrix)
+SORF (Yu et al., 2016) Structural ORF (weight matrix)
+QMC (Avron et al., 2016) Quasi- Monte Carlo (weight matrix)
+MM (Shen et al., 2017) (Liu et al., 2022) Moment Matching (weight matrix)
+SGQ (Dao et al., 2017) Sparse Grid Quadrature (weight matrix)
+FastFoodL(Yang et al., 2015) FastFood Learnable (weight matrix)
 2.1 Kernelized Attention
-OT attention [1] can be defined using the unified attention model [18] as follows.
+OT attention (Vaswani et al., 2017) can be defined using the unified attention model (Galassi et al., 2021) as follows.
 Given a collection of inputs and a learning objective, attention mechanism learns the attention matrix A which captures the associative information between pairs of tokens in a sequence of length $N$, with each row $\textbf{A}_{i}$ being the representation of a token $i$ with every other token $j\in[1,N]$.
-This is done by creating three learnable representations of each token $i$ called query ($\boldsymbol{q}_{i}$), key ($\boldsymbol{k}_{i}$), and value ($\boldsymbol{v}_{i}$).
-The attention representation $\textbf{A}_{i}$ is calculated as a weighted sum of value $\boldsymbol{v}_{j}$ given attention weights $a_{j}$.
+This is done by creating three learnable representations of each token $i$ called query ($\bm{q}_{i}$), key ($\bm{k}_{i}$), and value ($\bm{v}_{i}$).
+The attention representation $\textbf{A}_{i}$ is calculated as a weighted sum of value $\bm{v}_{j}$ given attention weights $a_{j}$.
 The general equation for $\textbf{A}_{i}$ is given as:
-$\begin{array}[]{cccc}\boldsymbol{A}_{i}=\sum_{j=1}^{N}\frac{f(\boldsymbol{q}_{i},\boldsymbol{k}_{j})}{\sum_{l}f(\boldsymbol{q}_{i},\boldsymbol{k}_{l})}\boldsymbol{v}_{j}=\sum_{j=1}^{N}a_{j}\boldsymbol{v}_{j};&a=g(e);&e=f(\boldsymbol{q}_{i},\boldsymbol{k}_{j})\end{array}$ (1)
+$\begin{array}[]{cccc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{f(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}f(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}=\sum_{j=1}^{N}a_{j}\bm{v}_{j};&a=g(e);&e=f(\bm{q}_{i},\bm{k}_{j})\end{array}$ (1)
 Attention weights are generated using a distribution function $g$ applied on an energy score $e$.
-The energy score is captured by applying a compatibility function on a query $\boldsymbol{q}_{i}$ and a set of keys K, in order to compute a score of a specific relationship between the pairs.
+The energy score is captured by applying a compatibility function on a query $\bm{q}_{i}$ and a set of keys K, in order to compute a score of a specific relationship between the pairs.
 The attention in the OT is called the ‘scaled dot-product’ which defines the compatibility function $f$ and the distribution function $g$ as follows:
-$\begin{array}[]{cc}f(\boldsymbol{q}_{i},\textbf{K})=\frac{\boldsymbol{q}_{i}^{T}\textbf{K}}{\sqrt{d_{k}}}&g(\boldsymbol{x}_{j})=\frac{\exp(\boldsymbol{x}_{j})}{\sum_{l}\exp(\boldsymbol{x}_{l})}\end{array}$ (2)
+$\begin{array}[]{cc}f(\bm{q}_{i},\textbf{K})=\frac{\bm{q}_{i}^{T}\textbf{K}}{\sqrt{d_{k}}}&g(\bm{x}_{j})=\frac{\exp(\bm{x}_{j})}{\sum_{l}\exp(\bm{x}_{l})}\end{array}$ (2)
 We can then have the attention matrix (with $\exp(.)$ being applied element-wise) as:
-$\begin{array}[]{cc}\boldsymbol{A}_{i}=\sum_{j=1}^{N}\frac{\exp(\boldsymbol{q}_{i}\boldsymbol{k}_{j}^{T}/\sqrt{d_{k}})}{\sum_{l}\exp(\boldsymbol{q}_{i}\boldsymbol{k}_{l}^{T}/\sqrt{d_{k}})}\boldsymbol{v}_{j}&A=softmax(\frac{\boldsymbol{QK}^{T}}{\sqrt{d_{k}}})\boldsymbol{V}\end{array}$ (3)
-with $\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V}$ being the matrix of queries, keys, and values respectively with each row being $\boldsymbol{q}_{i},\boldsymbol{k}_{i},\boldsymbol{v}_{i}$.
+$\begin{array}[]{cc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{\exp(\bm{q}_{i}\bm{k}_{j}^{T}/\sqrt{d_{k}})}{\sum_{l}\exp(\bm{q}_{i}\bm{k}_{l}^{T}/\sqrt{d_{k}})}\bm{v}_{j}&A=softmax(\frac{\bm{QK}^{T}}{\sqrt{d_{k}}})\bm{V}\end{array}$ (3)
+with $\bm{Q},\bm{K},\bm{V}$ being the matrix of queries, keys, and values respectively with each row being $\bm{q}_{i},\bm{k}_{i},\bm{v}_{i}$.
 We now discuss the intrinsic connection between kernel and attention.
-With the compatibility function $f$ as a kernel $f(\boldsymbol{q}_{i},\boldsymbol{k}_{j})=\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{j})$, Equation 1 can be rewritten as:
-$\boldsymbol{A}_{i}=\sum_{j=1}^{N}\frac{\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{j})}{\sum_{l}\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{l})}\boldsymbol{v}_{j}$ (4)
-As shown in [4], this is indeed the Nadaraya-Watson Kernel Estimator, with $\mathbb{E}_{p(\boldsymbol{k}_{j}|\boldsymbol{q}_{i})}[\boldsymbol{v}_{j}|X=\boldsymbol{k}_{j}]=\boldsymbol{A}_{i}$, $l_{i}\left(\boldsymbol{k}_{j})=p(\boldsymbol{k}_{j}|\boldsymbol{q}_{i}\right)$, $Y_{i}=\boldsymbol{v}_{j}$:
-$\begin{array}[]{ccc}\mathbb{E}_{l_{i}}[Y_{i}|X=\boldsymbol{u}]=\sum_{i=1}^{N}l_{i}(\boldsymbol{u})Y_{i};&l_{i}(\boldsymbol{u})=\frac{\mathcal{K}(\frac{\boldsymbol{u}-\boldsymbol{x}_{i}}{h})}{\sum_{j=1}^{n}\mathcal{K}(\frac{\boldsymbol{u}-\boldsymbol{x}_{j}}{h})}\end{array}$ (5)
+With the compatibility function $f$ as a kernel $f(\bm{q}_{i},\bm{k}_{j})=\mathcal{K}(\bm{q}_{i},\bm{k}_{j})$, Equation 1 can be rewritten as:
+$\bm{A}_{i}=\sum_{j=1}^{N}\frac{\mathcal{K}(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}\mathcal{K}(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}$ (4)
+As shown in Tsai et al.
+(2019), this is indeed the Nadaraya-Watson Kernel Estimator, with $\mathbb{E}_{p(\bm{k}_{j}|\bm{q}_{i})}[\bm{v}_{j}|X=\bm{k}_{j}]=\bm{A}_{i}$, $l_{i}\left(\bm{k}_{j})=p(\bm{k}_{j}|\bm{q}_{i}\right)$, $Y_{i}=\bm{v}_{j}$:
+$\begin{array}[]{ccc}\mathbb{E}_{l_{i}}[Y_{i}|X=\bm{u}]=\sum_{i=1}^{N}l_{i}(\bm{u})Y_{i};&l_{i}(\bm{u})=\frac{\mathcal{K}(\frac{\bm{u}-\bm{x}_{i}}{h})}{\sum_{j=1}^{n}\mathcal{K}(\frac{\bm{u}-\bm{x}_{j}}{h})}\end{array}$ (5)
 Specifically, the scaled dot-product attention in the OT in Equation 3 corresponds to the softmax kernel (see Equation 6).
-$\mathcal{K}_{softmax}(\boldsymbol{q}_{i},\boldsymbol{k}_{j})=\exp(\boldsymbol{q}_{i}^{T}\boldsymbol{k}_{j}/\sqrt{d_{k}})$ (6)
+$\mathcal{K}_{softmax}(\bm{q}_{i},\bm{k}_{j})=\exp(\bm{q}_{i}^{T}\bm{k}_{j}/\sqrt{d_{k}})$ (6)
 2.2 Linearizing Kernelized Attention via Random Features
-Since a kernel is the inner product of two vectors in some space with respect to some feature map $\phi$, i.e., $\mathcal{K}(\boldsymbol{x},\boldsymbol{y})=\phi(\boldsymbol{x})\phi(\boldsymbol{y})^{T}$, then we can rewrite Equation 4 as:
-$\begin{array}[]{c}\boldsymbol{A}_{i}=\frac{\sum_{j}\phi(\boldsymbol{q}_{i})\phi(\boldsymbol{k}_{j})^{T}\boldsymbol{v}_{j}}{\sum_{l}\phi(\boldsymbol{q}_{i})\phi(\boldsymbol{k}_{l})^{T}}=\frac{\phi(\boldsymbol{q}_{i})\sum_{j}\phi(\boldsymbol{k}_{j})\otimes\boldsymbol{v}_{j}}{\phi(\boldsymbol{q}_{i})\sum_{l}\phi(\boldsymbol{k}_{l})^{T}}\end{array}$ (7)
-If we pre-compute $\sum_{j}\phi(\boldsymbol{k}_{j})\otimes\boldsymbol{v}_{j}$ and $\sum_{l}\phi(\boldsymbol{k}_{l})^{T}$, the entire term becomes $O(1)$ and we only need to compute Equation 7 $N$ times for each $i$, resulting in an attention matrix calculated in linear time $O(N)$.
-We now refer to Equation 7 as linearized attention [6].
+Since a kernel is the inner product of two vectors in some space with respect to some feature map $\phi$, i.e., $\mathcal{K}(\bm{x},\bm{y})=\phi(\bm{x})\phi(\bm{y})^{T}$, then we can rewrite Equation 4 as:
+$\begin{array}[]{c}\bm{A}_{i}=\frac{\sum_{j}\phi(\bm{q}_{i})\phi(\bm{k}_{j})^{T}\bm{v}_{j}}{\sum_{l}\phi(\bm{q}_{i})\phi(\bm{k}_{l})^{T}}=\frac{\phi(\bm{q}_{i})\sum_{j}\phi(\bm{k}_{j})\otimes\bm{v}_{j}}{\phi(\bm{q}_{i})\sum_{l}\phi(\bm{k}_{l})^{T}}\end{array}$ (7)
+If we pre-compute $\sum_{j}\phi(\bm{k}_{j})\otimes\bm{v}_{j}$ and $\sum_{l}\phi(\bm{k}_{l})^{T}$, the entire term becomes $O(1)$ and we only need to compute Equation 7 $N$ times for each $i$, resulting in an attention matrix calculated in linear time $O(N)$.
+We now refer to Equation 7 as linearized attention (Katharopoulos et al., 2020).
 All that is left is to retrieve the feature map $\phi$, the process of which is called kernel approximation (see Section 2.3), which corresponds to the kernels we specify.
 There are many popular kernels from which we can choose, but typically, we want to approximate the softmax kernel in Equation 6.
-However, since the softmax relates to the $\mathcal{K}_{RBF}(\boldsymbol{x},\boldsymbol{y})=\exp(-\frac{||\boldsymbol{x}-\boldsymbol{y}||^{2}}{2})$ via the following equation:
-$\mathcal{K}_{softmax}(\boldsymbol{x},\boldsymbol{y})=\exp(\frac{||\boldsymbol{x}||^{2}}{2})\mathcal{K}_{RBF}(\boldsymbol{x},\boldsymbol{y})\exp(\frac{||\boldsymbol{y}||^{2}}{2})$ (8)
+However, since the softmax relates to the $\mathcal{K}_{RBF}(\bm{x},\bm{y})=\exp(-\frac{||\bm{x}-\bm{y}||^{2}}{2})$ via the following equation:
+$\mathcal{K}_{softmax}(\bm{x},\bm{y})=\exp(\frac{||\bm{x}||^{2}}{2})\mathcal{K}_{RBF}(\bm{x},\bm{y})\exp(\frac{||\bm{y}||^{2}}{2})$ (8)
 Approximating the RBF is easier than the softmax directly in the method we introduce in Section 2.3, hence we approximate the softmax indirectly via the RBF in Equation 8.
-Oftentimes, however, we might not want to specify a kernel, since kernel choice is indeed a hyperparameter and the softmax is by no means essential or necessary as shown in [4].
+Oftentimes, however, we might not want to specify a kernel, since kernel choice is indeed a hyperparameter and the softmax is by no means essential or necessary as shown in Tsai et al.
+(2019).
 Then kernel learning would be a more suitable option, the detail of which is discussed in Section 2.4.
@@
 2.3 Random Features for Kernel Approximation
-Of the kernel approximation family of methods, the most successful and appropriate for linearized attention in Transformer has been the random features approach [19].
-Estimating kernels using random features relies on a fundamental insight from harmonic analysis called Bochner’s theorem [20] stated as follows:
-A continuous shift-invariant kernel $\mathcal{K}(\boldsymbol{x},\boldsymbol{y})=\mathcal{K}(\boldsymbol{x}-\boldsymbol{y})$ on $\mathbb{R}^{d}$ is positive definite if and only if $\mathcal{K}(\boldsymbol{\delta})$ is the Fourier transform of a non-negative measure $p(\boldsymbol{\omega})$.
-If $\mathcal{K}(\boldsymbol{\delta})$ is properly scaled, that is $\mathcal{K}(0)=1$ then the measure $p(\boldsymbol{\omega})$ is a proper probability distribution.
+Of the kernel approximation family of methods, the most successful and appropriate for linearized attention in Transformer has been the random features approach (Rahimi & Recht, 2007).
+Estimating kernels using random features relies on a fundamental insight from harmonic analysis called Bochner’s theorem (Rudin, 1990) stated as follows:
+A continuous shift-invariant kernel $\mathcal{K}(\bm{x},\bm{y})=\mathcal{K}(\bm{x}-\bm{y})$ on $\mathbb{R}^{d}$ is positive definite if and only if $\mathcal{K}(\bm{\delta})$ is the Fourier transform of a non-negative measure $p(\bm{\omega})$.
+If $\mathcal{K}(\bm{\delta})$ is properly scaled, that is $\mathcal{K}(0)=1$ then the measure $p(\bm{\omega})$ is a proper probability distribution.
 inline]define unambiguously the dimension of x and phi here, also the dimension before and after approximation of the features map
-$\mathcal{K}(\boldsymbol{\delta})=\int p(\boldsymbol{\omega})\exp(i\boldsymbol{\omega\delta})\mathrm{d}\boldsymbol{\omega}$ (9)
-When approximating the integration in this equation via Monte Carlo sampling with $s$ samples, $\boldsymbol{\omega}\sim p(\boldsymbol{\omega})$, we can obtain the feature map:
-$\begin{array}[]{ll}\mathcal{K}(\boldsymbol{\delta})&=\mathbb{E}[\phi_{\boldsymbol{\omega}}(\boldsymbol{x})^{T}\phi_{\boldsymbol{\omega}}(\boldsymbol{y})]\approx<\phi_{\boldsymbol{\omega}}(\boldsymbol{x}),\phi_{\boldsymbol{\omega}}(\boldsymbol{y})>\\ \phi_{\boldsymbol{\omega}}(\boldsymbol{x})\!&=\!\frac{1}{\sqrt{s}}\big{[}\exp(-\mbox{i}{\boldsymbol{\omega}}^{\!\top}_{1}\boldsymbol{x}),\ldots,\exp(-\mbox{i}{\boldsymbol{\omega}}^{\!\top}_{s}\boldsymbol{x})]^{\!\top}\end{array}$ (10)
+$\mathcal{K}(\bm{\delta})=\int p(\bm{\omega})\exp(i\bm{\omega\delta})\mathrm{d}\bm{\omega}$ (9)
+When approximating the integration in this equation via Monte Carlo sampling with $s$ samples, $\bm{\omega}\sim p(\bm{\omega})$, we can obtain the feature map:
+$\begin{array}[]{ll}\mathcal{K}(\bm{\delta})&=\mathbb{E}[\phi_{\bm{\omega}}(\bm{x})^{T}\phi_{\bm{\omega}}(\bm{y})]\approx<\phi_{\bm{\omega}}(\bm{x}),\phi_{\bm{\omega}}(\bm{y})>\\ \phi_{\bm{\omega}}(\bm{x})\!&=\!\frac{1}{\sqrt{s}}\big{[}\exp(-\mbox{i}{\bm{\omega}}^{\!\top}_{1}\bm{x}),\ldots,\exp(-\mbox{i}{\bm{\omega}}^{\!\top}_{s}\bm{x})]^{\!\top}\end{array}$ (10)
 2.4 Random Features for Kernel Learning
@@
 However, kernels may be chosen using heuristics and by convention from a popular subset.
-Kernel is very much a hyperparameter.
-This could pose a problem since there is no free lunch in learning as much as in kernel choice [21].
-[4] has shown experimental results that this is also true in the context of attention in Transformer.
+The kernel is very much a hyperparameter.
+This could pose a problem since there is no free lunch in learning as much as in kernel choice (Schölkopf & Smola, 2018).
+Tsai et al.
+(2019) has shown experimental results that this is also true in the context of attention in Transformer.
 Hence, instead of picking a kernel for the attention, we could learn the kernel via kernel learning an established kernel methods technique.
-Kernel learning has shown to be effective in general [22, 23, 24] and in the case of attention via Gaussian Mixture Model (GMM), learnable weight matrices (FastFood), deep generative model (DGM) [14].
-Kernel learning is achieved by making the weight matrix $\boldsymbol{W}$ learnable where $\boldsymbol{\omega}_{i}$ is the output of a function parameterizing $p(\boldsymbol{\omega})$.
+Kernel learning has shown to be effective in general (Tompkins et al., 2019; Wilson & Adams, 2013; Oliva et al., 2016) and in the case of attention via Gaussian Mixture Model (GMM), learnable weight matrices (FastFood), deep generative model (DGM) (Chowdhury et al., 2022).
+Kernel learning is achieved by making the weight matrix $\bm{W}$ learnable where $\bm{\omega}_{i}$ is the output of a function parameterizing $p(\bm{\omega})$.
 The kernelized attention via random features discussed here is a popular approach, however it is by no means the only one.
@@
 3 Spectraformer
-Spectraformer is a unifed random feature framework that allows for the combination of any weight matrix with any component function to be used as kernelized attention in Transformer.
+Spectraformer is a unified random feature framework that allows for the combination of any weight matrix with any component function to be used as kernelized attention in Transformer.
 The general equation for random feature is shown in Equation 11, generalized from Equation 10, based on Liu et al.
-[7], Choromanski et al.
-[8].
-$\begin{array}[]{rl}\phi_{\boldsymbol{\omega}}(\boldsymbol{x})=&\frac{1}{\sqrt{s}}\big{[}a_{1}f_{1}(\boldsymbol{\omega}_{1},\boldsymbol{x}),...,a_{s}f_{1}(\boldsymbol{\omega}_{s},\boldsymbol{x}),...,\\ &a_{1}f_{l}(\boldsymbol{\omega}_{1},\boldsymbol{x}),...,a_{s}f_{l}(\boldsymbol{\omega}_{s},\boldsymbol{x})]^{\!\top}\\ \boldsymbol{W}=&[\boldsymbol{\omega}_{1},...,\boldsymbol{\omega}_{s}]^{T}\in\mathbb{R}^{s*d}\end{array}$ (11)
+(2022); Choromanski et al.
+(2021).
+$\begin{array}[]{rl}\phi_{\bm{\omega}}(\bm{x})=&\frac{1}{\sqrt{s}}\big{[}a_{1}f_{1}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{1}(\bm{\omega}_{s},\bm{x}),...,\\ &a_{1}f_{l}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{l}(\bm{\omega}_{s},\bm{x})]^{\!\top}\\ \bm{W}=&[\bm{\omega}_{1},...,\bm{\omega}_{s}]^{T}\in\mathbb{R}^{s*d}\end{array}$ (11)
 Spectraformer is shown in Figure 1.
-The attention of the OT (on the left-hand side) is replaced by the attention of the Spectraformer which pre-computes the Hadamard (element-wise) product of values $\boldsymbol{v}$ and transformed keys $\phi(\boldsymbol{k})$ scaled by $\phi(\boldsymbol{k})$, the product of which is multiplied with the transformed queries $\phi(\boldsymbol{q})$.
+The attention of the OT (on the left-hand side) is replaced by the attention of the Spectraformer which pre-computes the Hadamard (element-wise) product of values $\bm{v}$ and transformed keys $\phi(\bm{k})$ scaled by $\phi(\bm{k})$, the product of which is multiplied with the transformed queries $\phi(\bm{q})$.
 The linear map $\phi$ is composed from weight matrices and component functions.
-The linear map has $s$ components with $a=1$ (see Equation 11).
-The figure and the equation together highlight the three strands of research as mentioned previously in Section 1: (A) Weight matrix approximator: constructing $\boldsymbol{W}$ more effectively; (B) Weight matrix learner: parameterizing $\boldsymbol{W}$ instead of sampling; (C) Component function: constructing $f_{j}$ more effectively.
+The linear map has $s$ components with $a$ being the component weight (see Equation 11).
+The component weight $a$ can either be fixed (typically at $a=1$) or learnable as proposed by various kernel works (Liu et al., 2022).
+Due to the fact that the majority of random feature Transformers (e.g., Likhosherstov et al.
+(2022)) have $a=1$, we follow this setting for consistent comparison between combinations.
+The figure and the equation together highlight the three strands of research as mentioned previously in Section 1: (A) Weight matrix approximator: constructing $\bm{W}$ more effectively; (B) Weight matrix learner: parameterizing $\bm{W}$ instead of sampling; (C) Component function: constructing $f_{j}$ more effectively.
 3.1 Weight Matrices
-In Spectraformer, the weight matrix $\boldsymbol{\boldsymbol{W}}$ which is either an approximator (from the families of (e.g., ORF [9], SORF [9]), Unit-cube (e.g., QMC [25], MM [26]), Quadrature (e.g., SGQ [27]).
-or a learner (FastFoodL [14]).
+In Spectraformer, the weight matrix $\bm{W}$ which is either an approximator (from the families of (e.g., ORF and SORF (Yu et al., 2016)), Unit-cube (e.g., QMC (Avron et al., 2016), MM (Shen et al., 2017)), Quadrature (e.g., SGQ (Dao et al., 2017)).
+or a learner (FastFoodL (Chowdhury et al., 2022)).
 Figure 1: Spectraformer framework based on Equation 11
-We term the matrix $\boldsymbol{W}$ as weight matrix instead of random matrix since although $\boldsymbol{W}$ acts like a random matrix ($\boldsymbol{\omega}_{i}\sim p(.)$, see Equation 10), it is not guaranteed to be one.
-In the case of weight matrix approximator, they approximate a random matrix and in the case of weight matrix learner, they parameterize a weight matrix, thereby implicitly learning a distribution $p(.)$ associated with such $\boldsymbol{W}$.
+We term the matrix $\bm{W}$ as weight matrix instead of random matrix since although $\bm{W}$ acts like a random matrix ($\bm{\omega}_{i}\sim p(.)$, see Equation 10), it is not guaranteed to be one.
+In the case of weight matrix approximator, they approximate a random matrix and in the case of weight matrix learner, they parameterize a weight matrix, thereby implicitly learning a distribution $p(.)$ associated with such $\bm{W}$.
 The weight matrix is given in Definition 3.1.
 Definition 3.1.
-Provided that a matrix can substitute for a random matrix in Equation 11, and the solution for this equation given such matrix and $f$ as TrigRF being the solution for Equation 9, i.e., as an unbiased estimator of $\mathcal{K}$, and TrigRF being $f_{1}=f_{l}=\exp(-i\boldsymbol{\omega}^{T}\boldsymbol{x})$ (see Equation 10) then such a matrix is a weight matrix.
-It follows from Definition 3.1 that for any weight matrix $\boldsymbol{W}$, $f=TrigRF$ is a valid component function.
+Provided that a matrix can substitute for a random matrix in Equation 11, and the solution for this equation given such matrix and $f$ as TrigRF being the solution for Equation 9, i.e., as an unbiased estimator of $\mathcal{K}$, and TrigRF being $f_{1}=f_{l}=\exp(-i\bm{\omega}^{T}\bm{x})$ (see Equation 10) then such a matrix is a weight matrix.
+It follows from Definition 3.1 that for any weight matrix $\bm{W}$, $f=TrigRF$ is a valid component function.
 Spectraformer enables weight matrix approximation in terms of three families based on how they solve the intractable integral in Equation 9:
-  • Monte Carlo sampling-based weight matrix involves approximating the integral of Equation 9 using Monte Carlo sampling with the solution given by Equation 10 [19] with $\boldsymbol{W}$ being the ‘Base’ random matrix of $p(.)$, $\boldsymbol{\omega}_{i}\sim p(.)$.
-The random matrix $\boldsymbol{W}$ can be constructed more efficiently (either reducing time or space complexity) with FastFood [28].
-The random matrix $\boldsymbol{W}$ can also be enforced to form geometrical couplings such as orthogonality to reduce approximation error with ORF, SORF [9].
+  • Monte Carlo sampling-based weight matrix involves approximating the integral of Equation 9 using Monte Carlo sampling with the solution given by Equation 10 (Rahimi & Recht, 2007) with $\bm{W}$ being the ‘Base’ random matrix of $p(.)$, $\bm{\omega}_{i}\sim p(.)$.
+The random matrix $\bm{W}$ can be constructed more efficiently (either reducing time or space complexity) with FastFood (Le et al., 2013).
+The random matrix $\bm{W}$ can also be enforced to form geometrical couplings such as orthogonality to reduce approximation error with ORF, SORF (Yu et al., 2016).
 Both these methods produce valid random matrices.
   • Unit-cube sampling-based weight matrix transforms Equation 9 to an integral on the unit cube $[0,1]^{d}$, then performs the approximation with uniform and independent points, thus reducing variance.
-The first approach is QMC [25], where $\boldsymbol{W}:\boldsymbol{\omega}_{i}=\Phi^{-1}(\boldsymbol{t}_{i})$, $\Phi$ being the cumulative distribution function (CDF) associated with $p(.)$ and $\boldsymbol{t}_{i}$ being a low discrepancy sequence.
-$\Phi^{-1}(\boldsymbol{t}_{i})\sim p(.)$ whilst having a lower variance than direct sampling from $p(.)$.
-MM [26, 7] improves over QMC by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$ on $\Phi^{-1}$.
+The first approach is QMC (Avron et al., 2016), where $\bm{W}:\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$, $\Phi$ being the cumulative distribution function (CDF) associated with $p(.)$ and $\bm{t}_{i}$ being a low discrepancy sequence.
+$\Phi^{-1}(\bm{t}_{i})\sim p(.)$ whilst having a lower variance than direct sampling from $p(.)$.
+MM (Shen et al., 2017; Liu et al., 2022) improves over QMC by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$ on $\Phi^{-1}$.
   • Quadrature-based weight matrix uses quadrature rules with non-uniform deterministic weights.
-The main approach explored is SGQ [27] which uses one-dimensional Gaussian quadrature rule by assuming $\mathcal{K}$ factorizes with respect to the dimensions.
+The main approach explored is SGQ (Dao et al., 2017) which uses one-dimensional Gaussian quadrature rule by assuming $\mathcal{K}$ factorizes with respect to the dimensions.
 Smolyak rule is further implemented to alleviate the curse of dimensionality.
-[7] show that in various kernel benchmark datasets (among combination with TrigRF), QMC and ORF have consistently high performance.
+Liu et al.
+(2022) show that in various kernel benchmark datasets (among combination with TrigRF), QMC and ORF have consistently high performance.
 Due to the significant number of weight matrices in random features literature, we decide to cover only the most prominent ones.
 Appendix A.3 provides the technical details of the approximators discussed here and those that we left out.
-Spectraformer also allows to learn weight matrices by learning $p(.)$ via parameterizing $\boldsymbol{W}$.
-[29] introduce several of these parametrization schemes, which are then adapted into the attention setting by [14].
+Spectraformer also allows to learn weight matrices by learning $p(.)$ via parameterizing $\bm{W}$.
+Yang et al.
+(2015) introduce several of these parametrization schemes, which are then adapted into the attention setting by Chowdhury et al.
+(2022).
 The best performing method is via making the constituent matrices of FastFood learnable, which we term FastFoodL.
@@
 Technical details and other learners we have not covered are explored in Appendix A.4.
-We now explore existing works which have developed from Rahimi and Recht [19] (see Equation 9, 10) and how they lead to the development of Spectraformer.
+We now explore existing works which have developed from Rahimi & Recht (2007) (see Equation 9, 10) and how they lead to the development of Spectraformer.
 3.2 Component functions
 Spectraformer incorporates component functions $f$ such as PosRF, OPRF, SADERF.
-Component functions in Spectraformer combine each weight matrix row $\boldsymbol{\omega}_{i}$ with the input $\boldsymbol{x}$.
+Component functions in Spectraformer combine each weight matrix row $\bm{\omega}_{i}$ with the input $\bm{x}$.
 The base case of the component function is TrigRF.
@@
 Definition 3.2.
-A function of $\boldsymbol{x}$ with respect to $\boldsymbol{\omega}$ is a valid component function if and only if $\mathcal{K}\approx\phi_{\boldsymbol{\omega}}(\boldsymbol{x})^{T}\phi_{\boldsymbol{\omega}}(\boldsymbol{y})$ given that $\boldsymbol{\omega}_{i}\sim p(.)$ in Equation 10.
-It follows that all component functions have provided theoretical guarantees given that $\boldsymbol{W}$ is a random matrix.
-In addition, PosRF, OPRF, and SADERF have also been proven theoretically to approximate $\mathcal{K}_{RBF}$ given $\boldsymbol{W}$ being the ORF.
-While TrigRF[19] is the base component function and guarantees $\mathcal{K}(\boldsymbol{\delta})$ to be close to $\phi_{TrigRF}(\boldsymbol{x})\cdot\phi_{TrigRF}(\boldsymbol{y})$ due to uniform convergence, it is typically evaluated using only the real component of $\exp(-i\boldsymbol{\omega}_{j}^{T}\boldsymbol{x})$.
-This leads to $\cos$ and $\sin$ functions which can produce unstable behavior, especially when $K(\boldsymbol{\delta})$ is close to 0, where variance is infinite [8, 12].
+A function of $\bm{x}$ with respect to $\bm{\omega}$ is a valid component function if and only if $\mathcal{K}\approx\phi_{\bm{\omega}}(\bm{x})^{T}\phi_{\bm{\omega}}(\bm{y})$ given that $\bm{\omega}_{i}\sim p(.)$ in Equation 10.
+It follows that all component functions have provided theoretical guarantees given that $\bm{W}$ is a random matrix.
+In addition, PosRF, OPRF, and SADERF have also been proven theoretically to approximate $\mathcal{K}_{RBF}$ given $\bm{W}$ being the ORF.
+While TrigRF(Rahimi & Recht, 2007) is the base component function and guarantees $\mathcal{K}(\bm{\delta})$ to be close to $\phi_{TrigRF}(\bm{x})\cdot\phi_{TrigRF}(\bm{y})$ due to uniform convergence, it is typically evaluated using only the real component of $\exp(-i\bm{\omega}_{j}^{T}\bm{x})$.
+This leads to $\cos$ and $\sin$ functions which can produce unstable behavior, especially when $K(\bm{\delta})$ is close to 0, where variance is infinite (Choromanski et al., 2021; Likhosherstov et al., 2022).
 Therefore, TrigRF is avoided in Spectraformer.
-  • PosRF [8] is introduced in Spectraformer as a simple modification over TrigRF to counter this instability.
-PosRF guarantees positive function output and its variance approaching 0 as $K(\boldsymbol{\delta})$ approaches 0.
-[12] generalize TrigRF and PosRF to GERF.
-  • OPRF [12] is one of the solution to the reduction of the variance equation of GERF.
+  • PosRF (Choromanski et al., 2021) is introduced in Spectraformer as a simple modification over TrigRF to counter this instability.
+PosRF guarantees positive function output and its variance approaching 0 as $K(\bm{\delta})$ approaches 0.
+Likhosherstov et al.
+(2022) generalize TrigRF and PosRF to GERF.
+  • OPRF (Likhosherstov et al., 2022) is one of the solution to the reduction of the variance equation of GERF.
 OPRF is both positive and bounded (whereas PosRF is positive and unbounded).
-[13] generalize GERF to dense matrix instead of scalar parameters leading to DERF.
-  • SADERF [13] is one of the solutions to the reduction of the variance equation of DERF.
+Likhosherstov et al.
+(2023) generalize GERF to dense matrix instead of scalar parameters leading to DERF.
+  • SADERF (Likhosherstov et al., 2023) is one of the solutions to the reduction of the variance equation of DERF.
 Whilst DERF has many solutions, they all rely on operations which are incompatible with many deep learning libraries on GPU and TPU like SVD and eigen decompositions.
@@
 The alternatives for component functions allow Spectraformer to approximate the RBF with $p(.)$ corresponding to the Gaussian.
-[13] and [12] show that SADERF-ORF is the SOTA combination followed by OPRF-ORF and PosRF-ORF.
+Likhosherstov et al.
+(2023) and Likhosherstov et al.
+(2022) show that SADERF-ORF is the SOTA combination followed by OPRF-ORF and PosRF-ORF.
 Technical details and other component function(s) not covered are explored in Appendix A.5.
@@
 Rows are the component functions, columns are the weight matrices.
-For each cell, the left item is the cited work for this combination for typical kernel tasks, while the right item is the cited work for Transformer-based tasks.
+For each cell, the top item is the cited work for this combination for typical kernel tasks, while the bottom item is the cited work for Transformer-based tasks.
 $\times$ indicates no prior work for the combination.
 Spectraformer explores all combinations listed here, excluding ‘Base’ due to exhaustive work as seen in the table.
-f Monte Carlo Unit-cube Quadrature Learner
-Base ORF SORF QMC MM SGQ FastFoodL
-TrigRF [19], [8] [9], [8] [9], $\times$ [25], $\times$ [26], $\times$ [27], $\times$ [29], [14]
-PosRF [8], [8] [8], [8] $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, [14]
-OPRF [12], [12] [12], [12] $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$
-SADERF [13], [13] [13], [13] $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$
+TrigRF PosRF OPRF SADERF
+Monte Carlo Base (Rahimi & Recht, 2007) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
+(Choromanski et al., 2021) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
+ORF (Yu et al., 2016) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
+(Choromanski et al., 2021) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
+SORF (Yu et al., 2016) $\times$ $\times$ $\times$
+$\times$ $\times$ $\times$ $\times$
+Unit-cube QMC (Avron et al., 2016) $\times$ $\times$ $\times$
+$\times$ $\times$ $\times$ $\times$
+MM (Shen et al., 2017) $\times$ $\times$ $\times$
+$\times$ $\times$ $\times$ $\times$
+Quadrature SGQ (Dao et al., 2017) $\times$ $\times$ $\times$
+$\times$ $\times$ $\times$ $\times$
+Learner FastFoodL (Yang et al., 2015) $\times$ $\times$ $\times$
+(Chowdhury et al., 2022) (Chowdhury et al., 2022) $\times$ $\times$
 3.3 Utility
@@
 Spectraformer highlights this research gap.
-Excluding TrigRF due to its instability and Base due to its underperformance, we conduct experiments on 18 combinations.
+Excluding TrigRF due to its instability and Base due to its under-performance, we conduct experiments on 18 combinations.
 Spectraformer is a generic random feature framework that allows for the combination of any weight matrix with any component function provided that the weight matrix satisfies Definition 3.1 and the component function satisfies Definition 3.2.
@@
 Therefore, we seek to validate its feasibility and show its ability to discover novel and performant combination via experimental results.
+3.4 Adding New Code to Spectraformer
+We now discuss how to add a new component function or weight matrix to the base code for further implementation.
+New component functions need to satisfy Definition 3.2 and new weight matrices need to satisfy Definition 3.1.
+3.4.1 Component Function
+  1. 1.
+    Add the new component function f to src/models/component_functions.py, the arguments should include data (the input), and other optional parameters.
+  2. 2.
+    Import f and add a new entry to FastAttention.comp_functions[f_name] = f in src/models/attention_performer.py (line 176)
+3.4.2 Weight Matrix
+  1. 1.
+    Add the new weight matrix w to src/models/weight_matrix_approx.py or src/models/weight_matrix_learner.py, the arguments should include nb_rows (number of rows), nb_cols (number of columns) and device.
+  2. 2.
+    Import w and add the if clause and $w$ function call in src/models/attention_performer.py (line 208)
 4 Results
 4.1 Experiment Details
+Table 3: Experimental results of Spectraformer variants on the LRA benchmark, over five seeds.
+We report mean accuracy (on the test set), mean time (training time (hour)), and mean memory (peak memory consumption (GB)).
+$L$: ListOps, $T$: Text, $R$: Retrieval, and $\mu$ for the average of the three.
+Entries are ranked by $\mu$ accuracy.
+The previous RF SOTAs (SADERF-ORF (Likhosherstov et al., 2023) and OPRF-ORF (Likhosherstov et al., 2022)) are in bold.
+We note that via Spectraformer, we find various variants outperforming these previous RF SOTAs in terms of accuracy, with different combinations offering different trade-off between accuracy, time, and memory.
+Entries are sorted by mean accuracy.
+Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
+L T R $\mu$ L T R $\mu$ L T R $\mu$
+OPRF-FastFoodL 37.55 (0.48) 64.41 (0.62) 77.70 (0.33) 59.89 1.07 2.07 2.12 1.75 0.86 1.72 1.68 1.42
+OPRF-MM 38.08 (0.53) 60.40 (0.85) 81.09 (0.18) 59.86 0.68 1.26 1.24 1.06 1.36 2.71 2.56 2.21
+PosRF-MM 37.06 (0.37) 61.87 (1.79) 80.58 (0.53) 59.84 0.56 1.05 1.06 0.89 1.17 2.31 2.10 1.86
+OPRF-ORF 38.34 (0.22) 60.16 (0.79) 80.88 (0.17) 59.80 0.68 1.26 1.25 1.06 1.36 2.71 2.56 2.21
+SADERF-QMC 37.37 (0.38) 61.14 (1.48) 80.84 (0.11) 59.78 0.68 1.25 1.29 1.07 1.44 2.86 2.69 2.33
+PosRF-QMC 37.11 (0.09) 61.69 (0.96) 80.55 (0.13) 59.78 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
+SADERF-MM 37.10 (0.22) 60.68 (1.88) 81.13 (0.17) 59.64 0.68 1.25 1.29 1.07 1.44 2.86 2.69 2.33
+SADERF-ORF 37.10 (0.19) 60.39 (2.08) 81.05 (0.22) 59.51 0.68 1.25 1.28 1.07 1.44 2.86 2.69 2.33
+OPRF-QMC 37.69 (0.62) 59.94 (0.59) 80.38 (0.49) 59.34 0.68 1.26 1.26 1.07 1.36 2.71 2.56 2.21
+SADERF-SGQ 37.11 (0.21) 62.46 (0.54) 78.38 (0.25) 59.32 0.68 1.25 1.27 1.07 1.44 2.86 2.69 2.33
+SADERF-FastFoodL 36.02 (1.38) 64.63 (0.18) 76.99 (0.61) 59.21 1.07 2.08 2.16 1.77 0.92 1.84 1.80 1.52
+OPRF-SGQ 37.10 (0.23) 61.25 (0.54) 78.69 (0.54) 59.01 0.67 1.26 1.25 1.06 1.36 2.71 2.56 2.21
+PosRF-ORF 34.35 (5.96) 60.30 (0.97) 80.45 (0.22) 58.37 0.56 1.05 1.06 0.89 1.17 2.31 2.10 1.86
+PosRF-FastFoodL 33.46 (3.70) 64.65 (0.36) 76.95 (0.48) 58.35 1.02 1.98 2.03 1.68 0.79 1.57 1.53 1.30
+SADERF-SORF 33.30 (0.98) 64.70 (0.36) 74.71 (1.90) 57.57 0.68 1.24 1.28 1.07 1.44 2.86 2.69 2.33
+PosRF-SGQ 28.64 (7.54) 62.38 (0.53) 78.28 (0.20) 56.43 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
+OPRF-SORF 27.91 (3.26) 64.76 (0.66) 75.92 (1.74) 56.20 0.67 1.26 1.24 1.06 1.36 2.71 2.56 2.21
+PosRF-SORF 21.27 (6.65) 62.99 (0.40) 67.10 (1.11) 50.45 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
 The LRA covers tasks of different sequence lengths, difficulty, and objective and it is designed to evaluate efficient Transformers, of which Spectraformer belongs to.
-Specifically, we evaluate the models on three LRA tasks: ListOps [30], Byte-Level Text Classification (using the IMDb review dataset) [31], and Byte-Level Document Retrieval (using the ACL Anthology Network dataset) [32].
-All our models are experimented on Google Cloud Platform with n1-standard-4 (4 vCPUs, 15 GB RAM) and 1 NVIDIA T4 GPU.
-Our codebase is based on and includes the adapted or original implementation from [16], [14], [7], [8], [12], [13], [33] (Apache License, Version 2.0), [34] (Apache License, Version 2.0), [35], [36] (Apache License Version 2.0), [17].
-Due to the significant number of models that are trained, we do not run the experiments on different seeds nor perform hyper-parameter tuning.
-The parameters are identical to [16] and chosen to limit the parameter count to account for the marked training time of multiple models.
+Specifically, we evaluate the models on three textual LRA tasks: ListOps (Nangia & Bowman, 2018), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard & Ruder, 2018), and Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013).
+We discuss the task characteristics in greater details in Section 4.2.
+All our models are experimented on NVIDIA V100 and 12 CPU with 20GB of memory.
+Our codebase is based on and includes the adapted or original implementation from Chen et al.
+(2021); Chowdhury et al.
+(2022); Liu et al.
+(2022); Choromanski et al.
+(2021); Likhosherstov et al.
+(2022; 2023); Wang et al.
+(2020); Xiong et al.
+(2021), Kitaev et al.
+(2020); Thomas et al.
+(2018); Zhou et al.
+(2021) (Apache License, Version 2.0).
+We run our experiments on 5 different seeds due to computational limitation.
+The parameters are identical to Chen et al.
+(2021) and chosen to limit the parameter count to account for the marked training time of multiple models.
 All the parameters for Spectraformer variants are kept identical for fair comparison, with the only difference being their random feature component.
-Our code is implemented in Python 3.12 and Pytorch, we use the Transformer base from [16].
+Our code is implemented in Python 3.12 and Pytorch, we use the Transformer base from Chen et al.
+(2021).
 Our hyperparameters are available in Table 8.
 The model name convention is [component function]-[weight matrix] (e.g., OPRF-FastFoodL).
-Table 2: Experimental results of Spectraformer variants on the LRA benchmark.
-We report accuracy (on the test set), time (training time (hour)), and memory (peak memory consumption (GB)).
-$L$: ListOps, $T$: Text, $R$: Retrieval, and $\mu$ for the average of the three.
-The best variant (PosRF-MM) with highest average accuracy, and overall low training time/memory is higlighted in red.
-Existing RF techniques are cited in the model column.
-Novel combinations are marked by $\checkmark$ on the Novel column.
+Our result in Table 3 shows that the top mean accuracy models, ranked in descending order, are: OPRF-FastFoodL, OPRF-MM, PosRF-MM, OPRF-ORF, SADERF-QMC.
+Our mean accuracy result indicates that novel Spectraformer combinations have outperformed the highlighted previous SOTA, i.e., OPRF-ORF and SADERF-ORF.
+However, we are aware that mean accuracy is insufficient in offering a comprehensive picture.
+Therefore, we first proceed with analyzing the statistics of the tasks before analyzing the significance of these results.
+4.2 Task Characteristics
+We first analyze the characteristics of the three textual classification LRA (Tay et al., 2021) tasks: ListOps (Nangia & Bowman, 2018), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard & Ruder, 2018), and Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013).
+We take into consideration three main characteristics:
+  • Number of output classes.
+ListOps is a task which inputs a sequence containing brackets and simple mathematical operators, with the output being a integer between 1 and 10, hence it has 10 classes.
+Text is a sentiment classification task on an IMDb review dataset with the class being either negative or positive, hence it has 2 classes.
+Retrieval inputs two sequences and the output is either False (meaning the sequences are not of the same document) or True (meaning the sequences are of the same document).
+  • Maximum sequence size. ListOps is 2K; Text is 4K; whilst Retrieval requires the processing of two sequences each of 4K length making the total maximum sequence size being 8K.
+  • Required attention span.
+This is the ‘mean distance between the query token and the attended tokens, scaled by attention weights’.
+The average values for required attention span calculated by Tay et al.
+(2021) is: ListOps being 0.8, Text being 0.3, and Retrieval being 1.3.
+Table 4: Mean Pearson correlation between model performance statistics and each task characteristic across weight matrices and component functions.
+Model performance statistics include ‘acc’ (accuracy), ‘time’ (training time), ‘mem’ (peak memory consumption).
+Task characteristics are ‘size’ (maximum sequence size), ‘span’ (required attention span), and ‘classes’ (number of output classes).
+Correlation is calculated for each weight matrix or component function and the average value over weight matrix or component function respectively displayed.
+Left is weight matrix $W$, right is component function $f$.
+$W$ acc time mem
+size 0.93 0.74 0.64
+span 0.37 0.02 -0.10
+classes -0.91 -0.96 -0.94
+$f$ acc time mem
+size 0.93 0.56 0.60
+span 0.36 0.02 -0.10
+classes -0.91 -0.72 -0.89
+We then calculate Pearson correlation of accuracy, training time, and memory consumption against these characteristics.
+Our result is shown in Table 4.
+The table indicates that: (1) contrary to intuition, longer sequence size gives higher accuracy but this is not because of increasing required attention span and they somewhat increase training time and peak memory consumption; (2) increasing the number of classes decreases accuracy, training time, and peak memory consumption.
+4.3 Analysis of Combinations
+Table 5: Mean statistics across weight matrices for each task.
+Task is denoted on the top left cell of each table.
+Mean statistics include ‘acc’ (accuracy (%)), ‘time’ (training time (hour)), and ‘mem’ (peak memory consumption (GB)).
+ListOps acc time mem
+SORF 27.49 0.64 1.32
+SGQ 34.28 0.64 1.32
+FastFoodL 35.68 1.05 0.86
+ORF 36.59 0.64 1.32
+QMC 37.39 0.64 1.32
+MM 37.41 0.64 1.32
+Text acc time mem
+ORF 60.28 1.19 2.63
+QMC 60.92 1.19 2.63
+MM 60.98 1.19 2.63
+SGQ 62.03 1.19 2.63
+SORF 64.15 1.18 2.63
+FastFoodL 64.56 2.04 1.71
+Retrieval acc time mem
+SORF 72.58 1.19 2.45
+FastFoodL 77.22 2.10 1.67
+SGQ 78.45 1.19 2.45
+QMC 80.59 1.20 2.45
+ORF 80.79 1.20 2.45
+MM 80.93 1.20 2.45
+We now explore the suitability of weight matrices and component functions depending on task characteristics.
+Table 5 shows the mean statistics across weight matrices for each task.
+We notice that weight matrices form two different groups which align with the required attention span differences.
+ListOps and Retrieval are tasks with high required attention span, ListOps being 0.8 and Retrieval being 1.3.
+In both these tasks, we observe that ORF, QMC, and MM perform comparatively against each other and a lot better against SORF, SGQ, and FastFoodL.
+However, the Text task, which has a much lower required attention span of 0.3, has much higher performance from the SORF-SGQ-FastFoodL group, whilst ORF, QMC, and MM perform much poorly on this task.
+Based on this observation, we can conclude that ORF, QMC, and MM are suited for high required attention span; whilst SORF, SGQ, and FastFoodL are suited for low required attention span.
+Analyzing mean statistics per task across component functions do not show any task-specific trend (refer to Table 9 for more information).
+Table 6: Summary statistics of component functions $f$ (top) and weight matrices $W$ (bottom) in mean accuracy (on the test set), mean time (training time (hour)), and mean memory (peak memory consumption (GB)).
+$f$ and $W$ entries are sorted by mean accuracy respectively.
 Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
-Novel Model L T R $\mu$ L T R $\mu$ L T R $\mu$
-PosRF-ORF [8] 37.35 61.60 80.53 59.83 1.30 2.84 2.89 2.34 1.17 2.31 2.10 1.86
-$\checkmark$ PosRF-SORF 22.98 63.31 65.52 50.60 1.29 2.81 2.83 2.31 1.17 2.31 2.10 1.86
-$\checkmark$ PosRF-QMC 37.50 60.41 80.71 59.54 1.30 2.84 2.89 2.34 1.17 2.31 2.10 1.86
-$\checkmark$ PosRF-SGQ 37.45 62.68 78.37 59.50 1.30 2.83 2.89 2.34 1.17 2.31 2.10 1.86
-$\checkmark$ PosRF-MM 38.05 61.85 80.67 60.19 1.31 2.84 2.89 2.35 1.17 2.31 2.10 1.86
-PosRF-FastFoodL [14] 24.95 64.76 76.37 55.36 2.69 5.59 5.61 4.63 0.78 1.56 1.53 1.29
-OPRF-ORF [12] 37.50 59.35 80.90 59.25 1.61 3.45 3.50 2.86 1.36 2.71 2.56 2.21
-$\checkmark$ OPRF-SORF 32.11 64.34 77.47 57.97 1.57 3.37 3.41 2.78 1.36 2.71 2.56 2.21
-$\checkmark$ OPRF-QMC 38.41 60.32 80.80 59.84 1.61 3.46 3.51 2.86 1.36 2.71 2.56 2.21
-$\checkmark$ OPRF-MM 38.71 60.39 80.45 59.85 1.61 3.46 3.51 2.86 1.36 2.71 2.56 2.21
-$\checkmark$ OPRF-SGQ 22.53 61.34 79.29 54.39 1.60 3.45 3.47 2.84 1.36 2.71 2.56 2.21
-$\checkmark$ OPRF-FastFoodL 37.40 64.04 78.32 59.92 2.81 5.84 5.86 4.84 0.85 1.69 1.67 1.40
-SADERF-ORF [13] 37.35 61.37 80.87 59.86 1.61 3.52 3.58 2.90 1.44 2.86 2.69 2.33
-$\checkmark$ SADERF-SORF 32.31 64.49 76.43 57.74 1.58 3.44 3.49 2.84 1.44 2.86 2.69 2.33
-$\checkmark$ SADERF-QMC 37.70 59.89 80.65 59.41 1.61 3.52 3.58 2.90 1.44 2.86 2.69 2.33
-$\checkmark$ SADERF-MM 38.00 60.80 80.48 59.76 1.61 3.52 3.58 2.90 1.44 2.86 2.69 2.33
-$\checkmark$ SADERF-SGQ 36.79 63.55 77.31 59.22 1.61 3.52 3.57 2.90 1.44 2.86 2.69 2.33
-$\checkmark$ SADERF-FastFoodL 28.63 64.64 77.61 56.96 2.81 5.89 5.92 4.87 0.92 1.83 1.79 1.51
-4.2 Comparison of Weight Matrix-Component Function Alternatives using Spectraformer
-Our Spectraformer comparison results are available in Table 2.
-We now analyze the performance of component functions and weight matrices.
-We perform our analysis on both general trends of component functions across weight matrices and vice versa.
-We also make use of mean statistics (i.e., mean accuracy, mean training time, and mean peak memory consumption) which we give in brackets in the subsequent discussion.
-Figure 3: Peak memory consumption (x-axis), accuracy (y-axis) on three textual LRA [15] tasks: ListOps, Text, Retrieval.
-Color indicates training time, the darker the faster, the brighter the slower (see legend).
-Models compared are the best Spectraformer variant (PosRF-MM) (boldened), the OT, and other efficient Transformers (RF and non-RF) in the literature.
-Values for non-RF Transformers are taken from [16].
-Component functions.
-In terms of accuracy, SADERF (58.82%) typically performs better than OPRF (58.53%) and OPRF performs better than PosRF (57.50%) by small margin, agreeing with [12, 13].
-However, SADERF (3.21 hours, 2.19 GB) and OPRF (3.17 hours, 2.24 GB) consistently take longer to train and have higher memory consumption than PosRF (2.71%, 1.76 GB) across weight matrices.
-This is likely due to OPRF and SADERF having additional computation of the L2 norm of the weight matrix which PosRF does not calculate.
-Weight matrices.
-In terms of accuracy, SGQ (59.70%), ORF (59.64%), and QMC (59.59%) have the best performance.
-This is followed by MM (57.93%), FastFoodL (57.41%), and the worst being SORF (55.43%).
-In terms of training time, most weight matrices are rather uniform: SORF (2.64 hours) and MM, SGQ, QMC, ORF all at 2.69 hours.
-FastFoodL takes significant amount of time (4.78 hours).
-For peak memory consumption, FastFoodL is the most efficient at 1.73 GB whilst all other methods maintain a uniform value of 2.13 GB.
-FastFoodL reduces the space complexity at the cost of the time complexity.
-SORF reduces training time whilst compromising to a small extent its accuracy.
-This is interesting as it violates our initial assumptions that learners should perform better than approximators.
-Best model.
-Overall, we find that PosRF-MM has the best performance, with the top accuracy (60.19%), top-3 training time (2.35 hours) and top-3 peak memory consumption (1.86 GB), outperforming the SOTA SADERF-ORF [13] (59.86%, 2.90 hours, and 2.33 GB) and OPRF-ORF [12] (59.25%, 2.86 hours, 2.21 GB) in the existing RF Transformer family.
-PosRF-MM achieves this performance due to the combination of both an efficient component function (PosRF) and performative weight matrix (MM).
-PosRF-MM has higher average accuracy (60.19) than the previous SOTA, SADERF-ORF (59.86), whilst having 23.4% faster training time, and 25.2% lower memory consumption.
-Our result demonstrates the viability of Spectraformer as a unifying random feature framework of Transformer.
-Table 3: Comparison of the best Spectraformer variant (PosRF-MM), highlighted in red, against the OT and other efficient Transformers (RF and non-RF) in the literature.
-We report three metrics: accuracy (on the test set), time (training time (hour)), and memory (peak memory consumption (GB)).
-$L$ stands for ListOps, $T$ for Text, $R$ for Retrieval, and $\mu$ for the average of the three.
-Best results for each section and task are boldened.
-Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
-Model L T R $\mu$ L T R $\mu$ L T R $\mu$
-Original Transformer [1] 38.37 61.95 80.69 60.34 2.24 4.30 8.33 4.96 5.37 10.37 10.77 8.84
-Nyströmformer [17] 38.51 64.83 80.52 61.29 0.71 0.71 1.29 0.90 1.37 1.21 2.39 1.66
-Linformer [35] 37.45 58.93 78.19 58.19 0.60 0.65 1.13 0.79 0.99 0.99 1.89 1.29
-Informer [36] 32.53 62.64 77.57 57.58 1.19 1.60 2.91 1.90 4.85 5.12 5.77 5.25
-Reformer [33] 37.68 62.93 78.99 59.87 0.85 0.94 1.73 1.17 1.61 1.61 2.98 2.07
-BigBird [37] 39.25 63.86 80.28 61.13 1.88 2.00 3.81 2.56 2.71 2.83 4.97 3.50
-Skyformer [16] 38.69 64.70 82.06 61.82 1.29 1.02 1.86 1.39 1.75 1.59 3.15 2.16
-PosRF-ORF [8] 37.35 61.60 80.53 59.83 1.30 2.84 2.89 2.34 1.17 2.31 2.10 1.86
-OPRF-ORF [12] 37.50 59.35 80.90 59.25 1.61 3.45 3.50 2.86 1.36 2.71 2.56 2.21
-SADERF-ORF [13] 37.35 61.37 80.87 59.86 0.00 3.52 3.58 2.36 0.00 2.86 2.69 1.85
-PosRF-FastFoodL [14] 24.95 64.76 76.37 55.36 2.69 5.59 5.61 4.63 0.78 1.56 1.53 1.29
-Spectraformer (PosRF-MM) 38.05 61.85 80.67 60.19 1.31 2.84 2.89 2.35 1.17 2.31 2.10 1.86
-4.3 Comparison with other efficient Transformers
-When comparing with other non-random feature Transformers, we observe that PosRF-MM performs comparatively against the OT whilst reducing memory consumption and training time significantly (see Figure 3 and Table 3).
-On ListOps and Retrieval, PosRF-MM and other Spectraformer variants perform on par with other non-RF efficient Transformers.
-Results for non-RF efficient Transformers (above the horizontal divider) in Table 3 are taken from [16].
-Spectraformer does not outperform more performant Nyström-based transformers which are not based on random features.
-Having said that, we are the first to have conducted a comprehensive analysis of random feature methods in the Transformer via the Spectraformer framework, comparing it against other past works and the OT.
-Spectraformer can be extended in the future to include a larger range of component function and weight matrix options.
-We discuss limitations, and broader impact of our research in Appendix A.1.
+$f$ SADERF 59.17 1.19 2.19
+$f$ OPRF 59.02 1.18 2.08
+$f$ PosRF 57.20 1.02 1.77
+$W$ MM 59.78 1.01 2.13
+$W$ QMC 59.63 1.01 2.13
+$W$ ORF 59.22 1.01 2.13
+$W$ FastFoodL 59.15 1.73 1.41
+$W$ SGQ 58.25 1.01 2.13
+$W$ SORF 54.74 1.00 2.13
+In analyzing efficiency, we follow the definition of efficient Transformers in Tay et al.
+(2022) in terms of memory and computation, the latter interpreted as computational cost.
+The overall trend of accuracy, time, and memory of component functions is that they increase linearly from PosRF, OPRF, to SADERF (see Table 6).
+Hence, when limited computational resource is a concern, PosRF is a better option.
+Training time and peak memory consumption relies more heavily on component functions than weight matrices, with the exception of FastFoodL.
+FastFoodL takes significantly less memory consumption at the expense of training time.
+The accuracy trend for weight matrices is as described above.
+Table 6 also shows that whilst FastFoodL generally does not work as well due to its low required attention span, its high performance in the task of Text which is low in required attention span and the anomalously performant OPRF-FastFoodL leads to a data skew in the final ranking in Table 3.
+Comparing between combinations, we find that for ListOps, OPRF-ORF is the most accurate (38.34%), followed by OPRF-MM (38.08%) and OPRF-QMC (37.69%).
+OPRF-ORF, being the previous SOTA, is undoubtedly the best choice here since other models maintain the same training time and peak memory consumption with lower accuracy.
+For Text, the most accurate model is OPRF-SORF (64.76%), followed by SADERF-SORF (64.70%) and PosRF-FastFoodL (64.65%).
+These all outperform the previous SOTAs (OPRF-ORF and SADERF-ORF) by $4\%+$ whilst maintaining a diversity of training time and peak memory consumption options.
+Both OPRF-SORF and SADERF-SORF maintain the same level of training time and peak memory consumption as the SOTAs, PosRF-FastFoodL offers much less peak memory consumption at the expense of longer training time.
+For Retrieval, the most accurate model is SADERF-MM (81.13%), followed by OPRF-MM (81.09%) and SADERF-ORF (81.05%).
+Whilst these models only perform slightly better than the SOTA SADERF-ORF, OPRF-MM offers less peak memory consumption compared to SADERF-ORF.
 5 Conclusion
@@
 We experiment with 18 combinations, of which 14 are novel.
-Novel combinations outperform the previous SOTA in the family for multiple tasks.
-Our best model, PosRF-MM, has higher average accuracy (60.19) than the previous SOTA, SADERF-ORF (59.86), whilst having 23.4% faster training time, and 25.2% lower memory consumption.
-This shows the viability of Spectraformer as a unifying random feature framework of Transformer.
+Novel combinations perform competitively against, if not more performant than the previous SOTA in the family for LRA textual tasks.
+We show empirically that the choice of performant weight matrix depends on the required attention span of the task.
+Our results and the extensibility of our work shows the viability of Spectraformer as a unifying random feature framework of Transformer.
 Appendix A Appendix
@@
 Due to the significant number of experimented models, we are unable to perform hyper-parameter fine-tuning.
-Future work can benefit from extending the scope of the current research to other tasks, i.e., other non-textual LRA tasks (image classification, path-finder, and path-x) and short-context datasets/ tasks such as GLUE [38].
+Future work can benefit from extending the scope of the current research to other tasks, i.e., other non-textual LRA tasks (image classification, path-finder, and path-x) and short-context datasets/ tasks such as GLUE (Linzen et al., 2018).
 Transformer-based architecture can also benefit from Spectraformer.
-Past random feature Transformer works [12, 13] have experimented with conformer and vision Transformer with ViT.
-[39] provides a good foundation to extend our work.
+Past random feature Transformer works (Likhosherstov et al., 2022; 2023) have experimented with conformer and vision Transformer with ViT.
+Jeevan & Sethi (2022) provides a good foundation to extend our work.
 Spectraformer, although currently experimented in the Transformer setting, can also extend to non-parametric kernel classification.
-It could also benefit from ablation studies like the comparison variance in [14] or in [40]
-Given the complex nature of language models in the ethical, energy, and societal domains, our work should be used responsibly  [41]
+It could also benefit from ablation studies like the comparison variance in Chowdhury et al.
+(2022) or in Yao et al.
+(2023)
+Given the complex nature of language models in the ethical, energy, and societal domains, our work should be used responsibly (Weidinger et al., 2021).s
 A.2 Related Work
@@
 In addition to the kernelized formulation of attention as shown in Section 2.1, based on our literature survey, we have identified the following three alternative kernel-based formulations of attention.
-Nadaraya-Watson kernel estimator in integration form [42]: Given the Gaussian kernel, where the probability of a single variable and the joint probability of two variables are estimated using kernel density estimation with the Gaussian kernel, then we have the following equation:
+Nadaraya-Watson kernel estimator in integration form (Nguyen et al., 2022): Given the Gaussian kernel, where the probability of a single variable and the joint probability of two variables are estimated using kernel density estimation with the Gaussian kernel, then we have Equation 12, leading to the formulation of FourierFormer.
 […]
-Equation 12 lays the foundation for FourierFormer - a Transformer with attention based on this formulation [42].
-Gaussian decomposition [43]: the attention mechanism can also be decomposed into the Gaussian kernel directly from Equation 4 without appealing to the use of kernel estimator [43].
-$A_{i}=\sum_{j=1}^{N}\frac{1}{Z_{l}(\boldsymbol{q}_{i},\boldsymbol{K})}exp(\frac{-||\boldsymbol{q}_{i}-\boldsymbol{k}_{j}||^{2}_{2}}{2\sqrt{d_{k}}})exp(\frac{||\boldsymbol{q}_{i}||^{2}_{p=2}+||\boldsymbol{k}_{j}||^{2}_{p=2}}{2\sqrt{d_{k}}})\boldsymbol{v}_{j}$ (13)
-Equation 13 leads to Implicit Kernel Attention (IKA) - a Transformer with attention based on this formulation [43].
-Non-local operation [44] is popular decomposition of attention in computer vision.
+Gaussian decomposition (Song et al., 2021): the attention mechanism can also be decomposed into the Gaussian kernel directly from Equation 4 without appealing to the use of kernel estimator (Song et al., 2021).
+Equation 13 leads to Implicit Kernel Attention (IKA).
+$A_{i}=\sum_{j=1}^{N}\frac{1}{Z_{l}(\bm{q}_{i},\bm{K})}exp(\frac{-||\bm{q}_{i}-\bm{k}_{j}||^{2}_{2}}{2\sqrt{d_{k}}})exp(\frac{||\bm{q}_{i}||^{2}_{p=2}+||\bm{k}_{j}||^{2}_{p=2}}{2\sqrt{d_{k}}})\bm{v}_{j}$ (13)
+Non-local operation (Wang et al., 2018) is popular decomposition of attention in computer vision.
 This is the basis for Vision Transformer and related work.
-Non-local operation itself a generalisation of the non-local means used in image denoising [45].
+Non-local operation itself a generalisation of the non-local means used in image denoising (Buades et al., 2005).
 Non-local operation is essentially an operation that operates on the entire input range, instead of a specific window which is characteristic of convolutional operations (see Equation 14).
-Indeed, the attention vector output $\boldsymbol{y}$ can be defined using $f$ which is the softmax (see Equation 15) in the non-local operation.
-$\mathbf{\boldsymbol{y}}_{i}=\frac{1}{C(\boldsymbol{x})}\sum_{\forall j}f(\boldsymbol{x}_{i},\boldsymbol{x}_{j})g(\boldsymbol{x}_{j})$ (14)
-$\mathbf{\boldsymbol{y}}=softmax(\boldsymbol{x}^{T}\boldsymbol{W}_{\theta}^{T}\boldsymbol{W}_{\theta}\boldsymbol{x})g(\boldsymbol{x})$ (15)
+Indeed, the attention vector output $\bm{y}$ can be defined using $f$ which is the softmax (see Equation 15) in the non-local operation.
+$\mathbf{\bm{y}}_{i}=\frac{1}{C(\bm{x})}\sum_{\forall j}f(\bm{x}_{i},\bm{x}_{j})g(\bm{x}_{j})$ (14)
+$\mathbf{\bm{y}}=softmax(\bm{x}^{T}\bm{W}_{\theta}^{T}\bm{W}_{\theta}\bm{x})g(\bm{x})$ (15)
 We will show that the non-local operation in Equation 14 is a generalisation of the Nadaraya-Watson kernel estimator.
-Specifically, when defining $C(\boldsymbol{x})=\sum_{l}\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{l}),f=\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{l}),g=\boldsymbol{v}_{j}$, then we obtain the original Nadaraya-Watson kernel estimator in Equation 4.
+Specifically, when defining $C(\bm{x})=\sum_{l}\mathcal{K}(\bm{q}_{i},\bm{k}_{l}),f=\mathcal{K}(\bm{q}_{i},\bm{k}_{l}),g=\bm{v}_{j}$, then we obtain the original Nadaraya-Watson kernel estimator in Equation 4.
 The variety of kernel decompositions of attention demonstrate the robustness of the kernel interpretation of attention and hence opening up an alley in attention improvement through the lens of kernel.
@@
 However, there has been other kernel approximation techniques outside of random features.
-These include greedy basis selection [46], divide and conquer [47, 48], and nyström methods [49] (which led to the development of Skyformer [16] and Nyströmformer [17]).
-Figure 4: Spectraformer weight matrices. Red tick denotes techniques which have been explored in the literature. Green tick denotes techniques which are explored in this work.
+These include greedy basis selection (Smola & Schökopf, 2000), divide and conquer (Hsieh et al., 2014; Zhang et al., 2013), and nyström methods (Williams & Seeger, 2000) (which led to the development of Skyformer (Chen et al., 2021) and Nyströmformer (Xiong et al., 2021)).
+Figure 2: Spectraformer weight matrices.
+Red tick denotes techniques which have been explored in the literature.
+Green tick denotes techniques which are explored in this work.
+Please refer to Appendix A.3 for references.
+inline]MAYBE: Table here like table 2 comparison of different kernel space and time complexities combined with table 3 comparison of convergence rates and required random features for kernel approximation error in (Liu et al., 2022)
 A.3 Technical Details of Weight Approximators
@@
 We have not covered SSR in our experiments.
+inline]update the equations so that we use the same dimension for the number of components and so on inline]Generative should be DGM, add Base to the graph
 We now explore these weight matrices in the order specified above.
-Base [19] is the direct sampling from a distribution $\boldsymbol{P}_{ij}\sim p$, corresponding to the kernel $\mathcal{K}$ without any additional adjustment, and in approximating the RBF, $P=G$ where $G$ is the normal distribution.
-$W_{Base}=\frac{1}{\sigma}\boldsymbol{P}$ (16)
-FastFoodF [28] makes use of Hadamard and diagonal matrices to speed up Gaussian matrices $G$ constructions for RFF in $O(n\log{d})$ (100 times faster) with $O(n)$ space (1000 times less space) given $n\geq d$ against Base.
+Base (Rahimi & Recht, 2007) is the direct sampling from a distribution $\bm{P}_{ij}\sim p$, corresponding to the kernel $\mathcal{K}$ without any additional adjustment, and in approximating the RBF, $P=G$ where $G$ is the normal distribution.
+$W_{Base}=\frac{1}{\sigma}\bm{P}$ (16)
+FastFoodF (Le et al., 2013) makes use of Hadamard and diagonal matrices to speed up Gaussian matrices $G$ constructions for RFF in $O(n\log{d})$ (100 times faster) with $O(n)$ space (1000 times less space) given $n\geq d$ against Base.
 However, FastFoodF increases variance and approximation error and decreases the concentration bound.
-$W_{FastFood_{L}}=\frac{1}{\sigma}\boldsymbol{SHG\Pi HB}$ (17)
+$W_{FastFood_{L}}=\frac{1}{\sigma}\bm{SHG\Pi HB}$ (17)
 […]
-SCRF (Signed Circulant Matrix Projection) [50] maintains a high matrix construction speed and $O(s)$ space ($s=t*d$) whilst maintaining the variance of Base, an improvement over FastFoodF.
+SCRF (Signed Circulant Matrix Projection) (Feng et al., 2015) maintains a high matrix construction speed and $O(s)$ space ($s=t*d$) whilst maintaining the variance of Base, an improvement over FastFoodF.
 It does this by making use of circulant matrices.
-$\boldsymbol{W}_{SCRF}=[\boldsymbol{P}^{(1)},\boldsymbol{P}^{(2)},...,\boldsymbol{P}^{(t)}],t=d|D$ (18)
-where $\boldsymbol{P}^{(i)}$ is a signed circulant Gaussian matrix, with the first column of circulant random matrix $\boldsymbol{C}^{(i)_{[d]}}$ of $\boldsymbol{P}^{(i)}$ drawn randomly from a Gaussian distribution $\mathcal{N}(\boldsymbol{0},\boldsymbol{I}/\sigma^{2})$.
+$\bm{W}_{SCRF}=[\bm{P}^{(1)},\bm{P}^{(2)},...,\bm{P}^{(t)}],t=d|D$ (18)
+where $\bm{P}^{(i)}$ is a signed circulant Gaussian matrix, with the first column of circulant random matrix $\bm{C}^{(i)_{[d]}}$ of $\bm{P}^{(i)}$ drawn randomly from a Gaussian distribution $\mathcal{N}(\bm{0},\bm{I}/\sigma^{2})$.
 The circulant matrix via Discrete Fourier Transform can be defined as:
-$\begin{array}[]{cc}\boldsymbol{C}=\frac{1}{m}\boldsymbol{F}^{*}diag(\boldsymbol{Fc})\boldsymbol{F}\\ \boldsymbol{F}=[e^{i\frac{2\pi}{m}kn}]^{m-1}_{k,n=0}\end{array}$ (19)
-P-Model [51] generalizes over both circulant matrices (including SCRF), FastFoodF, along with other Gaussian and semi-Gaussian structured matrix methods.
-$\boldsymbol{W}_{P_{model}}=[\boldsymbol{g^{T}P_{1}},\boldsymbol{g^{T}P_{2}},...,\boldsymbol{g^{T}P_{s}}]^{T}$ (20)
-where $\boldsymbol{g}\in\mathbb{R}^{a}$ is a Gaussian vector and a sequence of $m$ $\boldsymbol{P}_{i}\in\mathbb{R}^{a*d}$ matrices with L2 norm columns, where $P_{i}$ represents a P-model.
+$\begin{array}[]{cc}\bm{C}=\frac{1}{m}\bm{F}^{*}diag(\bm{Fc})\bm{F}\\ \bm{F}=[e^{i\frac{2\pi}{m}kn}]^{m-1}_{k,n=0}\end{array}$ (19)
+P-Model (Choromanski & Sindhwani, 2016) generalizes over both circulant matrices (including SCRF), FastFoodF, along with other Gaussian and semi-Gaussian structured matrix methods.
+$\bm{W}_{P_{model}}=[\bm{g^{T}P_{1}},\bm{g^{T}P_{2}},...,\bm{g^{T}P_{s}}]^{T}$ (20)
+where $\bm{g}\in\mathbb{R}^{a}$ is a Gaussian vector and a sequence of $m$ $\bm{P}_{i}\in\mathbb{R}^{a*d}$ matrices with L2 norm columns, where $P_{i}$ represents a P-model.
 Semi-Gaussian structured matrices including Toeplitz, Hankel, and other Toeplitz-like matrices are an open area of research for random matrices.
-ORF (Orthogonal RF) [9] decreases the approximation error significantly compared to Base.
+ORF (Orthogonal RF) (Yu et al., 2016) decreases the approximation error significantly compared to Base.
 This is done via replacing $G$ with a properly scaled random orthogonal matrix.
 However, generating orthogonal matrices become costly quickly as the number of dimensions increases.
-$\boldsymbol{W}_{ORF}=\frac{1}{\sigma}\boldsymbol{SQ}$ (21)
-where $\boldsymbol{Q}$ is a uniformly distributed random orthogonal matrix (on the Stiefel manifold) obtained from the QR decomposition of $\boldsymbol{G}$, the set of rows of $\boldsymbol{Q}$ forming a basis in $\mathbb{R}^{d}$, and $\boldsymbol{S}$ is a diagonal matrix with entries sampled i.i.d from the $\chi$-distribution with $d$ degrees of freedom, thus making the rows of $\boldsymbol{SQ}$ and $\boldsymbol{G}$ identically distributed.
-SORF (Structured ORF) [9] decreases time and space complexity of ORF (from $O(d^{2})$ to $O(d\log{d})$ with almost no extra memory cost) by imposing structure on the orthogonal matrices, inspired by structural methods.
+$\bm{W}_{ORF}=\frac{1}{\sigma}\bm{SQ}$ (21)
+where $\bm{Q}$ is a uniformly distributed random orthogonal matrix (on the Stiefel manifold) obtained from the QR decomposition of $\bm{G}$, the set of rows of $\bm{Q}$ forming a basis in $\mathbb{R}^{d}$, and $\bm{S}$ is a diagonal matrix with entries sampled i.i.d from the $\chi$-distribution with $d$ degrees of freedom, thus making the rows of $\bm{SQ}$ and $\bm{G}$ identically distributed.
+SORF (Structured ORF) (Yu et al., 2016) decreases time and space complexity of ORF (from $O(d^{2})$ to $O(d\log{d})$ with almost no extra memory cost) by imposing structure on the orthogonal matrices, inspired by structural methods.
 SORF is unbiased with large $d$.
-We replace $\boldsymbol{S}$ with $\sqrt{d}$ and $\boldsymbol{Q}$ with structured matrix $\boldsymbol{HD}_{1}\boldsymbol{HD}_{2}\boldsymbol{HD}_{3}$
-$\boldsymbol{W}_{SORF}=\frac{\sqrt{d}}{\sigma}\boldsymbol{HD}_{1}\boldsymbol{HD}_{2}\boldsymbol{HD}_{3}$ (22)
-where $\boldsymbol{D}_{i}\in\mathbb{R}^{d*d},i=1,2,3$ is diagonal ’sign-flipping‘ matrices with diagonal entries sampled from the Rademacher distribution, $\boldsymbol{H}$ is the normalized Walsh-Hadamard matrix.
-ROM (Random Orthogonal Embeddings) [52] generalizes SORF ($t=3$) to:
-$\boldsymbol{W}_{ROM}=\frac{\sqrt{d}}{\sigma}\prod_{i=1}^{t}\boldsymbol{SD}_{i}$ (23)
-where $\boldsymbol{S}$ is a class of $L_{2}$-normalized versions of Kronecker product matrices, of which the Hadamard $\boldsymbol{H}$ is a representative of, and $\boldsymbol{D}$ being independent diagonal matrices.
-When $\boldsymbol{S}=\boldsymbol{H}$ and $t=3$, we obtain SORF used in our experiments.
+We replace $\bm{S}$ with $\sqrt{d}$ and $\bm{Q}$ with structured matrix $\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$
+$\bm{W}_{SORF}=\frac{\sqrt{d}}{\sigma}\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$ (22)
+where $\bm{D}_{i}\in\mathbb{R}^{d*d},i=1,2,3$ is diagonal ’sign-flipping‘ matrices with diagonal entries sampled from the Rademacher distribution, $\bm{H}$ is the normalized Walsh-Hadamard matrix.
+ROM (Random Orthogonal Embeddings) (Choromanski et al., 2017) generalizes SORF ($t=3$) to:
+$\bm{W}_{ROM}=\frac{\sqrt{d}}{\sigma}\prod_{i=1}^{t}\bm{SD}_{i}$ (23)
+where $\bm{S}$ is a class of $L_{2}$-normalized versions of Kronecker product matrices, of which the Hadamard $\bm{H}$ is a representative of, and $\bm{D}$ being independent diagonal matrices.
+When $\bm{S}=\bm{H}$ and $t=3$, we obtain SORF used in our experiments.
 Another ROM variant is S-Hybrid, which we do not experiment with, given as
-$\boldsymbol{W}_{S_{Hybrid}}=\boldsymbol{SD}_{k}^{(\mathcal{U})}\prod_{i=1}^{k-1}\boldsymbol{SD}_{i}^{(\mathcal{R})}$ (24)
-where $\boldsymbol{D}_{k}^{(\mathcal{U})}$ is a diagonal matrix with i.i.d.
+$\bm{W}_{S_{Hybrid}}=\bm{SD}_{k}^{(\mathcal{U})}\prod_{i=1}^{k-1}\bm{SD}_{i}^{(\mathcal{R})}$ (24)
+where $\bm{D}_{k}^{(\mathcal{U})}$ is a diagonal matrix with i.i.d.
 $Unif(\mathcal{S}^{1})$ on the unit circle of $\mathbb{C}$.
-When being used in the component function, instead of calculating $\phi(\boldsymbol{x})^{T}\phi(\boldsymbol{y})$ in Equation 10.
-We model the equation as $Real(\overline{\phi(\boldsymbol{x})}^{T}\phi(\boldsymbol{y}))$.
-SimRF (Simplex RF) [11] is developed to be a positive RF with the most optimal solution, the MSE of which is the lowest in the geometric methods.
+When being used in the component function, instead of calculating $\phi(\bm{x})^{T}\phi(\bm{y})$ in Equation 10.
+We model the equation as $Real(\overline{\phi(\bm{x})}^{T}\phi(\bm{y}))$.
+SimRF (Simplex RF) (Reid et al., 2023) is developed to be a positive RF with the most optimal solution, the MSE of which is the lowest in the geometric methods.
 This is an alternative to the ROM-based geometric methods.
@@
 SimRF is defined as:
-$\boldsymbol{W}_{simp}=\boldsymbol{DSR}$ (25)
-where $\boldsymbol{D}\in\mathbb{R}^{d*d}=diag(\boldsymbol{\omega}_{i}),\boldsymbol{R}\in\mathbb{R}^{d*d}$ being a random orthogonal matrix drawn from Haar measure on $O(d)$ and $\boldsymbol{S}\in\mathbb{R}^{d*d}$ with row $\boldsymbol{s}_{i}$:
-$\boldsymbol{s}_{i}=\begin{cases}\sqrt{\frac{d}{d-1}}\boldsymbol{e}_{i}-\frac{\sqrt{d}+1}{(d-1)^{3/2}}(1,...,1,0)^{T}&1\leq i<d\\ \frac{1}{\sqrt{d-1}}(1,1,...,1,0)^{T}&i=d\end{cases}$ (26)
+$\bm{W}_{simp}=\bm{DSR}$ (25)
+where $\bm{D}\in\mathbb{R}^{d*d}=diag(\bm{\omega}_{i}),\bm{R}\in\mathbb{R}^{d*d}$ being a random orthogonal matrix drawn from Haar measure on $O(d)$ and $\bm{S}\in\mathbb{R}^{d*d}$ with row $\bm{s}_{i}$:
+$\bm{s}_{i}=\begin{cases}\sqrt{\frac{d}{d-1}}\bm{e}_{i}-\frac{\sqrt{d}+1}{(d-1)^{3/2}}(1,...,1,0)^{T}&1\leq i<d\\ \frac{1}{\sqrt{d-1}}(1,1,...,1,0)^{T}&i=d\end{cases}$ (26)
 being unit vectors which are manifestly normalized and subtend obtuse angles.
-SimRF variants are defined based on how $\boldsymbol{\omega}_{i}$ is constructed.
-For SimRFindep, $\boldsymbol{\omega}_{i}\sim\chi_{d}$.
-For Simdep, we permit the random vector direction $\{\hat{\boldsymbol{\omega}}_{i}\}$ to be correlated with norms $\{\boldsymbol{\omega}_{i}\}$:
-$\boldsymbol{\omega}_{i}=-\frac{\sum_{j\neq i}\omega_{j}}{||\sum_{j\neq i}\omega_{j}||_{2}}\boldsymbol{\omega}_{i},i=1,...,d$ (27)
-QMC (Quasi-Monte Carlo) [25] evaluates on a low discrepancy sequence (e.g., Halton, Sobol’ Faure, and Niederreiter) of points instead of random points in Monte Carlo.
+SimRF variants are defined based on how $\bm{\omega}_{i}$ is constructed.
+For SimRFindep, $\bm{\omega}_{i}\sim\chi_{d}$.
+For Simdep, we permit the random vector direction $\{\hat{\bm{\omega}}_{i}\}$ to be correlated with norms $\{\bm{\omega}_{i}\}$:
+$\bm{\omega}_{i}=-\frac{\sum_{j\neq i}\omega_{j}}{||\sum_{j\neq i}\omega_{j}||_{2}}\bm{\omega}_{i},i=1,...,d$ (27)
+QMC (Quasi-Monte Carlo) (Avron et al., 2016) evaluates on a low discrepancy sequence (e.g., Halton, Sobol’ Faure, and Niederreiter) of points instead of random points in Monte Carlo.
 Although the approximation error is only reduced minimally, QMC has been shown to perform better than MC in high dimensions and does not have undesirable clustering effect.
-To calculate QMC, we first assume that $p(\boldsymbol{x})=\prod^{d}_{j=1}p_{j}(\boldsymbol{x}_{j})$ factorizes with respect to the dimensions with $p_{j}(.)$ being a univariate density function.
+To calculate QMC, we first assume that $p(\bm{x})=\prod^{d}_{j=1}p_{j}(\bm{x}_{j})$ factorizes with respect to the dimensions with $p_{j}(.)$ being a univariate density function.
 Then we define:
-$\Phi^{-1}(\boldsymbol{t})=\left(\Phi_{1}^{-1}\left(\boldsymbol{t}_{1}\right),\ldots,\Phi_{d}^{-1}\left(\boldsymbol{t}_{d}\right)\right)\in\mathbb{R}^{d}$ (28)
-where $\Phi_{j}$ being the cumulative distribution function of $p_{j}$, $\boldsymbol{t}_{1},\boldsymbol{t}_{2},...,\boldsymbol{t}_{s}\in[0,1]^{d}$ being a low discrepancy sequence, and $\boldsymbol{\omega}_{i}=\Phi^{-1}(\boldsymbol{t}_{i})$.
+$\Phi^{-1}(\bm{t})=\left(\Phi_{1}^{-1}\left(\bm{t}_{1}\right),\ldots,\Phi_{d}^{-1}\left(\bm{t}_{d}\right)\right)\in\mathbb{R}^{d}$ (28)
+where $\Phi_{j}$ being the cumulative distribution function of $p_{j}$, $\bm{t}_{1},\bm{t}_{2},...,\bm{t}_{s}\in[0,1]^{d}$ being a low discrepancy sequence, and $\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$.
 We can thus transform the integral on $\mathbb{R}^{d}$ in Equation 9 to an integral on the unit cube $[0,1]^{d}$ as
-$\mathcal{K}(\boldsymbol{x}-\boldsymbol{x}^{\prime})=\int_{[0,1]^{d}}\exp\big{(}\mbox{i}(\boldsymbol{x}-\boldsymbol{x}^{\prime})^{\!\top}\Phi^{-1}(\boldsymbol{t})\big{)}\mbox{d}\boldsymbol{t}$
+$\mathcal{K}(\bm{x}-\bm{x}^{\prime})=\int_{[0,1]^{d}}\exp\big{(}\mbox{i}(\bm{x}-\bm{x}^{\prime})^{\!\top}\Phi^{-1}(\bm{t})\big{)}\mbox{d}\bm{t}$
 The weight matrix then can be defined as:
-$\boldsymbol{W}_{\text{QMC}}=[\Phi^{-1}(\boldsymbol{t}_{1}),\Phi^{-1}(\boldsymbol{t}_{2}),\ldots,\Phi^{-1}(\boldsymbol{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}\;.$
-QMC can be further improved by a sub-grouped based rank-one lattice construction which improved complexity [53].
-MM (Moment Matching) [26] [7] improves over QMC by removing the undesirable clustering effect and having the same approximation error with less features.
+$\bm{W}_{\text{QMC}}=[\Phi^{-1}(\bm{t}_{1}),\Phi^{-1}(\bm{t}_{2}),\ldots,\Phi^{-1}(\bm{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}\;.$
+QMC can be further improved by a sub-grouped based rank-one lattice construction which improved complexity (Lyu et al., 2020).
+MM (Moment Matching) (Shen et al., 2017; Liu et al., 2022) improves over QMC by removing the undesirable clustering effect and having the same approximation error with less features.
 This is done by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$:
-$\boldsymbol{W}_{\text{MM}}=[\widetilde{\Phi}^{-1}(\boldsymbol{t}_{1}),\widetilde{\Phi}^{-1}(\boldsymbol{t}_{2}),\ldots,\widetilde{\Phi}^{-1}(\boldsymbol{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}$
-where $\widetilde{\Phi}^{-1}(\boldsymbol{t}_{i})=\tilde{\boldsymbol{A}}^{-1}(\Phi^{-1}(\boldsymbol{t}_{i})-\tilde{\boldsymbol{\mu}})$ can be constructed using moment matching with sample mean $\tilde{\boldsymbol{\mu}}=\frac{1}{s}\sum_{i=1}^{s}\Phi^{-1}(\boldsymbol{t}_{i})$ and the square root of the sample covariance matrix $\tilde{\boldsymbol{A}}$ satisfying $\tilde{\boldsymbol{A}}\tilde{\boldsymbol{A}}^{\!\top}=\text{Cov}(\Phi^{-1}(\boldsymbol{t}_{i})-\tilde{\boldsymbol{\mu}})$.
-SSF (Spherical Structured Features) [54] improves over QMC by including rotation-invariant kernel as well as in terms of complexity.
-Rotation-invariant property suggests the construction of feature maps using spherical equal weight approximation using Riesz s-energy on a d-dimensional sphere $\mathbb{S}^{d}:=\{\boldsymbol{x}\in\mathbb{R}^{d+1}|||\boldsymbol{x}||_{2}=1\}$.
-Specifically, we construct $\{\boldsymbol{v}_{i}\}^{s}_{i=1}$ asymptotically uniformly distributed on the sphere and obtain $\boldsymbol{V}:=[\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{s}]\in\mathbb{S}^{(d\times s}$, that $\Phi^{-1}(t)$ uses the one-dimensional QMC point, we have the weight matrix:
-$\boldsymbol{W}_{\text{SSF}}=[\Phi^{-1}(t)\boldsymbol{v}_{1},\Phi^{-1}(t)\boldsymbol{v}_{2},\ldots,\Phi^{-1}(t)\boldsymbol{v}_{s}]^{\!\top}\in\mathbb{R}^{s\times(d+1)},$
-SGQ (Sparse Grid Quadrature) [27] evolves from GQ.
-GQ (Gaussian Quadrature) [27] assumes that the kernel $k$ factorizes with respect to the dimensions and thus can be approximated by a one-dimensional Gaussian quadrature rule.
+$\bm{W}_{\text{MM}}=[\widetilde{\Phi}^{-1}(\bm{t}_{1}),\widetilde{\Phi}^{-1}(\bm{t}_{2}),\ldots,\widetilde{\Phi}^{-1}(\bm{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}$
+where $\widetilde{\Phi}^{-1}(\bm{t}_{i})=\tilde{\bm{A}}^{-1}(\Phi^{-1}(\bm{t}_{i})-\tilde{\bm{\mu}})$ can be constructed using moment matching with sample mean $\tilde{\bm{\mu}}=\frac{1}{s}\sum_{i=1}^{s}\Phi^{-1}(\bm{t}_{i})$ and the square root of the sample covariance matrix $\tilde{\bm{A}}$ satisfying $\tilde{\bm{A}}\tilde{\bm{A}}^{\!\top}=\text{Cov}(\Phi^{-1}(\bm{t}_{i})-\tilde{\bm{\mu}})$.
+SSF (Spherical Structured Features) (Lyu, 2017) improves over QMC by including rotation-invariant kernel as well as in terms of complexity.
+Rotation-invariant property suggests the construction of feature maps using spherical equal weight approximation using Riesz s-energy on a d-dimensional sphere $\mathbb{S}^{d}:=\{\bm{x}\in\mathbb{R}^{d+1}|||\bm{x}||_{2}=1\}$.
+Specifically, we construct $\{\bm{v}_{i}\}^{s}_{i=1}$ asymptotically uniformly distributed on the sphere and obtain $\bm{V}:=[\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{s}]\in\mathbb{S}^{(d\times s}$, that $\Phi^{-1}(t)$ uses the one-dimensional QMC point, we have the weight matrix:
+$\bm{W}_{\text{SSF}}=[\Phi^{-1}(t)\bm{v}_{1},\Phi^{-1}(t)\bm{v}_{2},\ldots,\Phi^{-1}(t)\bm{v}_{s}]^{\!\top}\in\mathbb{R}^{s\times(d+1)},$
+SGQ (Sparse Grid Quadrature) (Dao et al., 2017) evolves from GQ.
+GQ (Gaussian Quadrature) (Dao et al., 2017) assumes that the kernel $k$ factorizes with respect to the dimensions and thus can be approximated by a one-dimensional Gaussian quadrature rule.
 However the total number of points $s$ scale exponentially with the dimensions.
@@
 Assuming the third-degree SGQ using symmetric univariate quadrature points $\left\{-\hat{p}_{1},0,\hat{p}_{1}\right\}$ with weights $(\hat{a}_{1},\hat{a}_{0},\hat{a}_{1})$, then we have:
-$\boldsymbol{W}_{\text{SGQ}}\!=\![\boldsymbol{0}_{d},\hat{p}_{1}\boldsymbol{e}_{1},\ldots,\hat{p}_{1}\boldsymbol{e}_{d},-\hat{p}_{1}\boldsymbol{e}_{1},\ldots,-\hat{p}_{1}\boldsymbol{e}_{d}]^{\!\top}\!\in\!\mathbb{R}^{(2d+1)\times d}\!$
+$\bm{W}_{\text{SGQ}}\!=\![\bm{0}_{d},\hat{p}_{1}\bm{e}_{1},\ldots,\hat{p}_{1}\bm{e}_{d},-\hat{p}_{1}\bm{e}_{1},\ldots,-\hat{p}_{1}\bm{e}_{d}]^{\!\top}\!\in\!\mathbb{R}^{(2d+1)\times d}\!$
 given that $\mathbb{e}_{i}$ is the d-dimensional standard basis vector with the ith element being 1.
-SSR (Spherical-Radial Rules) [55] transforms Equation 10 into a double integral over a hyper-sphere and the real line. This leads to the following approximation:
-$\boldsymbol{W}_{\text{SSR}}=\boldsymbol{\vartheta}\otimes\begin{bmatrix}(\boldsymbol{QV})^{\top}\\ -(\boldsymbol{Q}\boldsymbol{V})^{\top}\end{bmatrix}\in\mathbb{R}^{(2d+1)\times d},$
-where $\boldsymbol{\vartheta}=[\vartheta_{1},\vartheta_{2},\ldots,\vartheta_{s}]$, $\boldsymbol{V}=[\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{d+1}]$, $\vartheta\sim\chi(d+2)$ and $\{\boldsymbol{v}_{i}\}_{i=1}^{d+1}$ being vertices of a unit regular d-simplex randomly rotated by $\boldsymbol{Q}$ (a random orthogonal matrix).
+SSR (Spherical-Radial Rules) (Munkhoeva et al., 2018) transforms Equation 10 into a double integral over a hyper-sphere and the real line. This leads to the following approximation:
+$\bm{W}_{\text{SSR}}=\bm{\vartheta}\otimes\begin{bmatrix}(\bm{QV})^{\top}\\ -(\bm{Q}\bm{V})^{\top}\end{bmatrix}\in\mathbb{R}^{(2d+1)\times d},$
+where $\bm{\vartheta}=[\vartheta_{1},\vartheta_{2},\ldots,\vartheta_{s}]$, $\bm{V}=[\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{d+1}]$, $\vartheta\sim\chi(d+2)$ and $\{\bm{v}_{i}\}_{i=1}^{d+1}$ being vertices of a unit regular d-simplex randomly rotated by $\bm{Q}$ (a random orthogonal matrix).
 A.4 Technical Details of Weight Learners
@@
 Weight matrix learner comes from a long line of work in kernel methods, namely, kernel learning via random features.
-Typically, a component function $f$ is chosen and a weight matrix $\boldsymbol{W}$ is produced via kernel approximation, providing us with the feature map $\phi$.
-$\phi$ can then be combined with input $\boldsymbol{x}$ and weight $\boldsymbol{v}$ to fit the learning objective.
-Kernel learning, however, imposes an additional objective: learning the weight matrix $\boldsymbol{W}$.
-The approach to the dual objectives separates the kernel learning methods into two-stage and one-stage methods [7].
-Two-stage methods (e.g., [56], [23], [57]) solves the dual objective separately: $\boldsymbol{W}$ is typically first learned via a kernel alignment scheme [58], then we solve for $\boldsymbol{v}$.
-One-stage methods (e.g., [14], [22]), on the other hand, solves the dual objective in parallel: $p(.)$ corresponding to $\boldsymbol{W}$ is parameterized then $\boldsymbol{v}$ is solved typically.
-It is not plausible to apply kernel alignment [58] on Transformer architectures.
+Typically, a component function $f$ is chosen and a weight matrix $\bm{W}$ is produced via kernel approximation, providing us with the feature map $\phi$.
+$\phi$ can then be combined with input $\bm{x}$ and weight $\bm{v}$ to fit the learning objective.
+Kernel learning, however, imposes an additional objective: learning the weight matrix $\bm{W}$.
+The approach to the dual objectives separates the kernel learning methods into two-stage and one-stage methods (Liu et al., 2022).
+Two-stage methods (e.g., Yu et al.
+(2015); Wilson & Adams (2013); Bullins et al.
+(2018)) solves the dual objective separately: $\bm{W}$ is typically first learned via a kernel alignment scheme (Wang et al., 2015), then we solve for $\bm{v}$.
+One-stage methods (e.g., Chowdhury et al.
+(2022); Tompkins et al.
+(2019)), on the other hand, solves the dual objective in parallel: $p(.)$ corresponding to $\bm{W}$ is parameterized then $\bm{v}$ is solved typically.
+It is not plausible to apply kernel alignment (Wang et al., 2015) on Transformer architectures.
 Therefore, only one-stage methods can be considered in the context of weight matrix learner.
-Two kernel learning methods, GMM and FastFoodL, have been experimented in the non-Transformer setting [29], then adapted into the Transformer setting [14].
-DGM (Deep Generative Model) has also been effective in modeling distributions [59].
-It too has been adapted by [14] for the Transformer setting.
-Due to the superior performance of FastFoodL in [14], we have only experimented with this technique.
+Two kernel learning methods, GMM and FastFoodL, have been experimented in the non-Transformer setting (Yang et al., 2015), then adapted into the Transformer setting (Chowdhury et al., 2022).
+DGM (Deep Generative Model) has also been effective in modeling distributions (Kingma & Welling, 2013).
+It too has been adapted by Chowdhury et al.
+(2022) for the Transformer setting.
+Due to the superior performance of FastFoodL with this technique, we have only experimented with this technique.
 We now explore them.
-FastFoodL (see Equation 17) is proposed to be learnable by [29] with two variations: FSARD (scaling matrix $\boldsymbol{S}$ previously sampled from chi-squared is made learnable) and FSGBARD (optimizes the marginal likelihood with respect to the diagonal matrices $\boldsymbol{G}$ and $\boldsymbol{B}$).
-It is further adapted by [14] to make $\boldsymbol{S}$ or $\boldsymbol{S},\boldsymbol{G},\boldsymbol{B}$ learnable parameters.
-In our experiments, we maintain the set up by [14].
-GMM (Gaussian Mixture Model) [23, 24] is a known universal approximator of probability distribution.
-Specifically, we can consider $\boldsymbol{W}_{GMM}$ being composed of $s$ components, with each component $c$ (with $C$ total components) being a Gaussian with mean vector $\mu_{c}$ and covariance matrix $\Sigma_{c}$ and component weight $\pi_{c}$.
-$p(\boldsymbol{\omega})$ can then be approximated as $p(\boldsymbol{\omega})=\sum_{c=1}^{C}\pi_{c}\mathcal{N}(\mu_{c},\Sigma_{c})$ [14].
-Assuming $\pi_{c}=\frac{1}{C}$, the weight matrix with learnable weights $\Sigma$ and $\mu$ can be parameterized as [14]:
-$\boldsymbol{W}_{GMM}=[\Sigma_{1}n+\mu_{1},...,\Sigma_{C}n+\mu_{C}]^{T},n\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ (29)
-Another formulation of GMM is proposed in [29] in combination with FastFood, albeit it extends beyond constructing the weight matrix.
-DGM (Deep Generative Model) [59] allows for a generic distribution modeling using neural network.
-DGM is initially proposed to make differentiating sampling processes possible via the reparametrization trick, where the generator $g$ creates samples from the target distribution $\boldsymbol{\omega}_{m}=g(\boldsymbol{n}_{m})$ via transforming from a noise distribution $\boldsymbol{n}_{m}\sim p_{0}(.)$.
+FastFoodL (see Equation 17) is proposed to be learnable by Yang et al.
+(2015) with two variations: FSARD (scaling matrix $\bm{S}$ previously sampled from chi-squared is made learnable) and FSGBARD (optimizes the marginal likelihood with respect to the diagonal matrices $\bm{G}$ and $\bm{B}$).
+It is further adapted by Chowdhury et al.
+(2022) to make $\bm{S}$ or $\bm{S},\bm{G},\bm{B}$ learnable parameters.
+In our experiments, we maintain this set up.
+GMM (Gaussian Mixture Model) (Wilson & Adams, 2013; Oliva et al., 2016) is a known universal approximator of probability distribution.
+Specifically, we can consider $\bm{W}_{GMM}$ being composed of $s$ components, with each component $c$ (with $C$ total components) being a Gaussian with mean vector $\mu_{c}$ and covariance matrix $\Sigma_{c}$ and component weight $\pi_{c}$.
+$p(\bm{\omega})$ can then be approximated as $p(\bm{\omega})=\sum_{c=1}^{C}\pi_{c}\mathcal{N}(\mu_{c},\Sigma_{c})$ (Chowdhury et al., 2022).
+Assuming $\pi_{c}=\frac{1}{C}$, the weight matrix with learnable weights $\Sigma$ and $\mu$ can be parameterized as:
+$\bm{W}_{GMM}=[\Sigma_{1}n+\mu_{1},...,\Sigma_{C}n+\mu_{C}]^{T},n\sim\mathcal{N}(\bm{0},\bm{I})$ (29)
+Another formulation of GMM is proposed in Yang et al. (2015) in combination with FastFood, albeit it extends beyond constructing the weight matrix.
+DGM (Deep Generative Model) (Kingma & Welling, 2013) allows for a generic distribution modeling using neural network.
+DGM is initially proposed to make differentiating sampling processes possible via the reparametrization trick, where the generator $g$ creates samples from the target distribution $\bm{\omega}_{m}=g(\bm{n}_{m})$ via transforming from a noise distribution $\bm{n}_{m}\sim p_{0}(.)$.
 In addition to the techniques discussed above, there are also some other prominent one-stage methods (e.g., INN, BBQ).
-INN (Invertible Neural Network) [60] is another universal approximator of probability distribution, consisting of invertible operations transforming samples from a known distribution to a different more complex distribution.
-BBQ (Black Box Quantile) [22] follows the QMC scheme and models $\Phi^{-1}$ using parameterized quantile function.
+INN (Invertible Neural Network) (Xu et al., 2021) is another universal approximator of probability distribution, consisting of invertible operations transforming samples from a known distribution to a different more complex distribution.
+BBQ (Black Box Quantile) (Tompkins et al., 2019) follows the QMC scheme and models $\Phi^{-1}$ using parameterized quantile function.
 A.5 Technical Details of Component Functions
-Figure 5: Spectraformer component functions. Red tick denotes techniques which have been explored in the literature. Green tick denotes techniques which are explored in this work.
+Figure 3: Spectraformer component functions.
+Red tick denotes techniques which have been explored in the literature.
+Green tick denotes techniques which are explored in this work.
+Please refer to Appendix A.5 for references.
 Component functions discussed in Section 3.2 all belong to CDERF.
@@
 We now investigate the theoretical details of these component functions, the CDERF family is introduced in their order of discovery.
-CDERF (Complex dense exponential random feature) is given in Equation 30, where $\boldsymbol{\omega},\boldsymbol{x}\in\mathbb{R}^{d}$, $p=\mathcal{N}(0,1)^{d}$, $k\in\{1,2\}$ and $\mathbb{S}_{d_{c}}$ being a set of $d*d$ complex symmetric matrices.
+CDERF (Complex dense exponential random feature) is given in Equation 30, where $\bm{\omega},\bm{x}\in\mathbb{R}^{d}$, $p=\mathcal{N}(0,1)^{d}$, $k\in\{1,2\}$ and $\mathbb{S}_{d_{c}}$ being a set of $d*d$ complex symmetric matrices.
 The parameters and constraints of different CDERF functions are specified in Table 7:
-$f_{DE}^{(k)}(\boldsymbol{\omega},\boldsymbol{x})=Real(D\exp(\boldsymbol{\omega}^{T}\boldsymbol{A}\boldsymbol{\omega}+\boldsymbol{\omega}^{T}\boldsymbol{B}^{(k)}\boldsymbol{x}+\boldsymbol{x}^{T}\boldsymbol{C}^{(k)}\boldsymbol{x}))$ (30)
+$f_{DE}^{(k)}(\bm{\omega},\bm{x})=Real(D\exp(\bm{\omega}^{T}\bm{A}\bm{\omega}+\bm{\omega}^{T}\bm{B}^{(k)}\bm{x}+\bm{x}^{T}\bm{C}^{(k)}\bm{x}))$ (30)
 Table 7: CDERF component functions
-$\boldsymbol{A}$ $\boldsymbol{B}^{(1)}$ $\boldsymbol{B}^{(2)}$ $\boldsymbol{C}^{(1)}$ $\boldsymbol{C}^{(2)}$ $D$ $\boldsymbol{x}$ $\boldsymbol{y}$
+$\bm{A}$ $\bm{B}^{(1)}$ $\bm{B}^{(2)}$ $\bm{C}^{(1)}$ $\bm{C}^{(2)}$ $D$ $\bm{x}$ $\bm{y}$
 […]
 […]
-SADERF [13] $A_{GE}\boldsymbol{I}_{d}$ $B_{GE}\boldsymbol{I}_{d}$ $C_{GE}\boldsymbol{I}_{d}$ $D_{GE}$ $\Psi\boldsymbol{x}$ $\Psi^{-1}\boldsymbol{y}$
-GERF [12] $A_{GE}\boldsymbol{I}_{d}$ $B_{GE}^{(1)}\boldsymbol{I}_{d}$ $B_{GE}^{(2)}\boldsymbol{I}_{d}$ $C_{GE}\boldsymbol{I}_{d}$ $D_{GE}$ $\boldsymbol{x}$ $\boldsymbol{y}$
-TrigRF [19] $0\boldsymbol{I}_{d}$ $i\boldsymbol{I}_{d}$ $-i\boldsymbol{I}_{d}$ $0\boldsymbol{I}_{d}$ $1$ $\boldsymbol{x}$ $\boldsymbol{y}$
-PosRF [8] $0\boldsymbol{I}_{d}$ $1\boldsymbol{I}_{d}$ $-1\boldsymbol{I}_{d}$ $1$ $\boldsymbol{x}$ $\boldsymbol{y}$
-OPRF [12] $0\boldsymbol{I}_{d}$ $1\boldsymbol{I}_{d}$ $-1\boldsymbol{I}_{d}$ $1$ $\boldsymbol{x}$ $\boldsymbol{y}$
-TrigRF (Trigonometric RF) [19] is the base RFF implementation.
-$\begin{array}[]{c}f_{TrigRF}^{(1)}(\boldsymbol{\omega},\boldsymbol{x})=\sqrt{2}\cos(\boldsymbol{\omega}^{T}\boldsymbol{x}+\boldsymbol{b})\\ f_{TrigRF}^{(2)}(\boldsymbol{\omega},\boldsymbol{y})=\sqrt{2}\sin(\boldsymbol{\omega}^{T}\boldsymbol{y}+\boldsymbol{b})\\ \boldsymbol{b}\sim Uniform(0,2\pi)\end{array}$ (31)
+SADERF (Likhosherstov et al., 2023) $A_{GE}\bm{I}_{d}$ $B_{GE}\bm{I}_{d}$ $C_{GE}\bm{I}_{d}$ $D_{GE}$ $\Psi\bm{x}$ $\Psi^{-1}\bm{y}$
+GERF (Likhosherstov et al., 2022) $A_{GE}\bm{I}_{d}$ $B_{GE}^{(1)}\bm{I}_{d}$ $B_{GE}^{(2)}\bm{I}_{d}$ $C_{GE}\bm{I}_{d}$ $D_{GE}$ $\bm{x}$ $\bm{y}$
+TrigRF (Rahimi & Recht, 2007) $0\bm{I}_{d}$ $i\bm{I}_{d}$ $-i\bm{I}_{d}$ $0\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
+PosRF (Choromanski et al., 2021) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $-1\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
+OPRF (Likhosherstov et al., 2022) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $-1\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
+TrigRF (Trigonometric RF) (Rahimi & Recht, 2007) is the base RFF implementation.
+$\begin{array}[]{c}f_{TrigRF}^{(1)}(\bm{\omega},\bm{x})=\sqrt{2}\cos(\bm{\omega}^{T}\bm{x}+\bm{b})\\ f_{TrigRF}^{(2)}(\bm{\omega},\bm{y})=\sqrt{2}\sin(\bm{\omega}^{T}\bm{y}+\bm{b})\\ \bm{b}\sim Uniform(0,2\pi)\end{array}$ (31)
 The use of the trigonometric sine and cosine functions leads to unstable behavior when the inputs have negative dimension-values.
 This can be further exacerbated when the values TrigRF try to approximate are close to 0 (since most values are of low significance).
-This causes the variance to approach infinity [8, 12].
+This causes the variance to approach infinity (Choromanski et al., 2021; Likhosherstov et al., 2022).
 Therefore, we do not want to use TrigRF to perform kernel approximation in attention.
-PosRF (Positive RF) [8] fixes the problem of TrigRF by enforcing positive component function output in the softmax.
+PosRF (Positive RF) (Choromanski et al., 2021) fixes the problem of TrigRF by enforcing positive component function output in the softmax.
 The variance of PosRF (in contrast to the variance of TrigRF approaching infinity) approaches 0 as the approximated value of the Softmax kernel approaches 0.
 PosRF has two forms: Positive RF Base (PosRF-B) (see Equation 32) and Positive RF Hyperbolic (PosRF-Hyp), which is multi-component, i.e., Equation 11: $l=2$, (see Equation 33).
-$\begin{array}[]{c}f_{PosRF_{B}}(\boldsymbol{\omega},\boldsymbol{x})=\exp(\boldsymbol{\omega}^{T}\boldsymbol{x}-\frac{||\boldsymbol{x}||^{2}}{2})\end{array}$ (32)
-$\begin{array}[]{ll}f_{1,PosRF_{Hyp}}(\boldsymbol{\omega},\boldsymbol{x})=\exp(\boldsymbol{\omega}^{T}\boldsymbol{x}-||\boldsymbol{x}||^{2})\\ f_{2,PosRF_{Hyp}}(\boldsymbol{\omega},\boldsymbol{x})=\exp(-\boldsymbol{\omega}^{T}\boldsymbol{x}-||\boldsymbol{x}||^{2})\end{array}$ (33)
+$\begin{array}[]{c}f_{PosRF_{B}}(\bm{\omega},\bm{x})=\exp(\bm{\omega}^{T}\bm{x}-\frac{||\bm{x}||^{2}}{2})\end{array}$ (32)
+$\begin{array}[]{ll}f_{1,PosRF_{Hyp}}(\bm{\omega},\bm{x})=\exp(\bm{\omega}^{T}\bm{x}-||\bm{x}||^{2})\\ f_{2,PosRF_{Hyp}}(\bm{\omega},\bm{x})=\exp(-\bm{\omega}^{T}\bm{x}-||\bm{x}||^{2})\end{array}$ (33)
 We only use PosRF-B in our experiments.
-GERF (Generalized exponential RF) [12] generalizes both TrigRF and PosRF with Equation 34.
-$\begin{array}[]{c}f_{GERF}^{(1)}(\boldsymbol{\omega},\boldsymbol{x})=D\exp(A||\boldsymbol{\omega}||^{2}+B\boldsymbol{\omega}^{T}\boldsymbol{x}+\boldsymbol{C}||\boldsymbol{x}||^{2})\\ f_{GERF}^{(2)}(\boldsymbol{\omega},\boldsymbol{y})=D\exp(\boldsymbol{A}||\boldsymbol{\omega}||^{2}+sB\boldsymbol{\omega}^{T}\boldsymbol{y}+\boldsymbol{C}||\boldsymbol{y}||^{2})\\ \\ Re(1-4A)>0,\boldsymbol{B}=\sqrt{s(1-4\boldsymbol{A})}\\ \boldsymbol{C}=-(s+1)/2,D=(\sqrt[4]{1-4\boldsymbol{A}})^{d}\\ \boldsymbol{A}\in\mathbb{C},s\in\{-1,+1\}\end{array}$ (34)
+GERF (Generalized exponential RF) (Likhosherstov et al., 2022) generalizes both TrigRF and PosRF with Equation 34.
+$\begin{array}[]{c}f_{GERF}^{(1)}(\bm{\omega},\bm{x})=D\exp(A||\bm{\omega}||^{2}+B\bm{\omega}^{T}\bm{x}+\bm{C}||\bm{x}||^{2})\\ f_{GERF}^{(2)}(\bm{\omega},\bm{y})=D\exp(\bm{A}||\bm{\omega}||^{2}+sB\bm{\omega}^{T}\bm{y}+\bm{C}||\bm{y}||^{2})\\ \\ Re(1-4A)>0,\bm{B}=\sqrt{s(1-4\bm{A})}\\ \bm{C}=-(s+1)/2,D=(\sqrt[4]{1-4\bm{A}})^{d}\\ \bm{A}\in\mathbb{C},s\in\{-1,+1\}\end{array}$ (34)
 where $\sqrt{.}$ and .$n]{.}$ denoting a principal root with a complex argument
-OPRF (Optimized positive RF) [12] is the solution to the minimization of the variance of GERF.
-Specifically it is defined as Equation 34 with $s=+1$, $||\boldsymbol{x}+\boldsymbol{y}||^{2}>0$, and $\boldsymbol{A}\in\mathbb{R}$ defined in terms of $p^{*}$ as:
-$\begin{array}[]{cc}\boldsymbol{A}=(1-1/p^{*})/8\\ p^{*}=(\sqrt{(2||\boldsymbol{x}+\boldsymbol{y}||^{2}+d)^{2}+8d||\boldsymbol{x}+\boldsymbol{y}||^{2}}-2||\boldsymbol{x}+\boldsymbol{y}||^{2}-d)\\ /(4||\boldsymbol{x}+\boldsymbol{y}||^{2})\end{array}$ (35)
-Whilst PosRF is not bounded, OPRF is. OPRF can provide $e^{60}$$\times$ variance reduction in estimating the Softmax compared to TrigRF [12].
-DERF (Dense-exponential random features) [13] extends GERF and replace $\boldsymbol{A},\boldsymbol{B},\boldsymbol{C}$ with dense matrices.
-DERF is CDERF when $\boldsymbol{B},C$ are in the real instead of the complex plane.
+OPRF (Optimized positive RF) (Likhosherstov et al., 2022) is the solution to the minimization of the variance of GERF.
+Specifically it is defined as Equation 34 with $s=+1$, $||\bm{x}+\bm{y}||^{2}>0$, and $\bm{A}\in\mathbb{R}$ defined in terms of $p^{*}$ as:
+$\begin{array}[]{cc}\bm{A}=(1-1/p^{*})/8\\ p^{*}=(\sqrt{(2||\bm{x}+\bm{y}||^{2}+d)^{2}+8d||\bm{x}+\bm{y}||^{2}}-2||\bm{x}+\bm{y}||^{2}-d)\\ /(4||\bm{x}+\bm{y}||^{2})\end{array}$ (35)
+Whilst PosRF is not bounded, OPRF is. OPRF can provide $e^{60}$$\times$ variance reduction in estimating the Softmax compared to TrigRF.
+DERF (Dense-exponential random features) (Likhosherstov et al., 2023) extends GERF and replace $\bm{A},\bm{B},\bm{C}$ with dense matrices.
+DERF is CDERF when $\bm{B},C$ are in the real instead of the complex plane.
 Minimizing the variance of DERF leads to two approaches: ADERF and SDERF.
@@
 Therefore SADERF is proposed.
-SADERF (simplified ADERF) [13] is a special case of ADERF and extends GERF, requiring only basic unary operations in addition.
-$\begin{array}[]{c}f_{SADE}^{(1)}(\boldsymbol{\omega},\boldsymbol{x})=f_{GE}^{(1)}(\boldsymbol{\omega},\Psi\boldsymbol{x}),f_{SADE}^{(2)}(\boldsymbol{\omega},\boldsymbol{y})=f_{GE}^{(2)}(\boldsymbol{\omega},\Psi^{-1}\boldsymbol{y})\\ \Psi_{l,l}^{*}=(\sum_{j}(\boldsymbol{y}_{l}^{(i)})^{2}/\sum_{i}(\boldsymbol{x}_{l}^{(i)})^{2})^{1/4}\end{array}$ (36)
-DIRF (Discretely-induced random features) [12] are based on the assumption of $p(\boldsymbol{\omega})$ being a discrete distribution with $\boldsymbol{\omega}_{1},...,\boldsymbol{\omega}_{d}$ being i.i.d., $\mathbb{P}(\boldsymbol{\omega}_{l}=k)=p_{k}$, $\sum_{k=0}^{\infty}p_{k}=1$, $p_{k}>0,k\in\{0\}\cup\mathbb{N}$.
+SADERF (simplified ADERF) (Likhosherstov et al., 2023) is a special case of ADERF and extends GERF, requiring only basic unary operations in addition.
+$\begin{array}[]{c}f_{SADE}^{(1)}(\bm{\omega},\bm{x})=f_{GE}^{(1)}(\bm{\omega},\Psi\bm{x}),f_{SADE}^{(2)}(\bm{\omega},\bm{y})=f_{GE}^{(2)}(\bm{\omega},\Psi^{-1}\bm{y})\\ \Psi_{l,l}^{*}=(\sum_{j}(\bm{y}_{l}^{(i)})^{2}/\sum_{i}(\bm{x}_{l}^{(i)})^{2})^{1/4}\end{array}$ (36)
+DIRF (Discretely-induced random features) (Likhosherstov et al., 2022) are based on the assumption of $p(\bm{\omega})$ being a discrete distribution with $\bm{\omega}_{1},...,\bm{\omega}_{d}$ being i.i.d., $\mathbb{P}(\bm{\omega}_{l}=k)=p_{k}$, $\sum_{k=0}^{\infty}p_{k}=1$, $p_{k}>0,k\in\{0\}\cup\mathbb{N}$.
 By making use of the Taylor series, we can define DIRF as:
-$f_{DI}(\boldsymbol{\omega},\boldsymbol{x})=\exp(-\frac{||\boldsymbol{x}||^{2}}{2})\prod_{l=1}^{d}\boldsymbol{x}_{i}^{\boldsymbol{\omega}_{l}}(\boldsymbol{\omega}_{l}!)^{\frac{-1}{2}}p_{\boldsymbol{\omega}_{l}}^{\frac{-1}{2}}$ (37)
+$f_{DI}(\bm{\omega},\bm{x})=\exp(-\frac{||\bm{x}||^{2}}{2})\prod_{l=1}^{d}\bm{x}_{i}^{\bm{\omega}_{l}}(\bm{\omega}_{l}!)^{\frac{-1}{2}}p_{\bm{\omega}_{l}}^{\frac{-1}{2}}$ (37)
 DIRF has two variations: PoisRF (Poisson RF) which requires $p(.)$ to be the Poisson distribution and GeomRF (Geometric RF) which requires $p(.)$ to be the geometric distribution.
 Since the many weight matrices introduced in Section 3.1 do not readily approximate these distributions, we decide to leave DIRF out of the paper.
-HRF (Hybrid RF) [10] combines multiple base estimators, similar to multiple kernel learning methods [61], in order to provide the most accurate approximation in regions of interest.
+HRF (Hybrid RF) (Choromanski et al., 2022) combines multiple base estimators, similar to multiple kernel learning methods (Bach et al., 2004), in order to provide the most accurate approximation in regions of interest.
 Choromanski et al.
 on solving softmax approximation using HRF.
-HRF is constructed using a weighted combination $\epsilon$ of $p+1$ base estimators $\hat{SM}^{k}(\boldsymbol{x},\boldsymbol{y})$ and $\lambda$-coefficients, $\Lambda$ $p$ binary estimators $\lambda^{k}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow[0,1]$, both constructed independently:
-$\hat{SM}^{\epsilon,\Lambda}=\sum_{k=1}^{p}\hat{\lambda}^{k}(\boldsymbol{x},\boldsymbol{y})\hat{SM}^{k}(\boldsymbol{x},\boldsymbol{y})+(1-\sum_{k=1}^{p}\hat{\lambda}^{k}(\boldsymbol{x},\boldsymbol{y}))\hat{SM}^{p+1}(\boldsymbol{x},\boldsymbol{y})$ (38)
-Since HRF requires the base estimators $SM(\boldsymbol{x},\boldsymbol{y})$ to be unbiased, the easiest solution is the complex exponential $\hat{SM}^{cexp}_{m}$ which can be derived directly from the expectation of the softmax as follows:
-$\begin{array}[]{cc}\hat{SM}^{cexp}_{m}(\boldsymbol{x},\boldsymbol{y})=\psi^{m}_{A}(\boldsymbol{x})^{T}\psi^{m}_{(\boldsymbol{A}^{T})^{-1}}(\boldsymbol{y})\\ \psi^{m}_{\boldsymbol{M}}(u)=\frac{1}{\sqrt{m}}\exp(-\frac{(\boldsymbol{Mu})^{2}}{2})\\ (\exp(\boldsymbol{\omega}^{T}_{i}\boldsymbol{Mu}),...,\exp(\boldsymbol{\omega}_{m}^{T}\boldsymbol{Mu}))^{T},\boldsymbol{\omega}_{i}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d})\end{array}$ (39)
+HRF is constructed using a weighted combination $\epsilon$ of $p+1$ base estimators $\hat{SM}^{k}(\bm{x},\bm{y})$ and $\lambda$-coefficients, $\Lambda$ $p$ binary estimators $\lambda^{k}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow[0,1]$, both constructed independently:
+$\hat{SM}^{\epsilon,\Lambda}=\sum_{k=1}^{p}\hat{\lambda}^{k}(\bm{x},\bm{y})\hat{SM}^{k}(\bm{x},\bm{y})+(1-\sum_{k=1}^{p}\hat{\lambda}^{k}(\bm{x},\bm{y}))\hat{SM}^{p+1}(\bm{x},\bm{y})$ (38)
+Since HRF requires the base estimators $SM(\bm{x},\bm{y})$ to be unbiased, the easiest solution is the complex exponential $\hat{SM}^{cexp}_{m}$ which can be derived directly from the expectation of the softmax as follows:
+$\begin{array}[]{cc}\hat{SM}^{cexp}_{m}(\bm{x},\bm{y})=\psi^{m}_{A}(\bm{x})^{T}\psi^{m}_{(\bm{A}^{T})^{-1}}(\bm{y})\\ \psi^{m}_{\bm{M}}(u)=\frac{1}{\sqrt{m}}\exp(-\frac{(\bm{Mu})^{2}}{2})\\ (\exp(\bm{\omega}^{T}_{i}\bm{Mu}),...,\exp(\bm{\omega}_{m}^{T}\bm{Mu}))^{T},\bm{\omega}_{i}\sim\mathcal{N}(\bm{0},\bm{I}_{d})\end{array}$ (39)
 To construct $\lambda$-coefficients, three methods are proposed: simultaneous accurate approximation of both small and large values, Gaussian Lambda Coefficients, and adaption to data admitting clustering structure.
 A.6 Experimental Details in Spectraformer
-Table 5: Hyper-parameters for Spectraformer, following the setting from [16] due to limited computational power.
+Table 8: Hyper-parameters for Spectraformer, following the setting from Chen et al.
+(2021) due to limited computational power.
 $L$ stands for ListOps, $T$ for Text, and $R$ for Retrieval.
@@
 Num. features 128 128 128
+Table 9: Mean statistics across component functions for each task.
+Task is denoted on the top left cell of each table.
+Mean statistics include ‘acc’ (accuracy), ‘time’ (training time), and ‘mem’ (peak memory consumption).
+ListOps acc time mem
+PosRF 31.98 0.63 1.10
+OPRF 36.11 0.74 1.28
+SADERF 36.33 0.75 1.35
+Text acc time mem
+OPRF 61.82 1.40 2.54
+PosRF 62.31 1.21 2.19
+SADERF 62.33 1.39 2.69
+Retrieval acc time mem
+PosRF 77.32 1.22 2.01
+SADERF 78.85 1.43 2.54
+OPRF 79.11 1.39 2.41

### v5

@@
 Spectraformer: A Unified Random Feature Framework for Transformer
-Abstract
+Abstract.
 Linearization of attention using various kernel approximation and kernel learning techniques has shown promise.
-Past methods use a subset of combinations of component functions and weight matrices within the random features paradigm.
+Past methods used a subset of combinations of component functions and weight matrices within the random feature paradigm.
 We identify the need for a systematic comparison of different combinations of weight matrices and component functions for attention learning in Transformer.
-In this work, we introduce Spectraformer, a unified framework for approximating and learning the kernel function in linearized attention of the Transformer.
-We experiment with broad classes of component functions and weight matrices for three textual tasks in the LRA benchmark.
-Our empirical findings indicate that different kernels are good at different tasks and that kernel choice is fundamental to performant models.
-Our code is available at: https://anonymous.4open.science/r/spectraformer-8A97.
-1 Introduction
+Hence, we introduce Spectraformer, a unified framework for approximating and learning the kernel function in the attention mechanism of the Transformer.
+Our empirical results demonstrate, for the first time, that a random feature-based approach can achieve performance comparable to top-performing sparse and low-rank methods on the challenging Long Range Arena benchmark.
+Thus, we establish a new state-of-the-art for random feature-based efficient Transformers.
+The framework also produces many variants that offer different advantages in accuracy, training time, and memory consumption.
+Our code is available at: https://github.com/cruiseresearchgroup/spectraformer.
+transformers, kernel, linearized attention, kernelized attention
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMccs: Computing methodologies Kernel methodsccs: Computing methodologies Neural networksccs: Computing methodologies Natural language processing
+1. Introduction
 Transformer (Vaswani et al., 2017) has revolutionized the landscape of natural language processing (NLP) and forms the basis of almost all state-of-the-art language models.
-Its influence has reached beyond NLP into computer vision (Han et al., 2023), speech processing (Lin et al., 2022) and other fields.
+Its influence has reached beyond NLP into computer vision (Han et al., 2023; Yu et al., 2024; Yuan et al., 2025; Hao et al., 2025), speech processing (Lin et al., 2022), and other fields.
 Compared to its precursors like the LSTMs, Transformer leverages parallelism since it is fully based on the attention mechanism.
@@
 Attention in the Original Transformer (referred to as ‘OT’ hereafter) uses the softmax function which is quadratic in time complexity.
-However, softmax is not fundamental to attention.
-The softmax serves the role of a compatibility function, which captures the degree of compatibility between two tokens (see Equations 1 and 2).
+To mitigate the quadratic complexity of the standard attention mechanism, a diverse range of efficient Transformer architectures have been developed (Tay et al., 2022).
+These include memory/ downsampling methods (Big Bird (Zaheer et al., 2020), Nystromformer (Xiong et al., 2021), and Skyformer (Chen et al., 2021)), learnable pattern methods (Reformer (Kitaev et al., 2020)), Informer (Zhou et al., 2021), Linformer (Wang et al., 2020).
+Our work, Spectraformer, focuses on another significant category: the linearization of attention through a unified framework of random features.
+We approach this linearization by rethinking the role of the softmax.
+The softmax acts as a compatibility function, capturing the similarity between pairs of tokens (see Equations 1 and 2).
 Extensive experiments with other compatibility functions show that the softmax is not the only possibility, and that the ‘best’ compatibility function depends on the task (Tsai et al., 2019).
@@
 Tsai et al.
-(2019) show that attention can be formulated using kernels and then be linearized (Katharopoulos et al., 2020), given the feature map of the respective attention kernel.
+(2019) show that attention can be formulated using kernels and then linearized (Katharopoulos et al., 2020), given the feature map of the respective attention kernel.
 Random feature-based algorithms (Liu et al., 2022) are a family of algorithms, inspired by spectral analysis, which provides the feature map associated with a kernel.
@@
 (2021).
-This is seen to lead to three strands of research: (A) Optimization of weight matrices either by making the matrices have lower time or space complexity, or by incorporating beneficial properties like orthogonality to reduce variance (Yu et al., 2016; Choromanski et al., 2021; 2022; Reid et al., 2023); (B) Enhancement of component functions by either: (i) finding a function with a tighter variance or approximation bound, or (ii) be engineered to output with numerical stability and be bounded (Choromanski et al., 2021; Likhosherstov et al., 2022; 2023); (C) Parameterization of weight matrices to perform kernel learning instead of kernel approximation (Chowdhury et al., 2022).
+This is seen to lead to three strands of research: (A) Optimization of weight matrices either (i) by making the matrices have lower time or space complexity, or (ii) by incorporating beneficial properties like orthogonality to reduce variance (Yu et al., 2016; Choromanski et al., 2021, 2022; Reid et al., 2023); (B) Enhancement of component functions by either: (i) finding a function with a tighter variance or approximation bound, or (ii) engineering in order to output with numerical stability and be bounded (Choromanski et al., 2021; Likhosherstov et al., 2022, 2023); (C) Parameterization of weight matrices to perform kernel learning instead of kernel approximation (Chowdhury et al., 2022).
 (A) and (B) seek to improve approximation quality via approximation error and variance reduction.
@@
 Additionally, past works only explore certain combinations of component functions and weight matrices, creating disjoint overlaps and gaps in the literature.
-Finally, the weight matrices are often those explicitly designed in the context of kernelized attention in the Transformer (Yu et al., 2016; Choromanski et al., 2021; 2022; Reid et al., 2023), ignoring those which are popular in the kernel methods literature (e.g., Liu et al.
+Finally, the weight matrices are often those explicitly designed in the context of kernelized attention in the Transformer (Yu et al., 2016; Choromanski et al., 2021, 2022; Reid et al., 2023), ignoring those which are popular in the kernel method literature (e.g., Liu et al.
 (2022)).
-These limitations necessitate a unified framework that allows the comparison of different combinations in search of the best-performing one.
-We introduce Spectraformer, a unified framework of random feature-based attention in Transformer.
+These limitations highlight the need for a unified framework to systematically compare combinations and identify optimal configurations.
+Table 1.
+Experimental results of Spectraformer variants against previous SOTA efficient Transformers on the LRA benchmark.
+All models were run for five seeds using 128 random features.
+We report mean accuracy (on the test set), mean time (training time (hour)), and mean memory (peak memory consumption (GB)).
+$L$: ListOps, $T$: Text, $R$: Retrieval, II: Image, $P$: Pathfinder, and $\mu$ for the average across tasks.
+Entries are sorted in descending order by $\mu$.
+Best models per category are boldened.
+Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
+L T R I P $\mu$ L T R I P $\mu$ L T R I P $\mu$
+Nystromformer 38.69 (0.59) 61.57 (0.38) 80.57 (0.30) 39.49 (0.89) 69.96 (1.27) 58.06 0.56 0.90 1.00 2.19 1.19 1.17 1.10 1.87 1.84 5.55 2.79 2.63
+Big Bird 38.82 (0.52) 61.04 (0.15) 80.53 (0.32) 37.59 (1.04) 68.28 (6.74) 57.25 1.60 3.17 3.24 5.72 2.89 3.32 2.24 4.57 3.80 8.42 4.22 4.65
+OPRF-FastFoodL 37.73 (0.28) 64.32 (0.61) 78.65 (1.94) 38.41 (0.66) 66.76 (0.39) 57.17 1.07 2.07 2.11 4.02 2.05 2.26 0.84 1.68 1.64 3.26 1.68 1.82
+Reformer 35.50 (4.09) 61.28 (0.45) 78.31 (0.29) 43.62 (0.81) 66.28 (2.35) 57.00 0.65 1.29 1.29 2.51 1.26 1.40 1.61 3.20 2.95 6.38 3.20 3.47
+Skyformer 38.96 (0.63) 60.67 (0.45) 81.90 (0.37) 32.93 (0.39) 69.81 (1.38) 56.85 0.78 1.35 1.49 3.09 1.63 1.67 1.67 3.01 2.92 7.90 3.96 3.89
+OT 38.49 (0.80) 59.72 (1.07) 80.86 (0.20) 37.48 (0.83) 67.56 (6.12) 56.82 1.96 7.27 7.07 4.30 2.20 4.56 4.37 16.72 8.74 9.42 4.72 8.79
+OPRF-SGQ 37.08 (0.56) 61.09 (0.77) 79.39 (0.94) 36.68 (0.28) 67.53 (2.79) 56.35 0.68 1.27 1.25 2.46 1.29 1.39 1.33 2.64 2.50 5.25 2.64 2.87
+PosRF-$M$ 37.13 (0.27) 62.88 (1.17) 80.58 (0.44) 33.82 (0.51) 67.11 (0.45) 56.30 0.56 1.05 1.06 2.02 1.05 1.15 1.14 2.26 2.05 4.49 2.26 2.44
+OPRF-OR 38.05 (0.78) 60.18 (0.49) 81.19 (0.23) 33.70 (0.57) 67.24 (0.88) 56.07 0.68 1.26 1.24 2.47 1.29 1.39 1.33 2.64 2.50 5.25 2.64 2.87
+SADERF-QMC 37.18 (0.47) 60.61 (2.06) 80.67 (0.44) 34.55 (0.63) 67.32 (0.40) 56.07 0.68 1.24 1.28 2.47 1.26 1.39 1.40 2.79 2.63 5.56 2.79 3.04
+PosRF-QMC 37.05 (0.16) 60.93 (0.59) 80.67 (0.29) 34.03 (0.94) 67.25 (0.44) 55.99 0.55 1.05 1.05 1.99 1.05 1.14 1.14 2.26 2.05 4.49 2.26 2.44
+OPRF-QMC 37.86 (0.40) 60.87 (1.85) 80.44 (0.20) 33.87 (0.69) 66.81 (0.68) 55.97 0.68 1.26 1.24 2.43 1.28 1.38 1.33 2.64 2.50 5.25 2.64 2.87
+SADERF-ORF 37.41 (0.45) 59.92 (0.90) 81.05 (0.18) 33.06 (1.02) 67.11 (0.56) 55.71 0.69 1.25 1.28 2.50 1.28 1.40 1.40 2.79 2.63 5.56 2.79 3.04
+OPRF-$M$ 38.31 (0.38) 60.01 (1.01) 81.03 (0.31) 33.93 (0.85) 65.17 (5.17) 55.69 0.68 1.26 1.29 2.47 1.28 1.40 1.33 2.64 2.50 5.25 2.64 2.87
+SADERF-$M$ 37.17 (0.27) 60.81 (1.82) 81.05 (0.20) 34.03 (1.04) 65.31 (4.85) 55.67 0.69 1.25 1.28 2.49 1.28 1.40 1.40 2.79 2.63 5.56 2.79 3.04
+SADERF-SGQ 37.22 (0.28) 62.86 (0.96) 78.51 (0.68) 37.82 (0.59) 61.27 (5.38) 55.54 0.68 1.25 1.28 2.50 1.27 1.39 1.40 2.79 2.63 5.56 2.79 3.04
+SADERF-FastFoodL 29.72 (7.27) 64.71 (0.45) 77.50 (0.96) 38.38 (0.82) 66.47 (1.47) 55.35 1.09 2.09 2.18 4.05 2.06 2.29 0.90 1.80 1.76 3.52 1.80 1.96
+PosRF-FastFoodL 29.54 (5.46) 64.61 (0.29) 77.10 (1.02) 38.28 (0.43) 66.38 (0.71) 55.18 1.02 2.00 2.01 3.88 1.98 2.18 0.77 1.54 1.50 3.01 1.54 1.67
+PosRF-ORF 34.50 (6.49) 61.10 (1.76) 80.53 (0.33) 33.72 (1.03) 65.52 (4.89) 55.07 0.56 1.05 1.06 2.02 1.06 1.15 1.14 2.26 2.05 4.49 2.26 2.44
+OPRF-SORF 29.59 (3.77) 64.81 (0.22) 77.12 (0.78) 37.42 (0.47) 63.98 (5.38) 54.58 0.68 1.26 1.24 2.42 1.28 1.38 1.33 2.64 2.50 5.25 2.64 2.87
+SADERF-SORF 34.31 (2.50) 64.82 (0.26) 75.24 (1.12) 36.76 (1.22) 60.98 (4.90) 54.42 0.68 1.25 1.28 2.47 1.26 1.39 1.40 2.79 2.63 5.56 2.79 3.04
+Linformer 36.94 (0.60) 56.99 (1.23) 77.86 (0.35) 39.09 (0.78) 59.24 (5.84) 54.02 0.41 0.74 0.77 1.28 0.71 0.78 0.86 1.70 1.61 3.39 1.71 1.85
+PosRF-SGQ 28.81 (7.77) 62.30 (0.47) 78.31 (0.39) 37.96 (0.76) 59.71 (7.87) 53.42 0.56 1.06 1.05 2.01 1.06 1.15 1.14 2.26 2.05 4.49 2.26 2.44
+Informer 31.45 (7.20) 62.06 (1.20) 76.92 (0.14) 37.90 (0.30) 57.07 (6.57) 53.08 0.88 1.92 2.23 3.21 1.78 2.01 2.66 5.57 3.27 5.20 2.61 3.86
+PosRF-SORF 21.91 (5.98) 62.06 (1.36) 66.33 (0.41) 28.70 (3.87) 52.37 (5.24) 46.27 0.55 1.06 1.06 2.00 1.05 1.14 1.14 2.26 2.05 4.49 2.26 2.44
+We introduce Spectraformer, a unified framework for approximating and learning the kernel function in the attention mechanism of the Transformer through the lens of random features.
+While other methods (e.g., Nystromformer (Xiong et al., 2021), BigBird (Zaheer et al., 2020)) have explored alternative approaches to improving efficiency, Spectraformer provides a systematic way to explore the vast design space of random feature-based attention.
 It utilizes spectral analysis-inspired random features to construct linearized attention in Transformers.
 Spectraformer allows for the experimentation of various combinations of weight matrices and component functions.
-We will open-source our codebase upon acceptance.
-Through our benchmarking experiments on three textual Long Range Arena (Tay et al., 2021) tasks and 18 combinations of weight matrices and component functions, Spectraformer enables novel combinations of component functions and weight matrices from the kernel literature to be experimented easily in the Transformer setting.
-This process makes possible the discovery of combinations that outperform existing random feature-based Transformers in terms of performance, training time, and peak memory consumption.
-Various novel Spectraformer combinations are more accurate than the previous SOTA models, SADERF-ORF (a.k.a., FAVOR# (Likhosherstov et al., 2023)), OPRF-ORF (a.k.a., FAVOR++ (Likhosherstov et al., 2022)), PosRF-ORF (a.k.a., FAVOR+ (Choromanski et al., 2021)), whilst significantly reducing training time and memory consumption over both.
-We hope that Spectraformer will enable the mechanism to benchmark more random feature-based Transformers for these and other tasks.
+Through our benchmarking experiments on the Long Range Arena (LRA) (Tay et al., 2021) and 18 combinations of weight matrices and component functions, Spectraformer produces novel combinations of component functions and weight matrices that perform competitively against previous efficient Transformers with different computational advantages.
+In addition, our work demonstrates that a well-configured kernelized attention model can perform on the same level as other classes of efficient Transformers, with our best Spectraformer variant proving highly competitive with leading methods like Nystromformer (Xiong et al., 2021) and BigBird (Zaheer et al., 2020).
+In doing so, it establishes a new state-of-the-art for random feature-based methods with a mean accuracy improvement of 0.35% over the Original Transformer (OT).
+Many Spectraformer variants also outperform existing efficient Transformers with random feature attention such as SADERF-ORF (a.k.a., FAVOR# (Likhosherstov et al., 2023)), OPRF-ORF (a.k.a., FAVOR++ (Likhosherstov et al., 2022)), PosRF-ORF (a.k.a., FAVOR+ (Choromanski et al., 2021)).
 Our contributions are:
-  1. 1.
-    Our novel framework, Spectraformer, which generalizes over all past works in linearized attention with kernel approximation and kernel learning.
-  2. 2.
-    We demonstrate how Spectraformer can discover novel SOTA models in the random feature Transformer family for the LRA benchmark and that these models are more performant than existing random feature Transformers.
-  3. 3.
-    We empirically identify the relationship between performant combinations and task characteristics, aiding the right kernel choice when tackling a task.
-  4. 4.
-    Our code is currently available for review, but they will be publicly released upon acceptance.
-We detail guide on how to incorporate future kernel works into Spectraformer.
-With this work, every future kernel work that is introduced can easily be incorporated into the transformer.
-2 Preliminaries
-We first cover the theoretical foundations to this work. A list of commonly used acronyms is available in Table 1.
-Table 1: Table of acronyms
-Acronym Description
-OT (Vaswani et al., 2017) Original Transformer
-RF (Rahimi & Recht, 2007) Random Feature method
-RBF Radial Basis Function kernel (also called the Gaussian kernel)
-TrigRF (Rahimi & Recht, 2007) Trigonometric RF (component function)
-PosRF (Choromanski et al., 2021) Positive RF (component function)
-OPRF (Likhosherstov et al., 2022) Optimized positive RF (component function)
-SADERF (Likhosherstov et al., 2023) Simplified Asymmetric Dense-Exponential RF (component function)
-ORF (Yu et al., 2016) Orthogonal RF (weight matrix)
-SORF (Yu et al., 2016) Structural ORF (weight matrix)
-QMC (Avron et al., 2016) Quasi- Monte Carlo (weight matrix)
-MM (Shen et al., 2017) (Liu et al., 2022) Moment Matching (weight matrix)
-SGQ (Dao et al., 2017) Sparse Grid Quadrature (weight matrix)
-FastFoodL(Yang et al., 2015) FastFood Learnable (weight matrix)
-2.1 Kernelized Attention
+  1. (1)
+    A novel framework, Spectraformer, that unifies and generalizes past work on linearized attention using both kernel approximation and kernel learning.
+  2. (2)
+    Establishment of a new state-of-the-art for random feature-based efficient Transformers, demonstrating that this model class can be highly competitive with leading sparse and low-rank methods on the Long Range Arena benchmark.
+  3. (3)
+    A comprehensive empirical analysis of 18 combinations of component functions and weight matrices, offering practical guidance on selecting configurations based on trade-offs between accuracy, memory, and training time.
+  4. (4)
+    A public and extensible implementation designed to facilitate future research by allowing for the easy integration of new kernel-based techniques into Transformers.
+2. Preliminaries
+We first cover the theoretical foundations of this work.
+2.1. Kernelized Attention
 OT attention (Vaswani et al., 2017) can be defined using the unified attention model (Galassi et al., 2021) as follows.
@@
 The general equation for $\textbf{A}_{i}$ is given as:
-$\begin{array}[]{cccc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{f(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}f(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}=\sum_{j=1}^{N}a_{j}\bm{v}_{j};&a=g(e);&e=f(\bm{q}_{i},\bm{k}_{j})\end{array}$ (1)
+(1) $\begin{array}[]{cccc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{f(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}f(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}=\sum_{j=1}^{N}a_{j}\bm{v}_{j};&a=g(e);&e=f(\bm{q}_{i},\bm{k}_{j})\end{array}$
 Attention weights are generated using a distribution function $g$ applied on an energy score $e$.
@@
 The attention in the OT is called the ‘scaled dot-product’ which defines the compatibility function $f$ and the distribution function $g$ as follows:
-$\begin{array}[]{cc}f(\bm{q}_{i},\textbf{K})=\frac{\bm{q}_{i}^{T}\textbf{K}}{\sqrt{d_{k}}}&g(\bm{x}_{j})=\frac{\exp(\bm{x}_{j})}{\sum_{l}\exp(\bm{x}_{l})}\end{array}$ (2)
+(2) $\begin{array}[]{cc}f(\bm{q}_{i},\textbf{K})=\frac{\bm{q}_{i}^{T}\textbf{K}}{\sqrt{d_{k}}}&g(\bm{x}_{j})=\frac{\exp(\bm{x}_{j})}{\sum_{l}\exp(\bm{x}_{l})}\end{array}$
 We can then have the attention matrix (with $\exp(.)$ being applied element-wise) as:
-$\begin{array}[]{cc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{\exp(\bm{q}_{i}\bm{k}_{j}^{T}/\sqrt{d_{k}})}{\sum_{l}\exp(\bm{q}_{i}\bm{k}_{l}^{T}/\sqrt{d_{k}})}\bm{v}_{j}&A=softmax(\frac{\bm{QK}^{T}}{\sqrt{d_{k}}})\bm{V}\end{array}$ (3)
+(3) $\begin{array}[]{cc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{\exp(\bm{q}_{i}\bm{k}_{j}^{T}/\sqrt{d_{k}})}{\sum_{l}\exp(\bm{q}_{i}\bm{k}_{l}^{T}/\sqrt{d_{k}})}\bm{v}_{j}&A=softmax(\frac{\bm{QK}^{T}}{\sqrt{d_{k}}})\bm{V}\end{array}$
 with $\bm{Q},\bm{K},\bm{V}$ being the matrix of queries, keys, and values respectively with each row being $\bm{q}_{i},\bm{k}_{i},\bm{v}_{i}$.
@@
 With the compatibility function $f$ as a kernel $f(\bm{q}_{i},\bm{k}_{j})=\mathcal{K}(\bm{q}_{i},\bm{k}_{j})$, Equation 1 can be rewritten as:
-$\bm{A}_{i}=\sum_{j=1}^{N}\frac{\mathcal{K}(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}\mathcal{K}(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}$ (4)
-As shown in Tsai et al.
-(2019), this is indeed the Nadaraya-Watson Kernel Estimator, with $\mathbb{E}_{p(\bm{k}_{j}|\bm{q}_{i})}[\bm{v}_{j}|X=\bm{k}_{j}]=\bm{A}_{i}$, $l_{i}\left(\bm{k}_{j})=p(\bm{k}_{j}|\bm{q}_{i}\right)$, $Y_{i}=\bm{v}_{j}$:
-$\begin{array}[]{ccc}\mathbb{E}_{l_{i}}[Y_{i}|X=\bm{u}]=\sum_{i=1}^{N}l_{i}(\bm{u})Y_{i};&l_{i}(\bm{u})=\frac{\mathcal{K}(\frac{\bm{u}-\bm{x}_{i}}{h})}{\sum_{j=1}^{n}\mathcal{K}(\frac{\bm{u}-\bm{x}_{j}}{h})}\end{array}$ (5)
+(4) $\bm{A}_{i}=\sum_{j=1}^{N}\frac{\mathcal{K}(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}\mathcal{K}(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}$
+As shown by Tsai et al. (2019), this is indeed the Nadaraya-Watson Kernel Estimator, with
+$\mathbb{E}_{p(\bm{k}_{j}|\bm{q}_{i})}[\bm{v}_{j}|X=\bm{k}_{j}]=\bm{A}_{i}$, $l_{i}\left(\bm{k}_{j})=p(\bm{k}_{j}|\bm{q}_{i}\right)$, $Y_{i}=\bm{v}_{j}$:
+(5) $\begin{array}[]{ccc}\mathbb{E}_{l_{i}}[Y_{i}|X=\bm{u}]=\sum_{i=1}^{N}l_{i}(\bm{u})Y_{i};&l_{i}(\bm{u})=\frac{\mathcal{K}(\frac{\bm{u}-\bm{x}_{i}}{h})}{\sum_{j=1}^{n}\mathcal{K}(\frac{\bm{u}-\bm{x}_{j}}{h})}\end{array}$
 Specifically, the scaled dot-product attention in the OT in Equation 3 corresponds to the softmax kernel (see Equation 6).
-$\mathcal{K}_{softmax}(\bm{q}_{i},\bm{k}_{j})=\exp(\bm{q}_{i}^{T}\bm{k}_{j}/\sqrt{d_{k}})$ (6)
-2.2 Linearizing Kernelized Attention via Random Features
+(6) $\mathcal{K}_{softmax}(\bm{q}_{i},\bm{k}_{j})=\exp(\bm{q}_{i}^{T}\bm{k}_{j}/\sqrt{d_{k}})$
+2.2. Linearizing Kernelized Attention via Random Features
 Since a kernel is the inner product of two vectors in some space with respect to some feature map $\phi$, i.e., $\mathcal{K}(\bm{x},\bm{y})=\phi(\bm{x})\phi(\bm{y})^{T}$, then we can rewrite Equation 4 as:
-$\begin{array}[]{c}\bm{A}_{i}=\frac{\sum_{j}\phi(\bm{q}_{i})\phi(\bm{k}_{j})^{T}\bm{v}_{j}}{\sum_{l}\phi(\bm{q}_{i})\phi(\bm{k}_{l})^{T}}=\frac{\phi(\bm{q}_{i})\sum_{j}\phi(\bm{k}_{j})\otimes\bm{v}_{j}}{\phi(\bm{q}_{i})\sum_{l}\phi(\bm{k}_{l})^{T}}\end{array}$ (7)
+(7) $\begin{array}[]{c}\bm{A}_{i}=\frac{\sum_{j}\phi(\bm{q}_{i})\phi(\bm{k}_{j})^{T}\bm{v}_{j}}{\sum_{l}\phi(\bm{q}_{i})\phi(\bm{k}_{l})^{T}}=\frac{\phi(\bm{q}_{i})\sum_{j}\phi(\bm{k}_{j})\otimes\bm{v}_{j}}{\phi(\bm{q}_{i})\sum_{l}\phi(\bm{k}_{l})^{T}}\end{array}$
 If we pre-compute $\sum_{j}\phi(\bm{k}_{j})\otimes\bm{v}_{j}$ and $\sum_{l}\phi(\bm{k}_{l})^{T}$, the entire term becomes $O(1)$ and we only need to compute Equation 7 $N$ times for each $i$, resulting in an attention matrix calculated in linear time $O(N)$.
 We now refer to Equation 7 as linearized attention (Katharopoulos et al., 2020).
-All that is left is to retrieve the feature map $\phi$, the process of which is called kernel approximation (see Section 2.3), which corresponds to the kernels we specify.
-There are many popular kernels from which we can choose, but typically, we want to approximate the softmax kernel in Equation 6.
-However, since the softmax relates to the $\mathcal{K}_{RBF}(\bm{x},\bm{y})=\exp(-\frac{||\bm{x}-\bm{y}||^{2}}{2})$ via the following equation:
-$\mathcal{K}_{softmax}(\bm{x},\bm{y})=\exp(\frac{||\bm{x}||^{2}}{2})\mathcal{K}_{RBF}(\bm{x},\bm{y})\exp(\frac{||\bm{y}||^{2}}{2})$ (8)
+[…]
+[…]
 Approximating the RBF is easier than the softmax directly in the method we introduce in Section 2.3, hence we approximate the softmax indirectly via the RBF in Equation 8.
-Oftentimes, however, we might not want to specify a kernel, since kernel choice is indeed a hyperparameter and the softmax is by no means essential or necessary as shown in Tsai et al.
+Oftentimes, however, we might not want to specify a kernel, since kernel choice is indeed a hyperparameter and the softmax is by no means essential or necessary as shown by Tsai et al.
 (2019).
@@
 Notwithstanding, we first introduce the main kernel approximation technique: random features.
-2.3 Random Features for Kernel Approximation
-Of the kernel approximation family of methods, the most successful and appropriate for linearized attention in Transformer has been the random features approach (Rahimi & Recht, 2007).
+2.3. Random Features for Kernel Approximation
+Of the kernel approximation family of methods, the most successful and appropriate for linearized attention in Transformer has been the random features approach (Rahimi and Recht, 2007).
 Estimating kernels using random features relies on a fundamental insight from harmonic analysis called Bochner’s theorem (Rudin, 1990) stated as follows:
@@
 If $\mathcal{K}(\bm{\delta})$ is properly scaled, that is $\mathcal{K}(0)=1$ then the measure $p(\bm{\omega})$ is a proper probability distribution.
-inline]define unambiguously the dimension of x and phi here, also the dimension before and after approximation of the features map
-$\mathcal{K}(\bm{\delta})=\int p(\bm{\omega})\exp(i\bm{\omega\delta})\mathrm{d}\bm{\omega}$ (9)
+(9) $\mathcal{K}(\bm{\delta})=\int p(\bm{\omega})\exp(i\bm{\omega\delta})\mathrm{d}\bm{\omega}$
 When approximating the integration in this equation via Monte Carlo sampling with $s$ samples, $\bm{\omega}\sim p(\bm{\omega})$, we can obtain the feature map:
-$\begin{array}[]{ll}\mathcal{K}(\bm{\delta})&=\mathbb{E}[\phi_{\bm{\omega}}(\bm{x})^{T}\phi_{\bm{\omega}}(\bm{y})]\approx<\phi_{\bm{\omega}}(\bm{x}),\phi_{\bm{\omega}}(\bm{y})>\\ \phi_{\bm{\omega}}(\bm{x})\!&=\!\frac{1}{\sqrt{s}}\big{[}\exp(-\mbox{i}{\bm{\omega}}^{\!\top}_{1}\bm{x}),\ldots,\exp(-\mbox{i}{\bm{\omega}}^{\!\top}_{s}\bm{x})]^{\!\top}\end{array}$ (10)
-2.4 Random Features for Kernel Learning
+(10) $\begin{array}[]{ll}\mathcal{K}(\bm{\delta})&=\mathbb{E}[\phi_{\bm{\omega}}(\bm{x})^{T}\phi_{\bm{\omega}}(\bm{y})]\approx<\phi_{\bm{\omega}}(\bm{x}),\phi_{\bm{\omega}}(\bm{y})>\\ \phi_{\bm{\omega}}(\bm{x})\!&=\!\frac{1}{\sqrt{s}}\big[\exp(-\mbox{i}{\bm{\omega}}^{T}_{1}\bm{x}),\ldots,\exp(-\mbox{i}{\bm{\omega}}^{T}_{s}\bm{x})]^{T}\end{array}$
+2.4. Random Features for Kernel Learning
 Previous random feature techniques are robust in approximating kernels for specific learning problems.
@@
 The kernel is very much a hyperparameter.
-This could pose a problem since there is no free lunch in learning as much as in kernel choice (Schölkopf & Smola, 2018).
+This could pose a problem since there is no free lunch in learning as much as in kernel choice (Schölkopf and Smola, 2018).
 Tsai et al.
-(2019) has shown experimental results that this is also true in the context of attention in Transformer.
+(2019) have shown experimental results that this is also true in the context of attention in Transformer.
 Hence, instead of picking a kernel for the attention, we could learn the kernel via kernel learning an established kernel methods technique.
-Kernel learning has shown to be effective in general (Tompkins et al., 2019; Wilson & Adams, 2013; Oliva et al., 2016) and in the case of attention via Gaussian Mixture Model (GMM), learnable weight matrices (FastFood), deep generative model (DGM) (Chowdhury et al., 2022).
+Kernel learning has shown to be effective in general (Tompkins et al., 2019; Wilson and Adams, 2013; Oliva et al., 2016) and in the case of attention via Gaussian Mixture Model (GMM), learnable weight matrices (FastFood), deep generative model (DGM) (Chowdhury et al., 2022).
 Kernel learning is achieved by making the weight matrix $\bm{W}$ learnable where $\bm{\omega}_{i}$ is the output of a function parameterizing $p(\bm{\omega})$.
 The kernelized attention via random features discussed here is a popular approach, however it is by no means the only one.
-A more detailed discussion is offered in related work in Section A.2.
-Specifically, Section A.2.1 provides alternative kernelized attention formulations and Section A.2.2 provides alternative kernel approximations to random features.
-3 Spectraformer
+A more detailed discussion is offered in related work in Section 5.
+Specifically, Section 5.1 provides alternative kernelized attention formulations and Section 5.2 provides alternative kernel approximations to random features.
+Figure 1. Spectraformer framework based on Equation 11
+3. Spectraformer
 Spectraformer is a unified random feature framework that allows for the combination of any weight matrix with any component function to be used as kernelized attention in Transformer.
@@
 (2021).
-$\begin{array}[]{rl}\phi_{\bm{\omega}}(\bm{x})=&\frac{1}{\sqrt{s}}\big{[}a_{1}f_{1}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{1}(\bm{\omega}_{s},\bm{x}),...,\\ &a_{1}f_{l}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{l}(\bm{\omega}_{s},\bm{x})]^{\!\top}\\ \bm{W}=&[\bm{\omega}_{1},...,\bm{\omega}_{s}]^{T}\in\mathbb{R}^{s*d}\end{array}$ (11)
+(11) $\begin{array}[]{rl}\phi_{\bm{\omega}}(\bm{x})=&\frac{1}{\sqrt{s}}\big[a_{1}f_{1}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{1}(\bm{\omega}_{s},\bm{x}),...,\\ &a_{1}f_{l}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{l}(\bm{\omega}_{s},\bm{x})]^{\!\top}\\ \bm{W}=&[\bm{\omega}_{1},...,\bm{\omega}_{s}]^{T}\in\mathbb{R}^{s*d}\end{array}$
 Spectraformer is shown in Figure 1.
@@
 The component weight $a$ can either be fixed (typically at $a=1$) or learnable as proposed by various kernel works (Liu et al., 2022).
-Due to the fact that the majority of random feature Transformers (e.g., Likhosherstov et al.
-(2022)) have $a=1$, we follow this setting for consistent comparison between combinations.
+We set the component weight a=1 to ensure a consistent and fair comparison with prior random feature Transformers (e.g., (Likhosherstov et al., 2022)).
 The figure and the equation together highlight the three strands of research as mentioned previously in Section 1: (A) Weight matrix approximator: constructing $\bm{W}$ more effectively; (B) Weight matrix learner: parameterizing $\bm{W}$ instead of sampling; (C) Component function: constructing $f_{j}$ more effectively.
-3.1 Weight Matrices
-In Spectraformer, the weight matrix $\bm{W}$ which is either an approximator (from the families of (e.g., ORF and SORF (Yu et al., 2016)), Unit-cube (e.g., QMC (Avron et al., 2016), MM (Shen et al., 2017)), Quadrature (e.g., SGQ (Dao et al., 2017)).
-or a learner (FastFoodL (Chowdhury et al., 2022)).
-Figure 1: Spectraformer framework based on Equation 11
+3.1. Weight Matrices
+In Spectraformer, the weight matrix $\bm{W}$ is either an approximator (from the families of Monte Carlo sampling (e.g., ORF and SORF (Yu et al., 2016)), Unit-cube sampling (e.g., QMC (Avron et al., 2016), MM (Shen et al., 2017)), Quadrature (e.g., SGQ (Dao et al., 2017)).
+or a learner (e.g., FastFoodL (Chowdhury et al., 2022)).
 We term the matrix $\bm{W}$ as weight matrix instead of random matrix since although $\bm{W}$ acts like a random matrix ($\bm{\omega}_{i}\sim p(.)$, see Equation 10), it is not guaranteed to be one.
@@
 Spectraformer enables weight matrix approximation in terms of three families based on how they solve the intractable integral in Equation 9:
-  • Monte Carlo sampling-based weight matrix involves approximating the integral of Equation 9 using Monte Carlo sampling with the solution given by Equation 10 (Rahimi & Recht, 2007) with $\bm{W}$ being the ‘Base’ random matrix of $p(.)$, $\bm{\omega}_{i}\sim p(.)$.
-The random matrix $\bm{W}$ can be constructed more efficiently (either reducing time or space complexity) with FastFood (Le et al., 2013).
-The random matrix $\bm{W}$ can also be enforced to form geometrical couplings such as orthogonality to reduce approximation error with ORF, SORF (Yu et al., 2016).
-Both these methods produce valid random matrices.
-  • Unit-cube sampling-based weight matrix transforms Equation 9 to an integral on the unit cube $[0,1]^{d}$, then performs the approximation with uniform and independent points, thus reducing variance.
+  • Monte Carlo sampling-based weight matrix involves approximating the integral of Equation 9 using Monte Carlo sampling with the solution given by Equation 10 (Rahimi and Recht, 2007) with $\bm{W}$ being the ‘Base’ random matrix of $p(.)$, $\bm{\omega}_{i}\sim p(.)$.
+There are additional methods which improve over Base: structural and geometric methods.
+Structural methods decompose the Base random matrices $P$ into smaller matrices which reduce time and/ or space complexity.
+Structural matrices includes FastFoodF (Chowdhury et al., 2022), SCRF (Feng et al., 2015) and P-Model (Choromanski and Sindhwani, 2016).
+Geometric methods enforce certain geometrical couplings along some dimensions in the construction of random matrices to reduce the approximation error.
+Geometric methods include: ROM (Choromanski et al., 2017) (ORF (Yu et al., 2016), SORF (Yu et al., 2016)) and SimRF (Reid et al., 2023).
+In addition to these methods, Monte Carlo has also been adapted as Random Fourier Signature Features (RFSF) (Tóth et al., 2025) for unbiased estimation of the signature kernel, designed for sequential data.
+RFSF also has more scalable variants, such as RFSF-DP (Diagonally Projected) and RFSF-TRP (Tensor Random Projection), that use additional dimensionality reduction.
+  • Unit-cube sampling-based weight matrix transforms Equation 9 to an integral on the unit cube $0,1]^{d}$, then performs the approximation with uniform and independent points, thus reducing variance.
 The first approach is QMC (Avron et al., 2016), where $\bm{W}:\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$, $\Phi$ being the cumulative distribution function (CDF) associated with $p(.)$ and $\bm{t}_{i}$ being a low discrepancy sequence.
@@
 Smolyak rule is further implemented to alleviate the curse of dimensionality.
-Liu et al.
-(2022) show that in various kernel benchmark datasets (among combination with TrigRF), QMC and ORF have consistently high performance.
-Due to the significant number of weight matrices in random features literature, we decide to cover only the most prominent ones.
-Appendix A.3 provides the technical details of the approximators discussed here and those that we left out.
 Spectraformer also allows to learn weight matrices by learning $p(.)$ via parameterizing $\bm{W}$.
@@
 (2022).
-The best performing method is via making the constituent matrices of FastFood learnable, which we term FastFoodL.
-In general, a weight matrix learner is valid as long as it can implicitly model any distribution.
-In addition the TrigRF, PosRF has also been applied on FastFoodL.
-Technical details and other learners we have not covered are explored in Appendix A.4.
-We now explore existing works which have developed from Rahimi & Recht (2007) (see Equation 9, 10) and how they lead to the development of Spectraformer.
-3.2 Component functions
+This comes from a long line of work in kernel learning.
+The approach to the dual objectives separates the kernel learning methods into two-stage and one-stage methods (Liu et al., 2022).
+Two-stage methods (e.g., (Yu et al., 2015; Wilson and Adams, 2013; Bullins et al., 2018)) solves the dual objective separately: $\bm{W}$ is typically first learned via a kernel alignment scheme (Wang et al., 2015), then we solve for $\bm{v}$.
+One-stage methods (e.g., (Chowdhury et al., 2022; Tompkins et al., 2019)), on the other hand, solves the dual objective in parallel: $p(.)$ corresponding to $\bm{W}$ is parameterized then $\bm{v}$ is solved typically.
+It is not plausible to apply kernel alignment (Wang et al., 2015) on Transformer architectures.
+Therefore, only one-stage methods can be considered in the context of weight matrix learner.
+Two kernel learning methods, GMM and FastFoodL, have been experimented in the non-Transformer setting (Yang et al., 2015), then adapted into the Transformer setting (Chowdhury et al., 2022).
+DGM (Deep Generative Model) has also been effective in modeling distributions (Kingma and Welling, 2013).
+It too has been adapted by Chowdhury et al.
+(2022) for the Transformer setting.
+Chowdhury et al.
+(2022) show the superior performance of FastFoodL over other kernel learning-based techniques, hence we have only experimented with FastFoodL among kernel learning techniques.
+Liu et al.
+(2022) show that in various kernel benchmark datasets (among combination with TrigRF), QMC and ORF have consistently high performance.
+Due to the significant number of weight matrices in random features literature and page limit, we decide to cover only the most prominent ones.
+Methods which we will not cover in detail includes: SCRF (Feng et al., 2015), P-Model (Choromanski and Sindhwani, 2016), ROM (Choromanski et al., 2017), SimRF (Choromanski et al., 2017), SSF (Lyu, 2017), SSR (Munkhoeva et al., 2018), GMM (Wilson and Adams, 2013; Oliva et al., 2016), DGM (Kingma and Welling, 2013).
+We still mention them and include them in the figures where appropriate to motivate future work.
+Figure 2 shows all weight matrices that are implemented and discussed.
+We now explore these weight matrices in the order specified above.
+Figure 2.
+Spectraformer weight matrices.
+Dark blue boxes are technique groupings.
+Arrow denotes which technique belongs to which grouping.
+Red boxes are techniques which have been explored in the literature but not in this work.
+Green boxes are techniques which have been explored in the literature and this work.
+Light blue boxes are techniques not studied empirically previously.
+3.1.1. Base
+(Rahimi and Recht, 2007) This is the direct sampling from a distribution $\bm{P}_{ij}\sim p$, corresponding to the kernel $\mathcal{K}$ without any additional adjustment, and in approximating the RBF, $P=G$ where $G$ is the normal distribution.
+(12) $W_{Base}=\frac{1}{\sigma}\bm{P}$
+3.1.2. FastFoodF
+FastFoodF (Le et al., 2013) makes use of Hadamard and diagonal matrices to speed up Gaussian matrices $G$ constructions for RFF in $O(n\log{d})$ (100 times faster) with $O(n)$ space (1000 times less space) given $n\geq d$ against Base.
+However, FastFoodF increases variance and approximation error and decreases the concentration bound.
+(13) $W_{FastFood_{L}}=\frac{1}{\sigma}\bm{SHG\Pi HB}$
+[…]
+3.1.3. ORF
+Orthogonal RF (Yu et al., 2016) decreases the approximation error significantly compared to Base.
+This is done via replacing $G$ with a properly scaled random orthogonal matrix.
+However, generating orthogonal matrices become costly quickly as the number of dimensions increases.
+(14) $\bm{W}_{ORF}=\frac{1}{\sigma}\bm{SQ}$
+where $\bm{Q}$ is a uniformly distributed random orthogonal matrix (on the Stiefel manifold) obtained from the QR decomposition of $\bm{G}$, the set of rows of $\bm{Q}$ forming a basis in $\mathbb{R}^{d}$, and $\bm{S}$ is a diagonal matrix with entries sampled i.i.d from the $\chi$-distribution with $d$ degrees of freedom, thus making the rows of $\bm{SQ}$ and $\bm{G}$ identically distributed.
+3.1.4. SORF
+Structured ORF (Yu et al., 2016) decreases time and space complexity of ORF (from $O(d^{2})$ to $O(d\log{d})$ with almost no extra memory cost) by imposing structure on the orthogonal matrices, inspired by structural methods.
+SORF is unbiased with large $d$.
+We replace $\bm{S}$ with $\sqrt{d}$ and $\bm{Q}$ with structured matrix $\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$
+(15) $\bm{W}_{SORF}=\frac{\sqrt{d}}{\sigma}\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$
+where $\bm{D}_{i}\in\mathbb{R}^{d*d},i=1,2,3$ is diagonal ’sign-flipping‘ matrices with diagonal entries sampled from the Rademacher distribution, $\bm{H}$ is the normalized Walsh-Hadamard matrix.
+3.1.5. QMC
+Quasi-Monte Carlo (Avron et al., 2016) evaluates on a low discrepancy sequence (e.g., Halton, Sobol’ Faure, and Niederreiter) of points instead of random points in Monte Carlo.
+Although the approximation error is only reduced minimally, QMC has been shown to perform better than MC in high dimensions and does not have undesirable clustering effect.
+To calculate QMC, we first assume that $p(\bm{x})=\prod^{d}_{j=1}p_{j}(\bm{x}_{j})$ factorizes with respect to the dimensions with $p_{j}(.)$ being a univariate density function.
+Then we define:
+(16) $\Phi^{-1}(\bm{t})=\left(\Phi_{1}^{-1}\left(\bm{t}_{1}\right),\ldots,\Phi_{d}^{-1}\left(\bm{t}_{d}\right)\right)\in\mathbb{R}^{d}$
+where $\Phi_{j}$ being the cumulative distribution function of $p_{j}$, $\bm{t}_{1},\bm{t}_{2},...,\bm{t}_{s}\in[0,1]^{d}$ being a low discrepancy sequence, and $\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$.
+We can thus transform the integral on $\mathbb{R}^{d}$ in Equation 9 to an integral on the unit cube $0,1]^{d}$ as
+$\mathcal{K}(\bm{x}-\bm{x}^{\prime})=\int_{[0,1]^{d}}\exp\big(\mbox{i}(\bm{x}-\bm{x}^{\prime})^{\!\top}\Phi^{-1}(\bm{t})\big)\mbox{d}\bm{t}$
+The weight matrix then can be defined as:
+$\bm{W}_{\text{QMC}}=[\Phi^{-1}(\bm{t}_{1}),\Phi^{-1}(\bm{t}_{2}),\ldots,\Phi^{-1}(\bm{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}\;.$
+QMC can be further improved by a sub-grouped based rank-one lattice construction which improved complexity (Lyu et al., 2020).
+3.1.6. MM
+Moment Matching (Shen et al., 2017; Liu et al., 2022) improves over QMC by removing the undesirable clustering effect and having the same approximation error with less features.
+This is done by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$:
+$\bm{W}_{\text{MM}}=[\widetilde{\Phi}^{-1}(\bm{t}_{1}),\widetilde{\Phi}^{-1}(\bm{t}_{2}),\ldots,\widetilde{\Phi}^{-1}(\bm{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}$
+[…]
+3.1.7. SGQ
+Sparse Grid Quadrature (Dao et al., 2017) evolves from GQ.
+GQ (Gaussian Quadrature) (Dao et al., 2017) assumes that the kernel $k$ factorizes with respect to the dimensions and thus can be approximated by a one-dimensional Gaussian quadrature rule.
+However the total number of points $s$ scale exponentially with the dimensions.
+However GQ suffers form the curse of dimensionality.
+This is alleviated using Smolyak rule resulting in SGQ.
+Assuming the third-degree SGQ using symmetric univariate quadrature points $\left\{-\hat{p}_{1},0,\hat{p}_{1}\right\}$ with weights $(\hat{a}_{1},\hat{a}_{0},\hat{a}_{1})$, then we have:
+$\bm{W}_{\text{SGQ}}\!=\![\bm{0}_{d},\hat{p}_{1}\bm{e}_{1},\ldots,\hat{p}_{1}\bm{e}_{d},-\hat{p}_{1}\bm{e}_{1},\ldots,-\hat{p}_{1}\bm{e}_{d}]^{\!\top}\!\in\!\mathbb{R}^{(2d+1)\times d}\!$
+given that $\mathbb{e}_{i}$ is the d-dimensional standard basis vector with the ith element being 1.
+3.1.8. FastFoodL
+This (see Equation 13) is proposed to be learnable by Yang et al.
+(2015) with two variations: FSARD (scaling matrix $\bm{S}$ previously sampled from chi-squared is made learnable) and FSGBARD (optimizes the marginal likelihood with respect to the diagonal matrices $\bm{G}$ and $\bm{B}$).
+It is further adapted by Chowdhury et al.
+(2022) to make $\bm{S}$ or $\bm{S},\bm{G},\bm{B}$ learnable parameters.
+In our experiments, we maintain this set up.
+Figure 3.
+Spectraformer component functions.
+Dark blue boxes are technique groupings.
+Arrow denotes which technique belongs to which grouping.
+Red boxes are techniques which have been explored in the literature but not in this work.
+Green boxes are techniques which have been explored in the literature and this work.
+Light blue boxes are techniques not studied empirically previously.
+3.2. Component functions
 Spectraformer incorporates component functions $f$ such as PosRF, OPRF, SADERF.
@@
 In addition, PosRF, OPRF, and SADERF have also been proven theoretically to approximate $\mathcal{K}_{RBF}$ given $\bm{W}$ being the ORF.
-While TrigRF(Rahimi & Recht, 2007) is the base component function and guarantees $\mathcal{K}(\bm{\delta})$ to be close to $\phi_{TrigRF}(\bm{x})\cdot\phi_{TrigRF}(\bm{y})$ due to uniform convergence, it is typically evaluated using only the real component of $\exp(-i\bm{\omega}_{j}^{T}\bm{x})$.
-This leads to $\cos$ and $\sin$ functions which can produce unstable behavior, especially when $K(\bm{\delta})$ is close to 0, where variance is infinite (Choromanski et al., 2021; Likhosherstov et al., 2022).
-Therefore, TrigRF is avoided in Spectraformer.
-  • PosRF (Choromanski et al., 2021) is introduced in Spectraformer as a simple modification over TrigRF to counter this instability.
-PosRF guarantees positive function output and its variance approaching 0 as $K(\bm{\delta})$ approaches 0.
-Likhosherstov et al.
-(2022) generalize TrigRF and PosRF to GERF.
-  • OPRF (Likhosherstov et al., 2022) is one of the solution to the reduction of the variance equation of GERF.
-OPRF is both positive and bounded (whereas PosRF is positive and unbounded).
-Likhosherstov et al.
-(2023) generalize GERF to dense matrix instead of scalar parameters leading to DERF.
-  • SADERF (Likhosherstov et al., 2023) is one of the solutions to the reduction of the variance equation of DERF.
-Whilst DERF has many solutions, they all rely on operations which are incompatible with many deep learning libraries on GPU and TPU like SVD and eigen decompositions.
-SADERF is an extension of GERF with a tighter variance bound on DERF.
 The alternatives for component functions allow Spectraformer to approximate the RBF with $p(.)$ corresponding to the Gaussian.
 Likhosherstov et al.
-(2023) and Likhosherstov et al.
-(2022) show that SADERF-ORF is the SOTA combination followed by OPRF-ORF and PosRF-ORF.
-Technical details and other component function(s) not covered are explored in Appendix A.5.
-Table 2: Summary of research gap in kernelized attention.
-Rows are the component functions, columns are the weight matrices.
+(2023, 2022) show that SADERF-ORF is the SOTA combination followed by OPRF-ORF and PosRF-ORF.
+Due to page limits, we will not discuss component functions which are not included in our empirical experiments.
+These techniques are still included in figures and mentioned where necessary to motivate future work in the area.
+These include DIRF (Likhosherstov et al., 2022) techniques (PoisRF, GeomRF) and HRF (Choromanski et al., 2022).
+3.2.1. CDERF
+Complex dense exponential random feature is given in Equation 17, where $\bm{\omega},\bm{x}\in\mathbb{R}^{d}$, $p=\mathcal{N}(0,1)^{d}$, $k\in\{1,2\}$ and $\mathbb{S}_{d_{c}}$ being a set of ddd*d complex symmetric matrices.
+This is the general formula for most of the component functions we are interested in.
+The parameters and constraints of different CDERF functions are specified in Table 2:
+(17) $f_{DE}^{(k)}(\bm{\omega},\bm{x})=Real(D\exp(\bm{\omega}^{T}\bm{A}\bm{\omega}+\bm{\omega}^{T}\bm{B}^{(k)}\bm{x}+\bm{x}^{T}\bm{C}^{(k)}\bm{x}))$
+Table 2. An overview of the parameters and constraints for different functions within the Complex Dense Exponential Random Feature (CDERF) family.
+$\bm{A}$ $\bm{B}^{(1)}$ $\bm{B}^{(2)}$ $\bm{C}^{(1)}$ $\bm{C}^{(2)}$ $D$ $\bm{x}$ $\bm{y}$
+CDERF $\mathbb{S}_{d_{C}}$ $\mathbb{C}^{d*d}$ $\mathbb{C}^{d*d}$ $\mathbb{R}$ $\bm{x}$ $\bm{y}$
+DERF (Likhosherstov et al., 2023) $\mathbb{S}_{d}$ $\mathbb{R}^{d*d}$ $\mathbb{R}^{d*d}$ $\mathbb{R}$ $\bm{x}$ $\bm{y}$
+SADERF (Likhosherstov et al., 2023) $A_{GE}\bm{I}_{d}$ $B_{GE}\bm{I}_{d}$ $C_{GE}\bm{I}_{d}$ $D_{GE}$ $\Psi\bm{x}$ $\Psi^{-1}\bm{y}$
+GERF (Likhosherstov et al., 2022) $A_{GE}\bm{I}_{d}$ $B_{GE}^{(1)}\bm{I}_{d}$ $B_{GE}^{(2)}\bm{I}_{d}$ $C_{GE}\bm{I}_{d}$ $D_{GE}$ $\bm{x}$ $\bm{y}$
+TrigRF (Rahimi and Recht, 2007) $0\bm{I}_{d}$ $i\bm{I}_{d}$ $i\bm{I}_{d}$ $0\bm{I}_{d}$ 11 $\bm{x}$ $\bm{y}$
+PosRF (Choromanski et al., 2021) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $\bm{I}_{d} 11$ $\bm{x}$ $\bm{y}$
+OPRF (Likhosherstov et al., 2022) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $\bm{I}_{d} 11$ $\bm{x}$ $\bm{y}$
+3.2.2. TrigRF
+Trigonometric RF (Rahimi and Recht, 2007) is the base RFF implementation.
+(18) $\begin{array}[]{c}f_{TrigRF}^{(1)}(\bm{\omega},\bm{x})=\sqrt{2}\cos(\bm{\omega}^{T}\bm{x}+\bm{b})\\ f_{TrigRF}^{(2)}(\bm{\omega},\bm{y})=\sqrt{2}\sin(\bm{\omega}^{T}\bm{y}+\bm{b})\\ \bm{b}\sim Uniform(0,2\pi)\end{array}$
+The use of the trigonometric sine and cosine functions leads to unstable behavior when the inputs have negative dimension-values.
+This can be further exacerbated when the values TrigRF try to approximate are close to 0 (since most values are of low significance).
+This causes the variance to approach infinity (Choromanski et al., 2021; Likhosherstov et al., 2022).
+Therefore, we do not want to use TrigRF to perform kernel approximation in attention.
+3.2.3. PosRF
+Positive RF (Choromanski et al., 2021) fixes the problem of TrigRF by enforcing positive component function output in the softmax.
+The variance of PosRF (in contrast to the variance of TrigRF approaching infinity) approaches 0 as the approximated value of the Softmax kernel approaches 0.
+PosRF has two forms: Positive RF Base (PosRF-B) (see Equation 19) and Positive RF Hyperbolic (PosRF-Hyp), which is multi-component, i.e., Equation 11: $l=2$, (see Equation 20).
+[…]
+[…]
+We only use PosRF-B in our experiments.
+3.2.4. GERF
+Generalized exponential RF (Likhosherstov et al., 2022) generalizes both TrigRF and PosRF with Equation 21.
+[…]
+where $\sqrt{.}$ and $\sqrt[n$]{.} denoting a principal root with a complex argument
+3.2.5. OPRF
+Optimized positive RF (Likhosherstov et al., 2022) is the solution to the minimization of the variance of GERF.
+Specifically it is defined as Equation 21 with $s=+1$, $\bm{x}+\bm{y}||^{2}>0$, and $\bm{A}\in\mathbb{R}$ defined in terms of $p^{*}$ as:
+[…]
+Whilst PosRF is not bounded, OPRF is. OPRF can provide $e^{60}$ $\times$ variance reduction in estimating the Softmax compared to TrigRF.
+3.2.6. DERF
+Dense-exponential random features (Likhosherstov et al., 2023) extends GERF and replace $\bm{A},\bm{B},\bm{C}$ with dense matrices.
+DERF is CDERF when $\bm{B},C$ are in the real instead of the complex plane.
+Minimizing the variance of DERF leads to two approaches: ADERF and SDERF.
+However both these approaches rely on SVD and eigen decompositions which are not extensively supported on GPU and deep learning libraries.
+Therefore SADERF is proposed.
+3.2.7. SADERF
+Simplified ADERF (Likhosherstov et al., 2023) is a special case of ADERF and extends GERF, requiring only basic unary operations in addition.
+(23) $\begin{array}[]{c}f_{SADE}^{(1)}(\bm{\omega},\bm{x})=f_{GE}^{(1)}(\bm{\omega},\Psi\bm{x}),f_{SADE}^{(2)}(\bm{\omega},\bm{y})=f_{GE}^{(2)}(\bm{\omega},\Psi^{-1}\bm{y})\\ \Psi_{l,l}^{*}=(\sum_{j}(\bm{y}_{l}^{(i)})^{2}/\sum_{i}(\bm{x}_{l}^{(i)})^{2})^{1/4}\end{array}$
+Table 3.
+A summary of research gap in the literature on kernelized attention.
+The table shows combinations of weight matrices (columns) and component functions (rows).
 For each cell, the top item is the cited work for this combination for typical kernel tasks, while the bottom item is the cited work for Transformer-based tasks.
-$\times$ indicates no prior work for the combination.
-Spectraformer explores all combinations listed here, excluding ‘Base’ due to exhaustive work as seen in the table.
-TrigRF PosRF OPRF SADERF
-Monte Carlo Base (Rahimi & Recht, 2007) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
-(Choromanski et al., 2021) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
-ORF (Yu et al., 2016) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
-(Choromanski et al., 2021) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
-SORF (Yu et al., 2016) $\times$ $\times$ $\times$
+$\times$ indicates that a combination has not been previously studied in either general kernel tasks or specifically for Transformers.
+Spectraformer explores all listed combinations (excluding ‘Base’ due to exhasutive pre-existing work)
+$f$ Base ORF SORF QMC $M$ SGQ FastFoodL
+TrigRF
+(Rahimi and Recht, 2007)
+(Choromanski et al., 2021)
+(Yu et al., 2016)
+(Choromanski et al., 2021)
+(Yu et al., 2016)
+$\times$
+(Avron et al., 2016)
+$\times$
+(Shen et al., 2017)
+$\times$
+(Dao et al., 2017)
+$\times$
+(Yang et al., 2015)
+(Chowdhury et al., 2022)
+PosRF
+(Choromanski et al., 2021)
+(Choromanski et al., 2021)
+(Choromanski et al., 2021)
+(Choromanski et al., 2021)
 $\times$ $\times$ $\times$ $\times$
-Unit-cube QMC (Avron et al., 2016) $\times$ $\times$ $\times$
-$\times$ $\times$ $\times$ $\times$
-MM (Shen et al., 2017) $\times$ $\times$ $\times$
-$\times$ $\times$ $\times$ $\times$
-Quadrature SGQ (Dao et al., 2017) $\times$ $\times$ $\times$
-$\times$ $\times$ $\times$ $\times$
-Learner FastFoodL (Yang et al., 2015) $\times$ $\times$ $\times$
-(Chowdhury et al., 2022) (Chowdhury et al., 2022) $\times$ $\times$
-3.3 Utility
+$\times$
+(Chowdhury et al., 2022)
+OPRF
+(Likhosherstov et al., 2022)
+(Likhosherstov et al., 2022)
+(Likhosherstov et al., 2022)
+(Likhosherstov et al., 2022)
+$\times$ $\times$ $\times$ $\times$ $\times$
+SADERF
+(Likhosherstov et al., 2023)
+(Likhosherstov et al., 2023)
+(Likhosherstov et al., 2023)
+(Likhosherstov et al., 2023)
+$\times$ $\times$ $\times$ $\times$ $\times$
+3.3. Utility
 There has been an unsystematic attempt at combining component functions and weight matrices.
@@
 Spectraformer is a generic random feature framework that allows for the combination of any weight matrix with any component function provided that the weight matrix satisfies Definition 3.1 and the component function satisfies Definition 3.2.
+Instructions on adding new component functions or weight matrices to Spectraformer are detailed step by step in the repository linked in the abstract.
 We note that Spectraformer is an experimental rather than a theoretical framework.
-Due to the large number of weight matrices and component functions, it is infeasible for the time being to derive a general mathematical proof for the bounds and convergence of Spectraformer.
-Therefore, we seek to validate its feasibility and show its ability to discover novel and performant combination via experimental results.
-3.4 Adding New Code to Spectraformer
-We now discuss how to add a new component function or weight matrix to the base code for further implementation.
-New component functions need to satisfy Definition 3.2 and new weight matrices need to satisfy Definition 3.1.
-3.4.1 Component Function
-  1. 1.
-    Add the new component function f to src/models/component_functions.py, the arguments should include data (the input), and other optional parameters.
-  2. 2.
-    Import f and add a new entry to FastAttention.comp_functions[f_name] = f in src/models/attention_performer.py (line 176)
-3.4.2 Weight Matrix
-  1. 1.
-    Add the new weight matrix w to src/models/weight_matrix_approx.py or src/models/weight_matrix_learner.py, the arguments should include nb_rows (number of rows), nb_cols (number of columns) and device.
-  2. 2.
-    Import w and add the if clause and $w$ function call in src/models/attention_performer.py (line 208)
-4 Results
-4.1 Experiment Details
-Table 3: Experimental results of Spectraformer variants on the LRA benchmark, over five seeds.
-We report mean accuracy (on the test set), mean time (training time (hour)), and mean memory (peak memory consumption (GB)).
-$L$: ListOps, $T$: Text, $R$: Retrieval, and $\mu$ for the average of the three.
-Entries are ranked by $\mu$ accuracy.
-The previous RF SOTAs (SADERF-ORF (Likhosherstov et al., 2023) and OPRF-ORF (Likhosherstov et al., 2022)) are in bold.
-We note that via Spectraformer, we find various variants outperforming these previous RF SOTAs in terms of accuracy, with different combinations offering different trade-off between accuracy, time, and memory.
-Entries are sorted by mean accuracy.
-Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
-L T R $\mu$ L T R $\mu$ L T R $\mu$
-OPRF-FastFoodL 37.55 (0.48) 64.41 (0.62) 77.70 (0.33) 59.89 1.07 2.07 2.12 1.75 0.86 1.72 1.68 1.42
-OPRF-MM 38.08 (0.53) 60.40 (0.85) 81.09 (0.18) 59.86 0.68 1.26 1.24 1.06 1.36 2.71 2.56 2.21
-PosRF-MM 37.06 (0.37) 61.87 (1.79) 80.58 (0.53) 59.84 0.56 1.05 1.06 0.89 1.17 2.31 2.10 1.86
-OPRF-ORF 38.34 (0.22) 60.16 (0.79) 80.88 (0.17) 59.80 0.68 1.26 1.25 1.06 1.36 2.71 2.56 2.21
-SADERF-QMC 37.37 (0.38) 61.14 (1.48) 80.84 (0.11) 59.78 0.68 1.25 1.29 1.07 1.44 2.86 2.69 2.33
-PosRF-QMC 37.11 (0.09) 61.69 (0.96) 80.55 (0.13) 59.78 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
-SADERF-MM 37.10 (0.22) 60.68 (1.88) 81.13 (0.17) 59.64 0.68 1.25 1.29 1.07 1.44 2.86 2.69 2.33
-SADERF-ORF 37.10 (0.19) 60.39 (2.08) 81.05 (0.22) 59.51 0.68 1.25 1.28 1.07 1.44 2.86 2.69 2.33
-OPRF-QMC 37.69 (0.62) 59.94 (0.59) 80.38 (0.49) 59.34 0.68 1.26 1.26 1.07 1.36 2.71 2.56 2.21
-SADERF-SGQ 37.11 (0.21) 62.46 (0.54) 78.38 (0.25) 59.32 0.68 1.25 1.27 1.07 1.44 2.86 2.69 2.33
-SADERF-FastFoodL 36.02 (1.38) 64.63 (0.18) 76.99 (0.61) 59.21 1.07 2.08 2.16 1.77 0.92 1.84 1.80 1.52
-OPRF-SGQ 37.10 (0.23) 61.25 (0.54) 78.69 (0.54) 59.01 0.67 1.26 1.25 1.06 1.36 2.71 2.56 2.21
-PosRF-ORF 34.35 (5.96) 60.30 (0.97) 80.45 (0.22) 58.37 0.56 1.05 1.06 0.89 1.17 2.31 2.10 1.86
-PosRF-FastFoodL 33.46 (3.70) 64.65 (0.36) 76.95 (0.48) 58.35 1.02 1.98 2.03 1.68 0.79 1.57 1.53 1.30
-SADERF-SORF 33.30 (0.98) 64.70 (0.36) 74.71 (1.90) 57.57 0.68 1.24 1.28 1.07 1.44 2.86 2.69 2.33
-PosRF-SGQ 28.64 (7.54) 62.38 (0.53) 78.28 (0.20) 56.43 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
-OPRF-SORF 27.91 (3.26) 64.76 (0.66) 75.92 (1.74) 56.20 0.67 1.26 1.24 1.06 1.36 2.71 2.56 2.21
-PosRF-SORF 21.27 (6.65) 62.99 (0.40) 67.10 (1.11) 50.45 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
-The LRA covers tasks of different sequence lengths, difficulty, and objective and it is designed to evaluate efficient Transformers, of which Spectraformer belongs to.
-Specifically, we evaluate the models on three textual LRA tasks: ListOps (Nangia & Bowman, 2018), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard & Ruder, 2018), and Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013).
+Aside from the challenge of generalizing the large number of weight matrices and component functions, estimating the “actual error of the approximation, or how this error will propagate into downstream learning tasks” is not possible (Yao et al., 2023).
+This is because, as Yao et al.
+(2023) pointed out: the RFF literature has largely use theoretical bounds which are largely impractical due to being either highly conservative or making use of “unknown qualities” as evident in Section 3.
+Thus, we seek to validate its feasibility and show its ability to discover novel and performant combination via empirical results and statistical analyses.
+4. Results
+4.1. Experimental Details
+The LRA covers tasks of different sequence lengths, difficulty, and objective, and is designed to evaluate efficient Transformers.
+The LRA is preferred over other benchmarks such as GLUE (Linzen et al., 2018) since efficient Transformers are designed to address tasks with long input length which reduces Transformer’s training and inference speed to the default attention’s quadratic time complexity.
+The LRA consists of ListOps (Nangia and Bowman, 2018) (testing the parsing ability of models), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard and Ruder, 2018), Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013), Image (Image Classification on Sequence of Pixels in 1D projected from the 2D of images in the CIFAR-10 dataset (Krizhevsky et al., 2009)), and Pathfinder (Long-Range Spatial Dependency) (Linsley et al., 2018; Kim et al., [n.
+d.]).
 We discuss the task characteristics in greater details in Section 4.3.
 All our models are experimented on NVIDIA V100 and 12 CPU with 20GB of memory.
-Our codebase is based on and includes the adapted or original implementation from Chen et al.
-(2021); Chowdhury et al.
-(2022); Liu et al.
-(2022); Choromanski et al.
-(2021); Likhosherstov et al.
-(2022; 2023); Wang et al.
-(2020); Xiong et al.
-(2021), Kitaev et al.
-(2020); Thomas et al.
-(2018); Zhou et al.
-(2021) (Apache License, Version 2.0).
-We run our experiments on 5 different seeds due to computational limitation.
-The parameters are identical to Chen et al.
-(2021) and chosen to limit the parameter count to account for the marked training time of multiple models.
-All the parameters for Spectraformer variants are kept identical for fair comparison, with the only difference being their random feature component.
-Our code is implemented in Python 3.12 and Pytorch, we use the Transformer base from Chen et al.
-(2021).
+Our codebase is based on and includes the adapted or original implementation from (Chen et al., 2021; Chowdhury et al., 2022; Liu et al., 2022; Choromanski et al., 2021; Likhosherstov et al., 2022, 2023; Wang et al., 2020; Xiong et al., 2021), (Kitaev et al., 2020; Thomas et al., 2018; Zhou et al., 2021) (Apache License, Version 2.0).
+The parameters are identical to (Chen et al., 2021) and chosen to limit the parameter count to account for the marked training time of multiple models.
+All the parameters for Spectraformer variants are kept identical for fair comparison, with the only difference being the number of random feature.
+We vary the number of random feature between 64, 128, 256, and 512 for comprehensive analysis.
+Dimension as 128 was used in (Chen et al., 2021) and 256 was used in (Choromanski et al., 2021; Likhosherstov et al., 2022, 2023).
 Our hyperparameters are available in Table 6.
+Our code is implemented in Python 3.12 and Pytorch, we use the Transformer base from (Chen et al., 2021).
 The model name convention is [component function]-[weight matrix] (e.g., OPRF-FastFoodL).
-Our result in Table 3 shows that the top mean accuracy models, ranked in descending order, are: OPRF-FastFoodL, OPRF-MM, PosRF-MM, OPRF-ORF, SADERF-QMC.
-Our mean accuracy result indicates that novel Spectraformer combinations have outperformed the highlighted previous SOTA, i.e., OPRF-ORF and SADERF-ORF.
-However, we are aware that mean accuracy is insufficient in offering a comprehensive picture.
-Therefore, we first proceed with analyzing the statistics of the tasks before analyzing the significance of these results.
-4.2 Task Characteristics
-We first analyze the characteristics of the three textual classification LRA (Tay et al., 2021) tasks: ListOps (Nangia & Bowman, 2018), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard & Ruder, 2018), and Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013).
+4.2. Overall Results
+Our comprehensive experiments on the Long Range Arena (LRA) benchmark reveal that the Spectraformer framework not only produces highly competitive models but also establishes a new state-of-the-art for random feature-based efficient Transformers.
+Before diving into a detailed breakdown, we first present a high-level comparison of our top-performing variants against a wide range of established efficient Transformers in Table 1.
+This includes methods based on downsampling (Nystromformer (Xiong et al., 2021), Skyformer (Chen et al., 2021), BigBird (Zaheer et al., 2020)), learnable patterns (Reformer (Kitaev et al., 2020)), and other linearization approaches (Linformer (Wang et al., 2020)).
+As shown in Table 1, our best model, OPRF-FastFoodL, achieves a mean accuracy of 57.17%, which is highly competitive with leading methods like Nystromformer (Xiong et al., 2021) (58.06%) and BigBird (Zaheer et al., 2020) (57.25%).
+This is a significant finding, as it demonstrates for the first time that a well-configured kernelized attention model can perform at the same level as other classes of efficient Transformers on this challenging benchmark.
+Notably, OPRF-FastFoodL is on par with BigBird (Zaheer et al., 2020) on mean accuracy while offering a substantial reduction in peak memory consumption (4.65 to 1.82 GB) and training time (3.32 hours to 2.26 hours)
+To provide full transparency and detail on how our top variants were identified, Table 7 presents the complete results for all 18 combinations of component functions and weight matrices explored within Spectraformer, tested with random feature dimensions of 64, 128, 256, and 512.
+The previous state-of-the-art random feature models, OPRF-ORF and SADERF-ORF, are highlighted in bold for direct comparison.
+The results in Table 8 confirm that several novel combinations discovered through our framework, such as OPRF-FastFoodL and PosRF-MM, consistently outperform the previous random feature SOTAs.
+To help navigate these extensive results, Table 4 summarises the best-performing model for each task according to accuracy, training time, and memory usage.
+This highlights the framework’s ability to generate variants that offer different trade-offs, catering to specific computational or performance requirements.
+We now proceed to a deeper analysis of these results.
+Table 4. A summary of the best-performing Spectraformer models, with tasks as columns and optimization categories as rows. Results are shown for 64, 128, 256, and 512 feature settings.
+Metric Feat Task
+L T R I P $\mu$
+Accuracy 64 OPRF-QMC OPRF-SORF OPRF-$M$ SADERF-FastFoodL PosRF-QMC OPRF-FastFoodL
+128 OPRF-$M$ SADERF-SORF OPRF-ORF OPRF-FastFoodL OPRF-SGQ OPRF-FastFoodL
+256 OPRF-FastFoodL SADERF-SORF SADERF-ORF OPRF-FastFoodL OPRF-FastFoodL OPRF-FastFoodL
+512 OPRF-FastFoodL SADERF-SORF OPRF-$M$ OPRF-FastFoodL PosRF-ORF OPRF-FastFoodL
+Time 64 PosRF-$M$ PosRF-QMC PosRF-SGQ PosRF-$M$ PosRF-$M$ PosRF-$M$
+128 PosRF-SORF PosRF-QMC PosRF-QMC PosRF-QMC PosRF-SORF PosRF-QMC
+256 PosRF-$M$ PosRF-SORF PosRF-$M$ PosRF-SGQ PosRF-QMC PosRF-$M$
+512 PosRF-FastFoodL
+Memory 64 PosRF-$M$
+128 PosRF-FastFoodL
+256 PosRF-FastFoodL
+512 PosRF-FastFoodL
+4.3. Task Characteristics
+The tasks in the LRA include: ListOps (Nangia and Bowman, 2018), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard and Ruder, 2018), Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013), Image (Image Classification), and Pathfinder (Long-Range Spatial Dependency) (Linsley et al., 2018; Kim et al., [n.
+d.]).
 We take into consideration three main characteristics:
@@
 Retrieval inputs two sequences and the output is either False (meaning the sequences are not of the same document) or True (meaning the sequences are of the same document).
-  • Maximum sequence size. ListOps is 2K; Text is 4K; whilst Retrieval requires the processing of two sequences each of 4K length making the total maximum sequence size being 8K.
+Image is based on the CIFAR-10  (Krizhevsky et al., 2009) and has 10 classes.
+Pathfinder (Linsley et al., 2018; Kim et al., [n.
+d.]) has 2 classes and requires the model to identify the target contour between one end and the other.
+  • Maximum sequence size.
+ListOps is 2K; Text is 4K; whilst Retrieval requires the processing of two sequences each of 4K length making the total maximum sequence size being 8K; Image flattens the 2D of CIFAR-10 images of 32-by-32 into a list with the size of 1024; Similary, Pathfinder also takes in a 1024-dimensional list from 32-by-32 images.
   • Required attention span.
 This is the ‘mean distance between the query token and the attended tokens, scaled by attention weights’.
-The average values for required attention span calculated by Tay et al.
-(2021) is: ListOps being 0.8, Text being 0.3, and Retrieval being 1.3.
-Table 4: Mean Pearson correlation between model performance statistics and each task characteristic across weight matrices and component functions.
-Model performance statistics include ‘acc’ (accuracy), ‘time’ (training time), ‘mem’ (peak memory consumption).
-Task characteristics are ‘size’ (maximum sequence size), ‘span’ (required attention span), and ‘classes’ (number of output classes).
-Correlation is calculated for each weight matrix or component function and the average value over weight matrix or component function respectively displayed.
-Left is weight matrix $W$, right is component function $f$.
-$W$ acc time mem
-size 0.93 0.74 0.64
-span 0.37 0.02 -0.10
-classes -0.91 -0.96 -0.94
-$f$ acc time mem
-size 0.93 0.56 0.60
-span 0.36 0.02 -0.10
-classes -0.91 -0.72 -0.89
-We then calculate Pearson correlation of accuracy, training time, and memory consumption against these characteristics.
-Our result is shown in Table 4.
-The table indicates that: (1) contrary to intuition, longer sequence size gives higher accuracy but this is not because of increasing required attention span and they somewhat increase training time and peak memory consumption; (2) increasing the number of classes decreases accuracy, training time, and peak memory consumption.
-4.3 Analysis of Combinations
-Table 5: Mean statistics across weight matrices for each task.
-Task is denoted on the top left cell of each table.
-Mean statistics include ‘acc’ (accuracy (%)), ‘time’ (training time (hour)), and ‘mem’ (peak memory consumption (GB)).
-ListOps acc time mem
-SORF 27.49 0.64 1.32
-SGQ 34.28 0.64 1.32
-FastFoodL 35.68 1.05 0.86
-ORF 36.59 0.64 1.32
-QMC 37.39 0.64 1.32
-MM 37.41 0.64 1.32
-Text acc time mem
-ORF 60.28 1.19 2.63
-QMC 60.92 1.19 2.63
-MM 60.98 1.19 2.63
-SGQ 62.03 1.19 2.63
-SORF 64.15 1.18 2.63
-FastFoodL 64.56 2.04 1.71
-Retrieval acc time mem
-SORF 72.58 1.19 2.45
-FastFoodL 77.22 2.10 1.67
-SGQ 78.45 1.19 2.45
-QMC 80.59 1.20 2.45
-ORF 80.79 1.20 2.45
-MM 80.93 1.20 2.45
+The average values for required attention span calculated by (Tay et al., 2021) is: ListOps being 0.7589, Text being 0.3194, Retrieval being 1.329, Image being 0.359 Pathfinder being 0.5371.
+4.4. Analysis of Methods
+Figure 4.
+Heatmaps of Spectraformer combinations for the 64, 128, 256, 512 features.
+$L$ and II have 10 classes, opposed to $T$, $R$, $P$ having 2 classes.
+Hence we adjust the vmin-vmax for the first group as 20-40 with the second group as 50-85.
+Table 5.
+Pearson correlation coefficients between model performance metrics (Mtr.) (Accuracy (Acc), Time (Time), Memory (Mem)) and task characteristics.
+Task characteristics include the number of classes (Cl.), input sequence length (Len.), and required attention span (Att.).
+The analysis is presented for component functions (left) and weight matrices (right).
+$f$ $W$
+Mtr. T PosRF OPRF SADERF FastFoodL QMC $M$ ORF SGQ SORF
+Acc Cl. -0.908 -0.924 -0.921 -0.950 -0.921 -0.919 -0.921 -0.905 -0.913
+Len. 0.706 0.720 0.722 0.664 0.735 0.741 0.728 0.754 0.696
+Att. 0.459 0.516 0.494 0.404 0.566 0.562 0.557 0.506 0.356
+Time Cl. 0.169 0.191 0.185 0.234 0.176 0.177 0.165 0.177 0.176
+Len. -0.169 -0.198 -0.186 -0.215 -0.180 -0.180 -0.175 -0.184 -0.180
+Att. -0.271 -0.275 -0.265 -0.364 -0.255 -0.255 -0.252 -0.259 -0.255
+Mem Cl. 0.160 0.144 0.143 0.241 0.160 0.160 0.149 0.161 0.160
+Len. -0.185 -0.157 -0.159 -0.253 -0.177 -0.177 -0.171 -0.181 -0.177
+Att. -0.265 -0.233 -0.238 -0.399 -0.259 -0.259 -0.255 -0.263 -0.259
 We now explore the suitability of weight matrices and component functions depending on task characteristics.
-Table 5 shows the mean statistics across weight matrices for each task.
-We notice that weight matrices form two different groups which align with the required attention span differences.
-ListOps and Retrieval are tasks with high required attention span, ListOps being 0.8 and Retrieval being 1.3.
-In both these tasks, we observe that ORF, QMC, and MM perform comparatively against each other and a lot better against SORF, SGQ, and FastFoodL.
-However, the Text task, which has a much lower required attention span of 0.3, has much higher performance from the SORF-SGQ-FastFoodL group, whilst ORF, QMC, and MM perform much poorly on this task.
-Based on this observation, we can conclude that ORF, QMC, and MM are suited for high required attention span; whilst SORF, SGQ, and FastFoodL are suited for low required attention span.
-Analyzing mean statistics per task across component functions do not show any task-specific trend (refer to Table 9 for more information).
-Table 6: Summary statistics of component functions $f$ (top) and weight matrices $W$ (bottom) in mean accuracy (on the test set), mean time (training time (hour)), and mean memory (peak memory consumption (GB)).
-$f$ and $W$ entries are sorted by mean accuracy respectively.
-Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
-$f$ SADERF 59.17 1.19 2.19
-$f$ OPRF 59.02 1.18 2.08
-$f$ PosRF 57.20 1.02 1.77
-$W$ MM 59.78 1.01 2.13
-$W$ QMC 59.63 1.01 2.13
-$W$ ORF 59.22 1.01 2.13
-$W$ FastFoodL 59.15 1.73 1.41
-$W$ SGQ 58.25 1.01 2.13
-$W$ SORF 54.74 1.00 2.13
-In analyzing efficiency, we follow the definition of efficient Transformers in Tay et al.
-(2022) in terms of memory and computation, the latter interpreted as computational cost.
-The overall trend of accuracy, time, and memory of component functions is that they increase linearly from PosRF, OPRF, to SADERF (see Table 6).
+Table 8 shows the mean statistics across component functions and weight matrices for each task as well as average performance across tasks.
+To ensure the performance of methods consistently across different feature settings, we also refer to Figure 4.
+First, we will analyze the component functions and weight matrices in terms of accuracy.
+For component functions, we can see that OPRF and SADERF outperform PosRF for all tasks except Text.
+For weight matrices, we observe the following:
+  • $T$: FastFoodL and SORF performs well in $T$ but other models also perform equally well in task $T$. This task is not too hard.
+  • $R$: ORF, QMC, and MM perform the best. FastFoodL and SGQ perform well moderately, and SORF performs the worst.
+  • $P$: Most models perform well (ORF, FastFoodL, QMC, MM). SGQ slightly less well than this group. SORF performs quite poorly.
+  • $L$: QMC and MM perform the best. ORF performs moderately well. SORF performs the worst. Meanwhile, SGQ’s performance varies with respect to the number of random features.
+  • II: FastFoodL performs the best, followed by SGQ. ORF, QMC and MM perform moderately. With SORF performing the worst.
+Based on the above observations, we note that ORF, QMC, and MM tend to have the same trend.
+These weight matrices along FastFoodL (albeit having a different trend) have high mean performance (see Table 8) and high performance per task.
+This is followed by SGQ.
+SORF has the worst performance.
+QMC and MM having the same trend is expected since MM is an improvement on QMC using the same base formula and assumption.
+Empirically, QMC also outperforms ORF marginally as expected, based on existing work (see (Liu et al., 2022), Table 9 (Appendix) w.r.t.
+Gaussian training error).
+QMC and MM are based on Equation 3.1.5 which is a reformulation of Equation 9 whilst reducing variance.
+Thus, theoretically QMC and MM should behave similarly yet better than MC-based methods like ORF.
+SORF is the faster version of ORF (Yu et al., 2016), and empirically it becomes nearly unbiased with the number of features above 32, essentially collapsing to the bias of ORF.
+Yet in practice, Liu et al.
+(2022) show that SORF does not perform as well as ORF in classification tasks, and this is reflected in our experiments.
+In addition, Liu et al.
+(2022) also show that FastFoodF (which is the FastFood approximator (FastFood fixed) as opposed to FastFoodL, the FastFood learner) performs as well as QMC.
+Based on (Chowdhury et al., 2022) however, it is expected that FastFoodL exceeds this.
+And being the weight learner class, its behaviors deviate from the general trend of weight matrices.
+Liu et al.
+(2022) show empirically that GQ performs as well as QMC.
+However, GQ cannot be used in practice, hence the use of SGQ.
+It was also not compared empirically in (Liu et al., 2022).
+Given that SGQ approximates a kernel $k$ that is subgaussian with parameter $b$, $A\geq 24eb^{2}M^{2}$, $d$ being the dimension of the data, $M\subset\mathbf{R}^{d}$ be the diameter between inputs $x$ and $y$, then SGQ’s sampling complexity is bounded by $2^{d}\left(\frac{12eb^{2}M^{2}}{A}\right)^{A}$.
+This means that the dimension of the token as well as the distance between two tokens being calculated will increase the complexity.
+Since our dimension of the token is quite high compared for most kernel tasks (64), and the distance between two tokens is quite high as well due to the tasks being from the LRA, this results in SGQ’s rather average performance.
+The relationship between the distance of the tokens and the approximation quality of SGQ is apparent in the moderate Pearson correlation coefficient of 0.506 between SGQ and attention span in Table 5.
+Unfortunately, we have less than an obvious idea for the relationship between the other weight matrices and token characteristics like SGQ since most of their sampling complexity is only dependent on the number of features (Liu et al., 2022).
+Table 5 shows that, as expected, the number of classes is the primary determinant of models’ performance.
+However, the input length also plays a role in determining the outcome of the performance of a weight matrix in a task as well.
+We observe a similar trend across ORF, QMC, and MM as stated above across correlation coefficients between accuracy w.r.t.
+number of classes, input length, and attention span.
+SGQ as suggested above performs similarly for the first two metrics but with a slightly lower attention span coefficient.
+FastFood has the lowest correlation coefficients for both input length and attention span, suggesting that its learnable weights behave contrary to task characteristics and learning might be attributed to this.
+Figure 5 shows the accuracy trend for different number of features for each task.
+We observe that accuracy peaks mostly at 128 and sometimes 256 number of features, thereafter maintaining or decreasingly stably.
+We also find that there is no relationship between input length and the number of features (also see Figure 6).
+This suggests that increasing the number of random features past 128 induces overfitting and 128 is the ideal number of random features.
+The best way to confirm whether overfitting is occurring is to look at the gap between the generalization accuracy (the accuracy on the development set) and the training accuracy.
+We term this accuracy as the generalization gap.
+We calculate the average generalization gap and averaging them for all seeds of the same task and model between 512 - 256, 256 - 128, and 128 - 64 features.
+The result is respectively 0.143%, 0.041%, -0.404%.
+This confirms our initial intuition that 64 features are too few (underfitting) but 256 features are too many (overfitting).
+The accuracy obtained is obtained with early stopping, accuracy obtained without early stopping can induce a much larger generalization gap difference.
+This confirms that overfitting occurs past 128 features.
+In addition, during experimentation, we also found that the 512 features is highly unstable due to gradient explosion from large matrix calculation leading to NaN values.
+This occurred for three seed (twice for PosRF-SGQ and once for SAERF-ORF).
+SORF in particular performs poorly at 64 number of features with accuracy improving drastically at 128 number of features.
+This validates previous findings that in transformer applications, 128 features is the optimal option.
+Figure 5. Accuracy vs Number of Features Across All Tasks
+4.5. Efficiency Analysis
+Figure 6. Performance Metrics vs. Number of Features
+In analyzing efficiency, we follow the definition of efficient Transformers in (Tay et al., 2022) in terms of memory and computation, the latter interpreted as computational cost.
+The overall trend of accuracy, time, and memory of component functions is that they increase linearly from PosRF, OPRF, to SADERF.
+Mean accuracy is respectively 53.71, 55.97 and 55.46.
+Mean time is respectively 1.32, 1.53 and 1.54.
+Mean memory is 2.31, 2.70, 2.86.
+Please refer to Table 8.
+These figures are for 128 random features but with other number of features, the results follow the same trend.
+SADERF and OPRF having a higher computational requirement than PosRF is due to them having additional computation of the L2 norm of the weight matrix which PosRF does not calculate.
 Hence, when limited computational resource is a concern, PosRF is a better option.
+Figure 6 shows that PosRF’s training tmie and memory consumption grows at a significantly lower rate than other methods.
+Otherwise, OPRF provides the best accuracy.
+We observe the increase in training time and peak memory consumption as expected due to increasing the number of random features sampled for weight matrices (see Figure 6).
 Training time and peak memory consumption relies more heavily on component functions than weight matrices, with the exception of FastFoodL.
 FastFoodL takes significantly less memory consumption at the expense of training time.
-The accuracy trend for weight matrices is as described above.
-Table 6 also shows that whilst FastFoodL generally does not work as well due to its low required attention span, its high performance in the task of Text which is low in required attention span and the anomalously performant OPRF-FastFoodL leads to a data skew in the final ranking in Table 3.
-Comparing between combinations, we find that for ListOps, OPRF-ORF is the most accurate (38.34%), followed by OPRF-MM (38.08%) and OPRF-QMC (37.69%).
-OPRF-ORF, being the previous SOTA, is undoubtedly the best choice here since other models maintain the same training time and peak memory consumption with lower accuracy.
-For Text, the most accurate model is OPRF-SORF (64.76%), followed by SADERF-SORF (64.70%) and PosRF-FastFoodL (64.65%).
-These all outperform the previous SOTAs (OPRF-ORF and SADERF-ORF) by $4\%+$ whilst maintaining a diversity of training time and peak memory consumption options.
-Both OPRF-SORF and SADERF-SORF maintain the same level of training time and peak memory consumption as the SOTAs, PosRF-FastFoodL offers much less peak memory consumption at the expense of longer training time.
-For Retrieval, the most accurate model is SADERF-MM (81.13%), followed by OPRF-MM (81.09%) and SADERF-ORF (81.05%).
-Whilst these models only perform slightly better than the SOTA SADERF-ORF, OPRF-MM offers less peak memory consumption compared to SADERF-ORF.
-5 Conclusion
-We present Spectraformer, a framework for approximating and learning the attention kernel in the Transformer.
-Our paper generalizes past works and presents empirical findings on different component function and weight matrix combinations.
-We experiment with 18 combinations, of which 14 are novel.
-Novel combinations perform competitively against, if not more performant than the previous SOTA in the family for LRA textual tasks.
-We show empirically that the choice of performant weight matrix depends on the required attention span of the task.
-Our results and the extensibility of our work shows the viability of Spectraformer as a unifying random feature framework of Transformer.
-Appendix A Appendix
-A.1 Future work, limitations, and broader impact
-Spectraformer provides an experimental framework for random features in the Transformer.
-However, we have not provided formal mathematical proofs for the bounds the these combinations due to the number of combinations discussed.
-We hope Spectraformer inspires future work to generalizes the proof in such a framework.
-Spectraformer has covered a significant number of combinations, however, we have not covered all weight matrices and component functions in the literature.
-Spectraformer can serve as a foundation for further random feature Transformer research using the techiques not yet experimented with but covered in Appendices A.3, A.4, A.5.
-Due to the significant number of experimented models, we are unable to perform hyper-parameter fine-tuning.
-Future work can benefit from extending the scope of the current research to other tasks, i.e., other non-textual LRA tasks (image classification, path-finder, and path-x) and short-context datasets/ tasks such as GLUE (Linzen et al., 2018).
-Transformer-based architecture can also benefit from Spectraformer.
-Past random feature Transformer works (Likhosherstov et al., 2022; 2023) have experimented with conformer and vision Transformer with ViT.
-Jeevan & Sethi (2022) provides a good foundation to extend our work.
-Spectraformer, although currently experimented in the Transformer setting, can also extend to non-parametric kernel classification.
-It could also benefit from ablation studies like the comparison variance in Chowdhury et al.
-(2022) or in Yao et al.
-(2023)
-Given the complex nature of language models in the ethical, energy, and societal domains, our work should be used responsibly (Weidinger et al., 2021).s
-A.2 Related Work
-Spectraformer is based on several theoretical assumptions and choice of mathematical framework and approach.
-This section discusses alternatives to the choices of Spectraformer.
-Specifically, Appendix A.2.1 discusses alternative kernelized formulations to Section 2.1; Appendix A.2.2 discusses alternative kernel approximation techniques to random features in Section A.2.2.
-There are also some component functions and weight matrices which are not covered in the experiments.
-We introduce them in the related work section in Appendices A.3, A.4, A.5.
-A.2.1 Alternative Kernelized Attentions
+However, whilst most methods increase training time and memory consumption significantly as the number of features increases, FastFoodL remains constant (see Figure 6).
+In addition, Table 5 shows that, in terms of weight matrices, training time and peak memory consumption is relatively stable and does not vary with respect to number of classes, input length, and the required attention span, with the exception for FastFoodL.
+This is largely due to FastFoodL modifying the attention matrix instead of approximating it.
+However, we do notice that there is a higher negative correlation between the required attention span and peak memory consumption.
+We also analyze efficiency in terms of the number of features. Figure 7 shows that memory usage decrease gradually as the sequence length increases and the training time remains relatively stable.
+4.6. Inductive Bias of Efficient Attention
+In addition, we also notice that, the top performing models (in Table 1) such as Spectraformer variants, Nystromformer (Xiong et al., 2021), Big Bird (Zaheer et al., 2020), Reformer (Kitaev et al., 2020), Skyformer (Chen et al., 2021) are efficient Transformers that outperform the OT.
+They can be grouped into the following two induction biases:
+  • Low-rank bias (Spectraformer, Nystromformer (Xiong et al., 2021), Skyformer (Chen et al., 2021)) assumes that the true attention matrix is simple, or low rank, in nature.
+  • Sparsity bias (Big Bird (Zaheer et al., 2020), Reformer (Kitaev et al., 2020)) assumes that only a small number of attention values in the attention matrix contains meaningful information, ignoring the computation of a significant number of token-to-token attention values, since they only contribute to noise.
+Efficient transformers do not just reduce the computational cost, but they also introduce strong induction bias about the structure of the attention matrix.
+Efficient transformers’ approximation of the OT’s true attention matrix should imply that their performance be upperbounded by the OT’s performance.
+However, as we have seen in Table 1, they consistently outperform the OT, implying that at least for the characteristics of the LRA, a low-rank bias or a sparsity bias, acting as regularizers on the attention matrix, are helpful induction biases.
+This aligns with the analysis done in Section 4.4.
+128 features will definitely produce a lower rank matrix than 256, 512 features hence models with an even lower-rank bias.
+This however questions why the OT outperforms 512 and 256-feature models.
+128, 256 and 512 feature models have lower rank than the OT and they necessarily produce noise due to the attention approximation process.
+Having more features likely enlarge the noise in the final approximated attention matrix, making these models fit to spurious noise.
+Having too few features (64), however, underfits and lead to inexact kernel approximation.
+4.7. Method Recommendation
+Based on the above experiments, our recommendation is to avoid SORF.
+SORF performs well on the Text task.
+However, this is not captured by our metrics (number of output classes, maximum input length, and attention span).
+There are potential task characteristics that we cannot account for with this exception.
+Overall, we suggest FastFoodL and QMC among the main weight matrix group of QMC, MM and ORF.
+They both are highly performant.
+FastFoodL offers the advantage of low memory consumption at the cost of longer training time, and QMC maintains the average training time and memory consumption.
+Our results establish OPRF-FastFoodL as the new state-of-the-art for random feature-based efficient Transformers.
+More importantly, our findings demonstrate that this class of models is now highly competitive with top-performing methods from other categories like Nystromformer (Xiong et al., 2021) and BigBird (Zaheer et al., 2020).
+Specifically, OPRF-FastFoodL achieves accuracy on par with BigBird (Zaheer et al., 2020) (57.17% vs.
+57.25%) while offering substantial reductions in both peak memory consumption (1.82 GB vs.
+4.65 GB) and training time.
+While Nystromformer attains a higher average accuracy (58.06%), our model provides a valuable alternative with greater memory efficiency (1.82 GB vs.
+2.63 GB).
+These results introduce new, valuable points to the accuracy-memory-time Pareto frontier, offering practitioners compelling options based on their specific performance and resource constraints.
+In addition to our top-performing model, the Spectraformer framework also produced other novel combinations like OPRF-SGQ and PosRF-MM, which outperform previous RF-based SOTAs (OPRF-ORF and SADERF-ORF) and provide a range of options with different computational profiles.
+The result mentioned here is available in Table 1.
+5. Related Work
+Spectraformer is based on several theoretical assumptions various mathematical frameworks.
+This section discusses alternatives to these assumptions.
+Specifically, Section 5.1 discusses alternative kernelized formulations to Section 2.1; Section 5.2 discusses alternative kernel approximation techniques to random features in Section 5.2.
+5.1. Alternative Kernelized Attentions
 In addition to the kernelized formulation of attention as shown in Section 2.1, based on our literature survey, we have identified the following three alternative kernel-based formulations of attention.
@@
 Equation 25 leads to Implicit Kernel Attention (IKA).
-$A_{i}=\sum_{j=1}^{N}\frac{1}{Z_{l}(\bm{q}_{i},\bm{K})}exp(\frac{-||\bm{q}_{i}-\bm{k}_{j}||^{2}_{2}}{2\sqrt{d_{k}}})exp(\frac{||\bm{q}_{i}||^{2}_{p=2}+||\bm{k}_{j}||^{2}_{p=2}}{2\sqrt{d_{k}}})\bm{v}_{j}$ (13)
+[…]
 Non-local operation (Wang et al., 2018) is popular decomposition of attention in computer vision.
@@
 Indeed, the attention vector output $\bm{y}$ can be defined using $f$ which is the softmax (see Equation 27) in the non-local operation.
-$\mathbf{\bm{y}}_{i}=\frac{1}{C(\bm{x})}\sum_{\forall j}f(\bm{x}_{i},\bm{x}_{j})g(\bm{x}_{j})$ (14)
-$\mathbf{\bm{y}}=softmax(\bm{x}^{T}\bm{W}_{\theta}^{T}\bm{W}_{\theta}\bm{x})g(\bm{x})$ (15)
+(26) $\mathbf{\bm{y}}_{i}=\frac{1}{C(\bm{x})}\sum_{\forall j}f(\bm{x}_{i},\bm{x}_{j})g(\bm{x}_{j})$
+(27) $\mathbf{\bm{y}}=softmax(\bm{x}^{T}\bm{W}_{\theta}^{T}\bm{W}_{\theta}\bm{x})g(\bm{x})$
 We will show that the non-local operation in Equation 26 is a generalisation of the Nadaraya-Watson kernel estimator.
@@
 The variety of kernel decompositions of attention demonstrate the robustness of the kernel interpretation of attention and hence opening up an alley in attention improvement through the lens of kernel.
-A.2.2 Kernel Approximation Beyond Random Features
-Kernel approximation is presented in Section 2.3 through the use of random features.
-However, there has been other kernel approximation techniques outside of random features.
-These include greedy basis selection (Smola & Schökopf, 2000), divide and conquer (Hsieh et al., 2014; Zhang et al., 2013), and nyström methods (Williams & Seeger, 2000) (which led to the development of Skyformer (Chen et al., 2021) and Nyströmformer (Xiong et al., 2021)).
-Figure 2: Spectraformer weight matrices.
-Red tick denotes techniques which have been explored in the literature.
-Green tick denotes techniques which are explored in this work.
-Please refer to Appendix A.3 for references.
-inline]MAYBE: Table here like table 2 comparison of different kernel space and time complexities combined with table 3 comparison of convergence rates and required random features for kernel approximation error in (Liu et al., 2022)
-A.3 Technical Details of Weight Approximators
-Spectraformer weight matrices are shown in Figure 2.
-Weight matrices are either approximators or learners.
-There are three approaches to approximators: Monte Carlo sampling, Unit-cube sampling, and quadrature.
-We have discussed these approaches in detail in Section 3.1.
-Here we provide a more specific classification schemes for each method.
-  • Monte Carlo sampling-based weight matrices are improvements over Base and subdivided into two types: structural and geometric methods.
-Structural methods decompose the Base random matrices $P$ into smaller matrices which reduce time and/ or space complexity.
-Structural matrices includes FastFoodF, SCRF and P-Model.
-Among these, we have explored the learner variant FastFoodL of FastFood.
-Geometric methods enforce certain geometrical couplings along some dimensions in the construction of random matrices to reduce the approximation error.
-Geometric methods include: ROM (ORF, SORF) and SimRF.
-Among these, we have explored the main ROM variants - ORF and SORF.
-  • Unit-cube sampling-based weight matrices reformulate the Bochner integral in Equation 9 as Equation A.3 with uniform and independent points.
-This approach includes QMC, MM, and SSF.
-We have not covered SSF in our experiments.
-  • Quadrature-based weight matrices estimates the intractable integral using non-uniform weights with deterministic rules.
-There are two approaches: SGQ and SSR.
-We have not covered SSR in our experiments.
-inline]update the equations so that we use the same dimension for the number of components and so on inline]Generative should be DGM, add Base to the graph
-We now explore these weight matrices in the order specified above.
-Base (Rahimi & Recht, 2007) is the direct sampling from a distribution $\bm{P}_{ij}\sim p$, corresponding to the kernel $\mathcal{K}$ without any additional adjustment, and in approximating the RBF, $P=G$ where $G$ is the normal distribution.
-$W_{Base}=\frac{1}{\sigma}\bm{P}$ (16)
-FastFoodF (Le et al., 2013) makes use of Hadamard and diagonal matrices to speed up Gaussian matrices $G$ constructions for RFF in $O(n\log{d})$ (100 times faster) with $O(n)$ space (1000 times less space) given $n\geq d$ against Base.
-However, FastFoodF increases variance and approximation error and decreases the concentration bound.
-$W_{FastFood_{L}}=\frac{1}{\sigma}\bm{SHG\Pi HB}$ (17)
-[…]
-SCRF (Signed Circulant Matrix Projection) (Feng et al., 2015) maintains a high matrix construction speed and $O(s)$ space ($s=t*d$) whilst maintaining the variance of Base, an improvement over FastFoodF.
-It does this by making use of circulant matrices.
-$\bm{W}_{SCRF}=[\bm{P}^{(1)},\bm{P}^{(2)},...,\bm{P}^{(t)}],t=d|D$ (18)
-where $\bm{P}^{(i)}$ is a signed circulant Gaussian matrix, with the first column of circulant random matrix $\bm{C}^{(i)_{[d]}}$ of $\bm{P}^{(i)}$ drawn randomly from a Gaussian distribution $\mathcal{N}(\bm{0},\bm{I}/\sigma^{2})$.
-The circulant matrix via Discrete Fourier Transform can be defined as:
-$\begin{array}[]{cc}\bm{C}=\frac{1}{m}\bm{F}^{*}diag(\bm{Fc})\bm{F}\\ \bm{F}=[e^{i\frac{2\pi}{m}kn}]^{m-1}_{k,n=0}\end{array}$ (19)
-P-Model (Choromanski & Sindhwani, 2016) generalizes over both circulant matrices (including SCRF), FastFoodF, along with other Gaussian and semi-Gaussian structured matrix methods.
-$\bm{W}_{P_{model}}=[\bm{g^{T}P_{1}},\bm{g^{T}P_{2}},...,\bm{g^{T}P_{s}}]^{T}$ (20)
-where $\bm{g}\in\mathbb{R}^{a}$ is a Gaussian vector and a sequence of $m$ $\bm{P}_{i}\in\mathbb{R}^{a*d}$ matrices with L2 norm columns, where $P_{i}$ represents a P-model.
-Semi-Gaussian structured matrices including Toeplitz, Hankel, and other Toeplitz-like matrices are an open area of research for random matrices.
-ORF (Orthogonal RF) (Yu et al., 2016) decreases the approximation error significantly compared to Base.
-This is done via replacing $G$ with a properly scaled random orthogonal matrix.
-However, generating orthogonal matrices become costly quickly as the number of dimensions increases.
-$\bm{W}_{ORF}=\frac{1}{\sigma}\bm{SQ}$ (21)
-where $\bm{Q}$ is a uniformly distributed random orthogonal matrix (on the Stiefel manifold) obtained from the QR decomposition of $\bm{G}$, the set of rows of $\bm{Q}$ forming a basis in $\mathbb{R}^{d}$, and $\bm{S}$ is a diagonal matrix with entries sampled i.i.d from the $\chi$-distribution with $d$ degrees of freedom, thus making the rows of $\bm{SQ}$ and $\bm{G}$ identically distributed.
-SORF (Structured ORF) (Yu et al., 2016) decreases time and space complexity of ORF (from $O(d^{2})$ to $O(d\log{d})$ with almost no extra memory cost) by imposing structure on the orthogonal matrices, inspired by structural methods.
-SORF is unbiased with large $d$.
-We replace $\bm{S}$ with $\sqrt{d}$ and $\bm{Q}$ with structured matrix $\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$
-$\bm{W}_{SORF}=\frac{\sqrt{d}}{\sigma}\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$ (22)
-where $\bm{D}_{i}\in\mathbb{R}^{d*d},i=1,2,3$ is diagonal ’sign-flipping‘ matrices with diagonal entries sampled from the Rademacher distribution, $\bm{H}$ is the normalized Walsh-Hadamard matrix.
-ROM (Random Orthogonal Embeddings) (Choromanski et al., 2017) generalizes SORF ($t=3$) to:
-$\bm{W}_{ROM}=\frac{\sqrt{d}}{\sigma}\prod_{i=1}^{t}\bm{SD}_{i}$ (23)
-where $\bm{S}$ is a class of $L_{2}$-normalized versions of Kronecker product matrices, of which the Hadamard $\bm{H}$ is a representative of, and $\bm{D}$ being independent diagonal matrices.
-When $\bm{S}=\bm{H}$ and $t=3$, we obtain SORF used in our experiments.
-Another ROM variant is S-Hybrid, which we do not experiment with, given as
-$\bm{W}_{S_{Hybrid}}=\bm{SD}_{k}^{(\mathcal{U})}\prod_{i=1}^{k-1}\bm{SD}_{i}^{(\mathcal{R})}$ (24)
-where $\bm{D}_{k}^{(\mathcal{U})}$ is a diagonal matrix with i.i.d.
-$Unif(\mathcal{S}^{1})$ on the unit circle of $\mathbb{C}$.
-When being used in the component function, instead of calculating $\phi(\bm{x})^{T}\phi(\bm{y})$ in Equation 10.
-We model the equation as $Real(\overline{\phi(\bm{x})}^{T}\phi(\bm{y}))$.
-SimRF (Simplex RF) (Reid et al., 2023) is developed to be a positive RF with the most optimal solution, the MSE of which is the lowest in the geometric methods.
-This is an alternative to the ROM-based geometric methods.
-Two variants of SimRF are: weight-independent and weight-dependent.
-SimRF is defined as:
-$\bm{W}_{simp}=\bm{DSR}$ (25)
-where $\bm{D}\in\mathbb{R}^{d*d}=diag(\bm{\omega}_{i}),\bm{R}\in\mathbb{R}^{d*d}$ being a random orthogonal matrix drawn from Haar measure on $O(d)$ and $\bm{S}\in\mathbb{R}^{d*d}$ with row $\bm{s}_{i}$:
-$\bm{s}_{i}=\begin{cases}\sqrt{\frac{d}{d-1}}\bm{e}_{i}-\frac{\sqrt{d}+1}{(d-1)^{3/2}}(1,...,1,0)^{T}&1\leq i<d\\ \frac{1}{\sqrt{d-1}}(1,1,...,1,0)^{T}&i=d\end{cases}$ (26)
-being unit vectors which are manifestly normalized and subtend obtuse angles.
-SimRF variants are defined based on how $\bm{\omega}_{i}$ is constructed.
-For SimRFindep, $\bm{\omega}_{i}\sim\chi_{d}$.
-For Simdep, we permit the random vector direction $\{\hat{\bm{\omega}}_{i}\}$ to be correlated with norms $\{\bm{\omega}_{i}\}$:
-$\bm{\omega}_{i}=-\frac{\sum_{j\neq i}\omega_{j}}{||\sum_{j\neq i}\omega_{j}||_{2}}\bm{\omega}_{i},i=1,...,d$ (27)
-QMC (Quasi-Monte Carlo) (Avron et al., 2016) evaluates on a low discrepancy sequence (e.g., Halton, Sobol’ Faure, and Niederreiter) of points instead of random points in Monte Carlo.
-Although the approximation error is only reduced minimally, QMC has been shown to perform better than MC in high dimensions and does not have undesirable clustering effect.
-To calculate QMC, we first assume that $p(\bm{x})=\prod^{d}_{j=1}p_{j}(\bm{x}_{j})$ factorizes with respect to the dimensions with $p_{j}(.)$ being a univariate density function.
-Then we define:
-$\Phi^{-1}(\bm{t})=\left(\Phi_{1}^{-1}\left(\bm{t}_{1}\right),\ldots,\Phi_{d}^{-1}\left(\bm{t}_{d}\right)\right)\in\mathbb{R}^{d}$ (28)
-where $\Phi_{j}$ being the cumulative distribution function of $p_{j}$, $\bm{t}_{1},\bm{t}_{2},...,\bm{t}_{s}\in[0,1]^{d}$ being a low discrepancy sequence, and $\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$.
-We can thus transform the integral on $\mathbb{R}^{d}$ in Equation 9 to an integral on the unit cube $[0,1]^{d}$ as
-$\mathcal{K}(\bm{x}-\bm{x}^{\prime})=\int_{[0,1]^{d}}\exp\big{(}\mbox{i}(\bm{x}-\bm{x}^{\prime})^{\!\top}\Phi^{-1}(\bm{t})\big{)}\mbox{d}\bm{t}$
-The weight matrix then can be defined as:
-$\bm{W}_{\text{QMC}}=[\Phi^{-1}(\bm{t}_{1}),\Phi^{-1}(\bm{t}_{2}),\ldots,\Phi^{-1}(\bm{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}\;.$
-QMC can be further improved by a sub-grouped based rank-one lattice construction which improved complexity (Lyu et al., 2020).
-MM (Moment Matching) (Shen et al., 2017; Liu et al., 2022) improves over QMC by removing the undesirable clustering effect and having the same approximation error with less features.
-This is done by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$:
-$\bm{W}_{\text{MM}}=[\widetilde{\Phi}^{-1}(\bm{t}_{1}),\widetilde{\Phi}^{-1}(\bm{t}_{2}),\ldots,\widetilde{\Phi}^{-1}(\bm{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}$
-where $\widetilde{\Phi}^{-1}(\bm{t}_{i})=\tilde{\bm{A}}^{-1}(\Phi^{-1}(\bm{t}_{i})-\tilde{\bm{\mu}})$ can be constructed using moment matching with sample mean $\tilde{\bm{\mu}}=\frac{1}{s}\sum_{i=1}^{s}\Phi^{-1}(\bm{t}_{i})$ and the square root of the sample covariance matrix $\tilde{\bm{A}}$ satisfying $\tilde{\bm{A}}\tilde{\bm{A}}^{\!\top}=\text{Cov}(\Phi^{-1}(\bm{t}_{i})-\tilde{\bm{\mu}})$.
-SSF (Spherical Structured Features) (Lyu, 2017) improves over QMC by including rotation-invariant kernel as well as in terms of complexity.
-Rotation-invariant property suggests the construction of feature maps using spherical equal weight approximation using Riesz s-energy on a d-dimensional sphere $\mathbb{S}^{d}:=\{\bm{x}\in\mathbb{R}^{d+1}|||\bm{x}||_{2}=1\}$.
-Specifically, we construct $\{\bm{v}_{i}\}^{s}_{i=1}$ asymptotically uniformly distributed on the sphere and obtain $\bm{V}:=[\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{s}]\in\mathbb{S}^{(d\times s}$, that $\Phi^{-1}(t)$ uses the one-dimensional QMC point, we have the weight matrix:
-$\bm{W}_{\text{SSF}}=[\Phi^{-1}(t)\bm{v}_{1},\Phi^{-1}(t)\bm{v}_{2},\ldots,\Phi^{-1}(t)\bm{v}_{s}]^{\!\top}\in\mathbb{R}^{s\times(d+1)},$
-SGQ (Sparse Grid Quadrature) (Dao et al., 2017) evolves from GQ.
-GQ (Gaussian Quadrature) (Dao et al., 2017) assumes that the kernel $k$ factorizes with respect to the dimensions and thus can be approximated by a one-dimensional Gaussian quadrature rule.
-However the total number of points $s$ scale exponentially with the dimensions.
-However GQ suffers form the curse of dimensionality.
-This is alleviated using Smolyak rule resulting in SGQ.
-Assuming the third-degree SGQ using symmetric univariate quadrature points $\left\{-\hat{p}_{1},0,\hat{p}_{1}\right\}$ with weights $(\hat{a}_{1},\hat{a}_{0},\hat{a}_{1})$, then we have:
-$\bm{W}_{\text{SGQ}}\!=\![\bm{0}_{d},\hat{p}_{1}\bm{e}_{1},\ldots,\hat{p}_{1}\bm{e}_{d},-\hat{p}_{1}\bm{e}_{1},\ldots,-\hat{p}_{1}\bm{e}_{d}]^{\!\top}\!\in\!\mathbb{R}^{(2d+1)\times d}\!$
-given that $\mathbb{e}_{i}$ is the d-dimensional standard basis vector with the ith element being 1.
-SSR (Spherical-Radial Rules) (Munkhoeva et al., 2018) transforms Equation 10 into a double integral over a hyper-sphere and the real line. This leads to the following approximation:
-$\bm{W}_{\text{SSR}}=\bm{\vartheta}\otimes\begin{bmatrix}(\bm{QV})^{\top}\\ -(\bm{Q}\bm{V})^{\top}\end{bmatrix}\in\mathbb{R}^{(2d+1)\times d},$
-where $\bm{\vartheta}=[\vartheta_{1},\vartheta_{2},\ldots,\vartheta_{s}]$, $\bm{V}=[\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{d+1}]$, $\vartheta\sim\chi(d+2)$ and $\{\bm{v}_{i}\}_{i=1}^{d+1}$ being vertices of a unit regular d-simplex randomly rotated by $\bm{Q}$ (a random orthogonal matrix).
-A.4 Technical Details of Weight Learners
-We give a non-mathematical definition of weight matrix learner in Section 3.1: weight matrix learner is any function which parameterize distributions.
-This allows for the discovery and future experimentation with other potential weight matrix learners in Spectraformer and random feature method in general.
-Weight matrix learner comes from a long line of work in kernel methods, namely, kernel learning via random features.
-Typically, a component function $f$ is chosen and a weight matrix $\bm{W}$ is produced via kernel approximation, providing us with the feature map $\phi$.
-$\phi$ can then be combined with input $\bm{x}$ and weight $\bm{v}$ to fit the learning objective.
-Kernel learning, however, imposes an additional objective: learning the weight matrix $\bm{W}$.
-The approach to the dual objectives separates the kernel learning methods into two-stage and one-stage methods (Liu et al., 2022).
-Two-stage methods (e.g., Yu et al.
-(2015); Wilson & Adams (2013); Bullins et al.
-(2018)) solves the dual objective separately: $\bm{W}$ is typically first learned via a kernel alignment scheme (Wang et al., 2015), then we solve for $\bm{v}$.
-One-stage methods (e.g., Chowdhury et al.
-(2022); Tompkins et al.
-(2019)), on the other hand, solves the dual objective in parallel: $p(.)$ corresponding to $\bm{W}$ is parameterized then $\bm{v}$ is solved typically.
-It is not plausible to apply kernel alignment (Wang et al., 2015) on Transformer architectures.
-Therefore, only one-stage methods can be considered in the context of weight matrix learner.
-Two kernel learning methods, GMM and FastFoodL, have been experimented in the non-Transformer setting (Yang et al., 2015), then adapted into the Transformer setting (Chowdhury et al., 2022).
-DGM (Deep Generative Model) has also been effective in modeling distributions (Kingma & Welling, 2013).
-It too has been adapted by Chowdhury et al.
-(2022) for the Transformer setting.
-Due to the superior performance of FastFoodL with this technique, we have only experimented with this technique.
-We now explore them.
-FastFoodL (see Equation 17) is proposed to be learnable by Yang et al.
-(2015) with two variations: FSARD (scaling matrix $\bm{S}$ previously sampled from chi-squared is made learnable) and FSGBARD (optimizes the marginal likelihood with respect to the diagonal matrices $\bm{G}$ and $\bm{B}$).
-It is further adapted by Chowdhury et al.
-(2022) to make $\bm{S}$ or $\bm{S},\bm{G},\bm{B}$ learnable parameters.
-In our experiments, we maintain this set up.
-GMM (Gaussian Mixture Model) (Wilson & Adams, 2013; Oliva et al., 2016) is a known universal approximator of probability distribution.
-Specifically, we can consider $\bm{W}_{GMM}$ being composed of $s$ components, with each component $c$ (with $C$ total components) being a Gaussian with mean vector $\mu_{c}$ and covariance matrix $\Sigma_{c}$ and component weight $\pi_{c}$.
-$p(\bm{\omega})$ can then be approximated as $p(\bm{\omega})=\sum_{c=1}^{C}\pi_{c}\mathcal{N}(\mu_{c},\Sigma_{c})$ (Chowdhury et al., 2022).
-Assuming $\pi_{c}=\frac{1}{C}$, the weight matrix with learnable weights $\Sigma$ and $\mu$ can be parameterized as:
-$\bm{W}_{GMM}=[\Sigma_{1}n+\mu_{1},...,\Sigma_{C}n+\mu_{C}]^{T},n\sim\mathcal{N}(\bm{0},\bm{I})$ (29)
-Another formulation of GMM is proposed in Yang et al. (2015) in combination with FastFood, albeit it extends beyond constructing the weight matrix.
-DGM (Deep Generative Model) (Kingma & Welling, 2013) allows for a generic distribution modeling using neural network.
-DGM is initially proposed to make differentiating sampling processes possible via the reparametrization trick, where the generator $g$ creates samples from the target distribution $\bm{\omega}_{m}=g(\bm{n}_{m})$ via transforming from a noise distribution $\bm{n}_{m}\sim p_{0}(.)$.
-In addition to the techniques discussed above, there are also some other prominent one-stage methods (e.g., INN, BBQ).
-INN (Invertible Neural Network) (Xu et al., 2021) is another universal approximator of probability distribution, consisting of invertible operations transforming samples from a known distribution to a different more complex distribution.
-BBQ (Black Box Quantile) (Tompkins et al., 2019) follows the QMC scheme and models $\Phi^{-1}$ using parameterized quantile function.
-A.5 Technical Details of Component Functions
-Figure 3: Spectraformer component functions.
-Red tick denotes techniques which have been explored in the literature.
-Green tick denotes techniques which are explored in this work.
-Please refer to Appendix A.5 for references.
-Component functions discussed in Section 3.2 all belong to CDERF.
-However, there are other component functions including DIRF and HRF (see Figure 3) which we do not explore in our experiments.
-We now investigate the theoretical details of these component functions, the CDERF family is introduced in their order of discovery.
-CDERF (Complex dense exponential random feature) is given in Equation 30, where $\bm{\omega},\bm{x}\in\mathbb{R}^{d}$, $p=\mathcal{N}(0,1)^{d}$, $k\in\{1,2\}$ and $\mathbb{S}_{d_{c}}$ being a set of $d*d$ complex symmetric matrices.
-The parameters and constraints of different CDERF functions are specified in Table 7:
-$f_{DE}^{(k)}(\bm{\omega},\bm{x})=Real(D\exp(\bm{\omega}^{T}\bm{A}\bm{\omega}+\bm{\omega}^{T}\bm{B}^{(k)}\bm{x}+\bm{x}^{T}\bm{C}^{(k)}\bm{x}))$ (30)
-Table 7: CDERF component functions
-$\bm{A}$ $\bm{B}^{(1)}$ $\bm{B}^{(2)}$ $\bm{C}^{(1)}$ $\bm{C}^{(2)}$ $D$ $\bm{x}$ $\bm{y}$
-[…]
-[…]
-SADERF (Likhosherstov et al., 2023) $A_{GE}\bm{I}_{d}$ $B_{GE}\bm{I}_{d}$ $C_{GE}\bm{I}_{d}$ $D_{GE}$ $\Psi\bm{x}$ $\Psi^{-1}\bm{y}$
-GERF (Likhosherstov et al., 2022) $A_{GE}\bm{I}_{d}$ $B_{GE}^{(1)}\bm{I}_{d}$ $B_{GE}^{(2)}\bm{I}_{d}$ $C_{GE}\bm{I}_{d}$ $D_{GE}$ $\bm{x}$ $\bm{y}$
-TrigRF (Rahimi & Recht, 2007) $0\bm{I}_{d}$ $i\bm{I}_{d}$ $-i\bm{I}_{d}$ $0\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
-PosRF (Choromanski et al., 2021) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $-1\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
-OPRF (Likhosherstov et al., 2022) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $-1\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
-TrigRF (Trigonometric RF) (Rahimi & Recht, 2007) is the base RFF implementation.
-$\begin{array}[]{c}f_{TrigRF}^{(1)}(\bm{\omega},\bm{x})=\sqrt{2}\cos(\bm{\omega}^{T}\bm{x}+\bm{b})\\ f_{TrigRF}^{(2)}(\bm{\omega},\bm{y})=\sqrt{2}\sin(\bm{\omega}^{T}\bm{y}+\bm{b})\\ \bm{b}\sim Uniform(0,2\pi)\end{array}$ (31)
-The use of the trigonometric sine and cosine functions leads to unstable behavior when the inputs have negative dimension-values.
-This can be further exacerbated when the values TrigRF try to approximate are close to 0 (since most values are of low significance).
-This causes the variance to approach infinity (Choromanski et al., 2021; Likhosherstov et al., 2022).
-Therefore, we do not want to use TrigRF to perform kernel approximation in attention.
-PosRF (Positive RF) (Choromanski et al., 2021) fixes the problem of TrigRF by enforcing positive component function output in the softmax.
-The variance of PosRF (in contrast to the variance of TrigRF approaching infinity) approaches 0 as the approximated value of the Softmax kernel approaches 0.
-PosRF has two forms: Positive RF Base (PosRF-B) (see Equation 32) and Positive RF Hyperbolic (PosRF-Hyp), which is multi-component, i.e., Equation 11: $l=2$, (see Equation 33).
-$\begin{array}[]{c}f_{PosRF_{B}}(\bm{\omega},\bm{x})=\exp(\bm{\omega}^{T}\bm{x}-\frac{||\bm{x}||^{2}}{2})\end{array}$ (32)
-$\begin{array}[]{ll}f_{1,PosRF_{Hyp}}(\bm{\omega},\bm{x})=\exp(\bm{\omega}^{T}\bm{x}-||\bm{x}||^{2})\\ f_{2,PosRF_{Hyp}}(\bm{\omega},\bm{x})=\exp(-\bm{\omega}^{T}\bm{x}-||\bm{x}||^{2})\end{array}$ (33)
-We only use PosRF-B in our experiments.
-GERF (Generalized exponential RF) (Likhosherstov et al., 2022) generalizes both TrigRF and PosRF with Equation 34.
-$\begin{array}[]{c}f_{GERF}^{(1)}(\bm{\omega},\bm{x})=D\exp(A||\bm{\omega}||^{2}+B\bm{\omega}^{T}\bm{x}+\bm{C}||\bm{x}||^{2})\\ f_{GERF}^{(2)}(\bm{\omega},\bm{y})=D\exp(\bm{A}||\bm{\omega}||^{2}+sB\bm{\omega}^{T}\bm{y}+\bm{C}||\bm{y}||^{2})\\ \\ Re(1-4A)>0,\bm{B}=\sqrt{s(1-4\bm{A})}\\ \bm{C}=-(s+1)/2,D=(\sqrt[4]{1-4\bm{A}})^{d}\\ \bm{A}\in\mathbb{C},s\in\{-1,+1\}\end{array}$ (34)
-where $\sqrt{.}$ and .$n]{.}$ denoting a principal root with a complex argument
-OPRF (Optimized positive RF) (Likhosherstov et al., 2022) is the solution to the minimization of the variance of GERF.
-Specifically it is defined as Equation 34 with $s=+1$, $||\bm{x}+\bm{y}||^{2}>0$, and $\bm{A}\in\mathbb{R}$ defined in terms of $p^{*}$ as:
-$\begin{array}[]{cc}\bm{A}=(1-1/p^{*})/8\\ p^{*}=(\sqrt{(2||\bm{x}+\bm{y}||^{2}+d)^{2}+8d||\bm{x}+\bm{y}||^{2}}-2||\bm{x}+\bm{y}||^{2}-d)\\ /(4||\bm{x}+\bm{y}||^{2})\end{array}$ (35)
-Whilst PosRF is not bounded, OPRF is. OPRF can provide $e^{60}$$\times$ variance reduction in estimating the Softmax compared to TrigRF.
-DERF (Dense-exponential random features) (Likhosherstov et al., 2023) extends GERF and replace $\bm{A},\bm{B},\bm{C}$ with dense matrices.
-DERF is CDERF when $\bm{B},C$ are in the real instead of the complex plane.
-Minimizing the variance of DERF leads to two approaches: ADERF and SDERF.
-However both these approaches rely on SVD and eigen decompositions which are not extensively supported on GPU and deep learning libraries.
-Therefore SADERF is proposed.
-SADERF (simplified ADERF) (Likhosherstov et al., 2023) is a special case of ADERF and extends GERF, requiring only basic unary operations in addition.
-$\begin{array}[]{c}f_{SADE}^{(1)}(\bm{\omega},\bm{x})=f_{GE}^{(1)}(\bm{\omega},\Psi\bm{x}),f_{SADE}^{(2)}(\bm{\omega},\bm{y})=f_{GE}^{(2)}(\bm{\omega},\Psi^{-1}\bm{y})\\ \Psi_{l,l}^{*}=(\sum_{j}(\bm{y}_{l}^{(i)})^{2}/\sum_{i}(\bm{x}_{l}^{(i)})^{2})^{1/4}\end{array}$ (36)
-DIRF (Discretely-induced random features) (Likhosherstov et al., 2022) are based on the assumption of $p(\bm{\omega})$ being a discrete distribution with $\bm{\omega}_{1},...,\bm{\omega}_{d}$ being i.i.d., $\mathbb{P}(\bm{\omega}_{l}=k)=p_{k}$, $\sum_{k=0}^{\infty}p_{k}=1$, $p_{k}>0,k\in\{0\}\cup\mathbb{N}$.
-By making use of the Taylor series, we can define DIRF as:
-$f_{DI}(\bm{\omega},\bm{x})=\exp(-\frac{||\bm{x}||^{2}}{2})\prod_{l=1}^{d}\bm{x}_{i}^{\bm{\omega}_{l}}(\bm{\omega}_{l}!)^{\frac{-1}{2}}p_{\bm{\omega}_{l}}^{\frac{-1}{2}}$ (37)
-DIRF has two variations: PoisRF (Poisson RF) which requires $p(.)$ to be the Poisson distribution and GeomRF (Geometric RF) which requires $p(.)$ to be the geometric distribution.
-Since the many weight matrices introduced in Section 3.1 do not readily approximate these distributions, we decide to leave DIRF out of the paper.
-HRF (Hybrid RF) (Choromanski et al., 2022) combines multiple base estimators, similar to multiple kernel learning methods (Bach et al., 2004), in order to provide the most accurate approximation in regions of interest.
-Choromanski et al.
-on solving softmax approximation using HRF.
-HRF is constructed using a weighted combination $\epsilon$ of $p+1$ base estimators $\hat{SM}^{k}(\bm{x},\bm{y})$ and $\lambda$-coefficients, $\Lambda$ $p$ binary estimators $\lambda^{k}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow[0,1]$, both constructed independently:
-$\hat{SM}^{\epsilon,\Lambda}=\sum_{k=1}^{p}\hat{\lambda}^{k}(\bm{x},\bm{y})\hat{SM}^{k}(\bm{x},\bm{y})+(1-\sum_{k=1}^{p}\hat{\lambda}^{k}(\bm{x},\bm{y}))\hat{SM}^{p+1}(\bm{x},\bm{y})$ (38)
-Since HRF requires the base estimators $SM(\bm{x},\bm{y})$ to be unbiased, the easiest solution is the complex exponential $\hat{SM}^{cexp}_{m}$ which can be derived directly from the expectation of the softmax as follows:
-$\begin{array}[]{cc}\hat{SM}^{cexp}_{m}(\bm{x},\bm{y})=\psi^{m}_{A}(\bm{x})^{T}\psi^{m}_{(\bm{A}^{T})^{-1}}(\bm{y})\\ \psi^{m}_{\bm{M}}(u)=\frac{1}{\sqrt{m}}\exp(-\frac{(\bm{Mu})^{2}}{2})\\ (\exp(\bm{\omega}^{T}_{i}\bm{Mu}),...,\exp(\bm{\omega}_{m}^{T}\bm{Mu}))^{T},\bm{\omega}_{i}\sim\mathcal{N}(\bm{0},\bm{I}_{d})\end{array}$ (39)
-To construct $\lambda$-coefficients, three methods are proposed: simultaneous accurate approximation of both small and large values, Gaussian Lambda Coefficients, and adaption to data admitting clustering structure.
-A.6 Experimental Details in Spectraformer
-Table 8: Hyper-parameters for Spectraformer, following the setting from Chen et al.
-(2021) due to limited computational power.
-$L$ stands for ListOps, $T$ for Text, and $R$ for Retrieval.
-The seed that can be used to reproduce our experiments is 42.
-L T R
-Embedding dim. 64 64 64
-Transformer dim 64 64 64
-Hidden dim 128 128 128
-Head dim 32 32 32
-Num. heads 2 2 2
-Num. layers 2 2 2
-Vocabulary size 32 512 512
-Sequence length 2000 4000 4000
-Dropout rate 0.1 0.1 0.1
-Att. dropout rate 0.1 0.1 0.1
+5.2. Kernel Approximation Beyond Random Features
+Kernel approximation is presented in Section 2.3 through the use of random features based on Bochner’s theorem limiting to only symmetric, positive-definite (PD) kernels.
+While Spectraformer focuses on approximating the symmetric softmax kernel, other research has extended the RFF paradigm to handle other classes of functions.
+Notably, the AsK-RFFs (Asymmetric Kernel RFFs) (He et al., 2024) method generalizes RFFs to handle asymmetric kernels.
+The key theoretical innovation is the use of a complex measure expanding on the Bochner theorem, which can be decomposed into four finite positive measures.
+This allows AsK-RFFs to create two distinct feature mappings that can approximate an asymmetric kernel.
+A different class of kernel functions are the Generalized Zonal Kernels (GZKs) which are dot-product kernels on the unit sphere, which is different from shift-invariant kernels which are the focal point of this paper.
+Nevertheless, the Gaussian is a part of this class.
+Random Gegenbauer Features (Han et al., 2022) are designed to approximate the GZKs.
+This is done by performing a low-rank approximation via sampling the GZK’s feature map, defined by an infinite series of Gegenbauer polynomials instead of using the entire feature map.
+In addition, there has been other kernel approximation techniques outside of random features.
+These include greedy basis selection (Smola and Schökopf, 2000), divide and conquer (Hsieh et al., 2014; Zhang et al., 2013), nyström methods (Williams and Seeger, 2000) (which led to the development of Skyformer (Chen et al., 2021) and Nyströmformer (Xiong et al., 2021)).
+Nystrom methods are a type of low rank attention of which Low-Rank Transformer (Winata et al., 2020) is a part of.
+There is also sparse attention (Big Bird (Zaheer et al., 2020) which combines windowed, global and random attention, Reformer (Kitaev et al., 2020) which uses locality-sensitive hashing, Longformer (Beltagy et al., 2020) which combines band attention and global attention on special tokens like [CLS], ETC (Encoding Long and Structured Inputs) (Ainslie et al., 2020), Routing Transformer (Roy et al., 2021), Star Transformer (Guo et al., 2019)), time-series inspired methods (Informer (Zhou et al., 2021)), or Linformer (Wang et al., 2020).
+The inclusion of many of these models in our LRA benchmark comparison (Table 1) provides a broader comparison for evaluating the performance of Spectraformer against other efficient Transformers beyond random features.
+For a more comprehensive overview, please see the survey in (Elouargui et al., 2023).
+An alternative to using random features is directly specifying $\phi(x)$ without approximating the softmax via random features. cosFormer (Qin et al., 2022) specifies $\phi(x)$ as LeakyReLU and ReLU.
+6. Conclusion
+We present Spectraformer, an open-source framework for approximating and learning the attention kernel in the Transformer.
+Our paper generalizes past works, and presents empirical findings on different component function and weight matrix combinations.
+We experiment with 18 combinations, of which 14 are novel.
+Spectraformer has produced variants that are competitive against a wide range of efficient Transformers, including those based on sparse attention and other approximation methods as shown in our expanded LRA benchmark.
+We have demonstrated that a novel combination discovered through our framework, OPRF-FastFoodL, performs on par with leading models like Nystromformer and BigBird, establishing a new SOTA among random feature-based methods and a strong contender in the overall efficient Transformer landscape.
+We suggest the use of weight matrices FastFoodL and QMC combined with component functions based on specific computational requirements (PosRF for less and OPRF for more computational requirement).
+Our work shows the validity of Spectraformer for studying and producing novel variants based on existing and future research in random features and distribution sampling.
+7. Future work, limitations, and broader impact
+Spectraformer provides an experimental framework for random features in efficient Transformer with kernelized attention.
+We are aware that many works cited do not provide comparable error approximation bounds.
+In addition, as (Yao et al., 2023) point out, most theoretical bounds are either highly conservative or involving unknown qualities, therefore making comparisons impossible.
+Therefore, we have validated our framework on extensive tasks, and provided comprehensive analyses and intuitive explanation.
+We welcome future work to generalize and build on Spectraformer.
+We have experimented with a significant number of combinations but not the exhaustive list in Sections 3.1 and 3.2.
+Future experiments can work with these additional techniques.
+In addition, providing a definitive theoretical reason for the success of specific method-task combinations, such as PosRF-QMC on ListOps, remains a significant open question that aligns with broader challenges in the kernel and random features literature.
+A rigorous theoretical investigation into these interactions is a valuable direction for future work.
+Transformer-based architecture can also benefit from Spectraformer.
+Past random feature Transformer works (Likhosherstov et al., 2022, 2023) have experimented with conformer and vision Transformer with ViT.
+Jeevan and Sethi (2022) provide a good foundation to extend our work.
+Spectraformer, although currently experimented in the Transformer setting, can also extend to non-parametric kernel classification.
+It could also benefit from ablation studies like the comparison variance in (Chowdhury et al., 2022) or in (Yao et al., 2023).
+Given the complex nature of language models in the ethical, energy, and societal domains, our work should be used responsibly (Weidinger et al., 2021).
+Acknowledgments
+The production of the code used in this paper was partially reliant on generative AI code assistant.
+All AI-generation has been validated by the authors.
+This research was undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government.
+We would also like to acknowledge the ARC Centre of Excellence for Automated Decision Making and Society (CE200100005).
+Appendix
+Table 6.
+Hyper-parameters for all Spectraformer experiments on the Long Range Arena (LRA) benchmark tasks.
+The settings follow Chen et al.
+(2021) to ensure a fair comparison under limited computational power.
+ListOps (L) Text (T) Retrieval (R) Image (I) Pathfinder (P)
+Embedding dim. 64
+Transformer dim 64
+Hidden dim 128
+Head dim 32
+Num. heads 2
+Num. layers 2
+Vocabulary size 32 512 512 256 512
+Sequence length 2000 4000 8000 1024 1024
+Dropout rate 0.1
+Att. dropout rate 0.1
 Pooling mode mean
-Batch size 32 32 16
-Learning rate 0.0001 0.0001 0.0002
-Warmup steps 1000 80 800
+Batch size 32 32 16 256 128
+Learning rate 0.0001 0.0001 0.0002 0.0001 0.0002
+Warmup steps 1000 80 800 175 312
 Learning rate decay linear
-Weight decay 0 0 0
-Evaluation freq. 500 500 1000
-Num. epochs 50k 50K 50k
-Num. init steps 1k 3K 3k
-Num. eval steps 62 200 300
-Patience 10 10 10
-Num. features 128 128 128
-Table 9: Mean statistics across component functions for each task.
-Task is denoted on the top left cell of each table.
-Mean statistics include ‘acc’ (accuracy), ‘time’ (training time), and ‘mem’ (peak memory consumption).
-ListOps acc time mem
-PosRF 31.98 0.63 1.10
-OPRF 36.11 0.74 1.28
-SADERF 36.33 0.75 1.35
-Text acc time mem
-OPRF 61.82 1.40 2.54
-PosRF 62.31 1.21 2.19
-SADERF 62.33 1.39 2.69
-Retrieval acc time mem
-PosRF 77.32 1.22 2.01
-SADERF 78.85 1.43 2.54
-OPRF 79.11 1.39 2.41
+Weight decay 0
+Evaluation freq. 500 500 1000 50 500
+Num. epochs 50k
+Num. init steps 1k 3K 3k 0 3.5k
+Num. eval steps 62 200 300 20 156
+Patience 10
+Num. features {64, 128, 256, 512}
+Table 7.
+Detailed experimental results for all variants on the LRA benchmark over five seeds.
+We report mean accuracy, training time, and peak memory for feature dimensions ($d$) of 64, 128, 256, and 512.
+L: ListOps, T: Text, R: Retrieval, I: Image, P: Pathfinder.
+Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
+Model Variant $d$ L T R I P $\mu$ L T R I P $\mu$ L T R I P $\mu$
+PosRF-FastFoodL 64 31.50 (6.32) 64.61 (0.29) 77.13 (0.99) 38.28 (0.43) 66.38 (0.71) 55.58 1.04 2.08 2.17 3.95 2.02 2.25 0.77 1.54 1.50 3.01 1.54 1.67
+128 29.54 (5.46) 64.61 (0.29) 77.10 (1.02) 38.28 (0.43) 66.38 (0.71) 55.18 1.02 2.00 2.01 3.88 1.98 2.18 0.77 1.54 1.50 3.01 1.54 1.67
+256 29.54 (5.46) 64.61 (0.29) 76.94 (0.77) 38.28 (0.43) 66.17 (0.29) 55.11 1.01 1.98 2.02 3.88 1.98 2.17 0.77 1.54 1.50 3.01 1.54 1.67
+512 31.50 (6.32) 64.61 (0.29) 77.13 (0.99) 38.28 (0.43) 66.38 (0.71) 55.58 1.02 1.99 2.02 3.89 1.98 2.18 0.77 1.54 1.50 3.01 1.54 1.67
+PosRF-ORF 64 37.51 (0.37) 60.46 (0.34) 80.68 (0.47) 34.41 (1.11) 66.93 (0.59) 56.00 0.48 0.80 0.81 1.52 0.80 0.88 0.76 1.51 1.42 2.99 1.50 1.64
+128 34.50 (6.49) 61.10 (1.76) 80.53 (0.33) 33.72 (1.03) 65.52 (4.89) 55.07 0.56 1.05 1.06 2.02 1.06 1.15 1.14 2.26 2.05 4.49 2.26 2.44
+256 37.31 (0.51) 62.19 (1.82) 80.73 (0.59) 32.82 (0.42) 64.22 (4.62) 55.46 0.78 1.54 1.55 3.06 1.55 1.70 1.89 3.76 3.31 7.50 3.76 4.04
+512 37.13 (0.15) 61.34 (1.23) 80.30 (0.27) 33.79 (1.11) 67.35 (0.31) 55.98 1.22 2.55 2.49 5.07 2.53 2.77 3.40 6.77 5.82 13.52 6.77 7.26
+PosRF-SORF 64 25.26 (8.01) 63.50 (0.63) 67.91 (0.39) 28.45 (0.81) 49.88 (0.13) 47.00 0.48 0.80 0.81 1.52 0.80 0.89 0.76 1.51 1.42 2.99 1.50 1.64
+128 21.91 (5.98) 62.06 (1.36) 66.33 (0.41) 28.70 (3.87) 52.37 (5.24) 46.27 0.55 1.06 1.06 2.00 1.05 1.14 1.14 2.26 2.05 4.49 2.26 2.44
+256 20.63 (3.02) 61.10 (1.11) 65.81 (0.78) 26.52 (2.40) 50.12 (0.36) 44.84 0.78 1.54 1.56 3.05 1.55 1.69 1.89 3.76 3.30 7.50 3.76 4.04
+512 20.21 (2.70) 61.48 (0.40) 65.12 (1.11) 25.77 (0.64) 50.05 (0.28) 44.53 1.22 2.56 2.50 5.07 2.54 2.78 3.40 6.77 5.82 13.52 6.77 7.26
+PosRF-QMC 64 37.02 (0.17) 61.25 (1.67) 80.71 (0.58) 34.42 (1.36) 67.17 (0.35) 56.11 0.48 0.80 0.81 1.52 0.81 0.89 0.76 1.51 1.42 2.99 1.50 1.64
+128 37.05 (0.16) 60.93 (0.59) 80.67 (0.29) 34.03 (0.94) 67.25 (0.44) 55.99 0.55 1.05 1.05 1.99 1.05 1.14 1.14 2.26 2.05 4.49 2.26 2.44
+256 37.22 (0.13) 61.53 (1.43) 80.67 (0.32) 34.47 (0.71) 62.93 (5.13) 55.36 0.78 1.54 1.56 3.04 1.55 1.69 1.89 3.76 3.30 7.50 3.76 4.04
+512 37.06 (0.63) 61.92 (1.91) 80.66 (0.24) 34.03 (0.55) 66.85 (0.70) 56.10 1.22 2.56 2.50 5.07 2.53 2.78 3.40 6.77 5.82 13.52 6.77 7.26
+PosRF-$M$ 64 37.06 (0.10) 60.35 (1.02) 80.83 (0.18) 33.55 (1.07) 64.76 (4.63) 55.31 0.48 0.80 0.81 1.51 0.80 0.88 0.76 1.51 1.42 2.99 1.50 1.64
+128 37.13 (0.27) 62.88 (1.17) 80.58 (0.44) 33.82 (0.51) 67.11 (0.45) 56.30 0.56 1.05 1.06 2.02 1.05 1.15 1.14 2.26 2.05 4.49 2.26 2.44
+256 37.14 (0.16) 62.66 (1.48) 80.10 (0.42) 33.71 (1.41) 64.43 (5.05) 55.61 0.77 1.54 1.54 3.03 1.55 1.69 1.89 3.76 3.30 7.50 3.76 4.04
+512 34.31 (6.23) 62.78 (1.49) 80.32 (0.29) 33.14 (1.20) 64.97 (4.38) 55.10 1.22 2.55 2.50 5.09 2.54 2.78 3.40 6.77 5.82 13.52 6.77 7.26
+PosRF-SGQ 64 34.40 (6.21) 63.04 (0.70) 78.11 (0.46) 37.26 (0.63) 57.72 (8.55) 54.10 0.48 0.80 0.81 1.52 0.80 0.88 0.76 1.51 1.42 2.99 1.50 1.64
+128 28.81 (7.77) 62.30 (0.47) 78.31 (0.39) 37.96 (0.76) 59.71 (7.87) 53.42 0.56 1.06 1.05 2.01 1.06 1.15 1.14 2.26 2.05 4.49 2.26 2.44
+256 31.58 (7.71) 62.26 (0.36) 78.28 (0.27) 37.90 (0.58) 57.46 (5.25) 53.50 0.78 1.54 1.55 3.03 1.56 1.69 1.89 3.76 3.30 7.50 3.76 4.04
+512 31.52 (7.60) 62.57 (0.64) 78.02 (0.33) 37.42 (0.51) 53.87 (3.06) 52.68 1.22 2.55 2.49 5.07 2.54 2.78 3.40 6.77 5.82 13.52 6.77 7.26
+OPRF-FastFoodL 64 37.92 (0.39) 64.32 (0.61) 77.86 (1.37) 38.18 (0.71) 66.76 (0.39) 57.01 1.06 2.06 2.11 4.02 2.05 2.26 0.84 1.68 1.64 3.26 1.68 1.82
+128 37.73 (0.28) 64.32 (0.61) 78.65 (1.94) 38.41 (0.66) 66.76 (0.39) 57.17 1.07 2.07 2.11 4.02 2.05 2.26 0.84 1.68 1.64 3.26 1.68 1.82
+256 37.90 (0.36) 64.32 (0.61) 77.32 (0.22) 38.41 (0.66) 66.60 (0.15) 56.91 1.06 2.06 2.11 4.01 2.05 2.26 0.84 1.68 1.64 3.26 1.67 1.82
+512 37.92 (0.39) 64.32 (0.61) 77.86 (1.37) 38.41 (0.66) 66.76 (0.39) 57.05 1.06 2.06 2.12 4.05 2.07 2.27 0.84 1.68 1.64 3.26 1.68 1.82
+OPRF-ORF 64 38.03 (0.41) 60.29 (0.58) 80.81 (0.35) 34.43 (0.49) 66.82 (0.74) 56.08 0.55 0.92 0.93 1.76 0.93 1.02 0.86 1.70 1.65 3.37 1.70 1.86
+128 38.05 (0.78) 60.18 (0.49) 81.19 (0.23) 33.70 (0.57) 67.24 (0.88) 56.07 0.68 1.26 1.24 2.47 1.29 1.39 1.33 2.64 2.50 5.25 2.64 2.87
+256 37.67 (0.26) 60.95 (1.38) 80.76 (0.50) 33.29 (0.50) 64.66 (4.82) 55.47 0.98 1.92 1.86 4.21 2.11 2.22 2.27 4.52 4.19 9.01 4.52 4.90
+512 37.68 (0.27) 60.07 (0.49) 80.61 (0.20) 34.01 (1.34) 66.92 (0.26) 55.86 1.63 3.34 3.11 7.72 3.55 3.87 4.15 8.28 7.58 16.53 8.28 8.96
+OPRF-SORF 64 27.60 (3.10) 65.22 (0.18) 70.73 (1.65) 35.49 (2.67) 50.41 (0.50) 49.89 0.55 0.93 0.94 1.77 0.93 1.02 0.86 1.70 1.65 3.37 1.70 1.86
+128 29.59 (3.77) 64.81 (0.22) 77.12 (0.78) 37.42 (0.47) 63.98 (5.38) 54.58 0.68 1.26 1.24 2.42 1.28 1.38 1.33 2.64 2.50 5.25 2.64 2.87
+256 28.33 (5.15) 64.72 (0.31) 77.59 (0.95) 37.34 (1.16) 63.13 (3.80) 54.22 0.98 1.92 1.86 4.23 2.11 2.22 2.27 4.52 4.19 9.01 4.52 4.90
+512 25.64 (1.89) 64.82 (0.27) 76.97 (0.68) 37.89 (0.70) 65.99 (0.50) 54.26 1.63 3.34 3.10 7.73 3.54 3.87 4.15 8.28 7.58 16.53 8.28 8.96
+OPRF-QMC 64 38.07 (0.69) 60.52 (0.62) 80.45 (0.69) 34.33 (1.02) 67.04 (0.18) 56.08 0.55 0.92 0.94 1.77 0.94 1.03 0.86 1.70 1.65 3.37 1.70 1.86
+128 37.86 (0.40) 60.87 (1.85) 80.44 (0.20) 33.87 (0.69) 66.81 (0.68) 55.97 0.68 1.26 1.24 2.43 1.28 1.38 1.33 2.64 2.50 5.25 2.64 2.87
+256 37.55 (0.06) 60.16 (0.65) 80.45 (0.22) 34.35 (0.40) 62.75 (5.33) 55.05 0.98 1.93 1.87 4.24 2.12 2.23 2.27 4.52 4.19 9.01 4.52 4.90
+512 37.83 (0.58) 59.84 (0.87) 80.24 (0.83) 33.80 (0.56) 66.57 (0.90) 55.66 1.64 3.34 3.11 7.74 3.53 3.87 4.15 8.28 7.58 16.53 8.28 8.96
+OPRF-$M$ 64 37.90 (0.66) 60.29 (0.61) 80.97 (0.31) 33.61 (0.62) 66.08 (2.11) 55.77 0.55 0.92 0.93 1.76 0.93 1.02 0.86 1.70 1.65 3.37 1.70 1.86
+128 38.31 (0.38) 60.01 (1.01) 81.03 (0.31) 33.93 (0.85) 65.17 (5.17) 55.69 0.68 1.26 1.29 2.47 1.28 1.40 1.33 2.64 2.50 5.25 2.64 2.87
+256 37.59 (0.47) 60.27 (0.67) 80.94 (0.46) 33.47 (0.87) 63.15 (5.61) 55.09 0.98 1.92 1.87 4.23 2.10 2.22 2.27 4.52 4.19 9.01 4.52 4.90
+512 37.73 (0.25) 60.39 (2.04) 80.82 (0.28) 33.49 (0.76) 66.43 (1.27) 55.77 1.63 3.33 3.09 7.71 3.54 3.86 4.15 8.28 7.58 16.53 8.28 8.96
+OPRF-SGQ 64 37.66 (0.52) 61.25 (0.50) 78.68 (0.83) 36.64 (1.33) 59.77 (6.20) 54.80 0.55 0.92 0.93 1.76 0.93 1.02 0.86 1.70 1.65 3.37 1.70 1.86
+128 37.08 (0.56) 61.09 (0.77) 79.39 (0.94) 36.68 (0.28) 67.53 (2.79) 56.35 0.68 1.27 1.25 2.46 1.29 1.39 1.33 2.64 2.50 5.25 2.64 2.87
+256 37.16 (0.28) 61.36 (0.36) 79.24 (0.11) 36.34 (1.11) 65.11 (5.13) 55.84 0.98 1.92 1.87 4.22 2.11 2.22 2.27 4.52 4.19 9.01 4.52 4.90
+512 37.20 (0.51) 61.37 (0.25) 79.27 (0.37) 36.91 (1.24) 64.75 (4.87) 55.90 1.63 3.34 3.10 7.72 3.53 3.86 4.15 8.28 7.58 16.53 8.28 8.96
+SADERF-FastFoodL 64 29.72 (7.27) 64.71 (0.45) 75.92 (1.17) 38.38 (0.82) 65.97 (0.88) 54.94 1.09 2.12 2.16 4.05 2.05 2.30 0.90 1.80 1.75 3.52 1.80 1.95
+128 29.72 (7.27) 64.71 (0.45) 77.50 (0.96) 38.38 (0.82) 66.47 (1.47) 55.35 1.09 2.09 2.18 4.05 2.06 2.29 0.90 1.80 1.76 3.52 1.80 1.96
+256 29.72 (7.27) 64.71 (0.45) 76.25 (0.73) 38.05 (1.18) 65.97 (0.88) 54.94 1.07 2.07 2.16 4.04 2.05 2.28 0.90 1.80 1.75 3.52 1.80 1.95
+512 29.72 (7.27) 64.71 (0.45) 76.64 (1.59) 38.38 (0.82) 65.97 (0.88) 55.08 1.06 2.07 2.16 4.06 2.06 2.29 0.90 1.80 1.76 3.52 1.80 1.95
+SADERF-ORF 64 37.19 (0.61) 60.33 (2.13) 80.83 (0.25) 33.68 (1.23) 66.96 (0.81) 55.80 0.56 0.92 0.98 1.76 0.92 1.03 0.93 1.85 1.78 3.68 1.85 2.02
+128 37.41 (0.45) 59.92 (0.90) 81.05 (0.18) 33.06 (1.02) 67.11 (0.56) 55.71 0.69 1.25 1.28 2.50 1.28 1.40 1.40 2.79 2.63 5.56 2.79 3.04
+256 33.98 (6.41) 60.05 (0.78) 80.96 (0.17) 33.27 (0.89) 62.91 (5.16) 54.23 0.96 1.92 1.93 4.26 2.10 2.23 2.34 4.67 4.32 9.32 4.67 5.07
+512 34.45 (6.19) 61.68 (2.08) 80.62 (0.16) 33.31 (1.60) 64.53 (4.48) 54.92 1.59 3.34 3.13 7.75 3.53 3.87 4.23 8.43 7.71 16.84 8.43 9.13
+SADERF-SORF 64 26.42 (4.53) 65.22 (0.20) 68.74 (1.05) 33.39 (1.21) 50.31 (0.49) 48.82 0.56 0.92 0.98 1.76 0.92 1.03 0.93 1.85 1.78 3.68 1.85 2.02
+128 34.31 (2.50) 64.82 (0.26) 75.24 (1.12) 36.76 (1.22) 60.98 (4.90) 54.42 0.68 1.25 1.28 2.47 1.26 1.39 1.40 2.79 2.63 5.56 2.79 3.04
+256 32.36 (6.74) 64.92 (0.18) 76.65 (1.04) 36.73 (0.78) 58.76 (3.42) 53.88 0.96 1.91 1.91 4.25 2.10 2.23 2.34 4.67 4.32 9.32 4.67 5.07
+512 31.60 (2.32) 64.85 (0.16) 76.99 (0.49) 37.23 (1.08) 61.54 (4.79) 54.44 1.60 3.35 3.14 7.74 3.52 3.87 4.23 8.43 7.71 16.84 8.43 9.13
+SADERF-QMC 64 37.35 (0.40) 60.76 (1.74) 80.93 (0.20) 34.55 (1.23) 66.60 (1.77) 56.04 0.56 0.91 0.99 1.78 0.93 1.04 0.93 1.85 1.78 3.68 1.85 2.02
+128 37.18 (0.47) 60.61 (2.06) 80.67 (0.44) 34.55 (0.63) 67.32 (0.40) 56.07 0.68 1.24 1.28 2.47 1.26 1.39 1.40 2.79 2.63 5.56 2.79 3.04
+256 37.42 (0.45) 60.01 (0.63) 80.94 (0.31) 34.50 (0.64) 64.89 (4.75) 55.55 0.96 1.92 1.91 4.25 2.09 2.23 2.34 4.67 4.32 9.32 4.67 5.07
+512 36.95 (0.46) 59.56 (1.23) 80.36 (0.76) 34.21 (1.28) 61.78 (5.53) 54.57 1.60 3.34 3.13 7.75 3.56 3.88 4.23 8.43 7.71 16.84 8.43 9.13
+SADERF-$M$ 64 34.65 (5.94) 60.48 (3.13) 80.76 (0.29) 32.92 (0.95) 64.85 (4.74) 54.73 0.56 0.92 0.98 1.76 0.92 1.03 0.93 1.85 1.78 3.68 1.85 2.02
+128 37.17 (0.27) 60.81 (1.82) 81.05 (0.20) 34.03 (1.04) 65.31 (4.85) 55.67 0.69 1.25 1.28 2.49 1.28 1.40 1.40 2.79 2.63 5.56 2.79 3.04
+256 37.15 (0.64) 60.69 (1.64) 80.83 (0.19) 33.06 (0.82) 62.91 (5.97) 54.93 0.95 1.91 1.91 4.24 2.08 2.22 2.34 4.67 4.32 9.32 4.67 5.07
+512 37.69 (0.74) 61.38 (2.25) 80.81 (0.26) 33.63 (0.84) 64.97 (4.43) 55.70 1.59 3.33 3.13 7.73 3.53 3.86 4.23 8.43 7.71 16.84 8.43 9.13
+SADERF-SGQ 64 34.19 (6.27) 62.69 (0.75) 77.96 (0.68) 37.81 (0.43) 59.31 (6.87) 54.39 0.56 0.92 0.98 1.77 0.92 1.03 0.93 1.85 1.78 3.68 1.85 2.02
+128 37.22 (0.28) 62.86 (0.96) 78.51 (0.68) 37.82 (0.59) 61.27 (5.38) 55.54 0.68 1.25 1.28 2.50 1.27 1.39 1.40 2.79 2.63 5.56 2.79 3.04
+256 37.18 (0.31) 63.66 (0.22) 77.94 (0.44) 37.62 (0.55) 62.61 (3.12) 55.80 0.97 1.94 1.93 4.28 2.11 2.24 2.34 4.67 4.32 9.32 4.67 5.07
+512 37.08 (0.10) 62.92 (0.79) 78.40 (0.80) 37.89 (0.30) 65.79 (0.84) 56.42 1.60 3.35 3.14 7.75 3.53 3.88 4.23 8.43 7.71 16.84 8.43 9.13
+Table 8.
+Summary statistics across component functions ($f$) and weight matrices ($W$) for 64, 128, 256, and 512 feature settings.
+The original four tables have been consolidated for clarity and compactness.
+We report mean accuracy (test set %), mean training time (hours), and mean peak memory consumption (GB).
+Tasks are abbreviated as $L$: ListOps, $T$: Text, $R$: Retrieval, II: Image, $P$: Pathfinder, and $\mu$ for the average.
+Type Method Features Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
+L T R I P $\mu$ L T R I P $\mu$ L T R I P $\mu$
+$f$ PosRF 64 33.79 62.20 77.56 34.39 62.14 54.02 0.57 1.02 1.04 1.93 1.01 1.11 0.76 1.51 1.43 2.99 1.51 1.64
+128 31.49 62.32 77.25 34.42 63.06 53.71 0.63 1.21 1.21 2.32 1.21 1.32 1.08 2.14 1.96 4.25 2.14 2.31
+256 32.24 62.39 77.09 33.95 60.89 53.31 0.82 1.61 1.63 3.18 1.62 1.77 1.70 3.39 3.00 6.75 3.39 3.65
+512 31.95 62.45 76.89 33.74 61.58 53.16 1.19 2.46 2.41 4.88 2.44 2.68 2.96 5.90 5.07 11.77 5.90 6.33
+OPRF 64 36.20 61.98 78.25 35.45 62.81 54.94 0.63 1.11 1.13 2.14 1.12 1.23 0.86 1.70 1.65 3.35 1.69 1.85
+128 36.44 61.88 79.64 35.67 66.25 55.97 0.74 1.40 1.40 2.71 1.41 1.53 1.25 2.48 2.36 4.92 2.48 2.70
+256 36.03 61.97 79.38 35.53 64.23 55.43 0.99 1.94 1.91 4.19 2.10 2.23 2.03 4.05 3.77 8.05 4.04 4.39
+512 35.67 61.80 79.30 35.75 66.24 55.75 1.54 3.12 2.94 7.11 3.29 3.60 3.60 7.18 6.59 14.32 7.18 7.77
+SADERF 64 33.25 62.36 77.53 35.12 62.34 54.12 0.65 1.12 1.18 2.15 1.11 1.24 0.93 1.84 1.78 3.66 1.84 2.01
+128 35.50 62.29 79.01 35.77 64.74 55.46 0.75 1.39 1.43 2.75 1.40 1.54 1.32 2.63 2.48 5.22 2.63 2.86
+256 34.63 62.34 78.93 35.54 63.01 54.89 0.98 1.94 1.96 4.22 2.09 2.24 2.10 4.19 3.89 8.36 4.19 4.55
+512 34.58 62.52 78.97 35.86 64.10 55.33 1.51 3.13 2.97 7.11 3.29 3.58 3.67 7.32 6.72 14.54 7.33 7.87
+$W$ FastFoodL 64 33.05 64.55 76.97 38.28 66.37 55.84 1.06 2.09 2.15 4.01 2.04 2.27 0.84 1.67 1.63 3.26 1.67 1.82
+128 32.33 64.55 77.75 38.35 66.54 55.90 1.06 2.05 2.10 3.99 2.03 2.25 0.84 1.67 1.63 3.26 1.67 1.82
+256 32.38 64.55 76.83 38.25 66.25 55.65 1.05 2.03 2.10 3.98 2.03 2.24 0.84 1.67 1.63 3.26 1.67 1.82
+512 33.05 64.55 77.21 38.35 66.37 55.91 1.05 2.04 2.10 4.00 2.04 2.25 0.84 1.67 1.63 3.26 1.67 1.82
+$M$ 64 36.54 60.37 80.85 33.36 65.23 55.27 0.53 0.88 0.91 1.68 0.88 0.97 0.85 1.69 1.62 3.35 1.68 1.84
+128 37.53 61.23 80.89 33.93 65.86 55.89 0.64 1.19 1.21 2.33 1.21 1.32 1.29 2.56 2.39 5.10 2.56 2.78
+256 37.29 61.21 80.62 33.41 63.50 55.21 0.90 1.79 1.77 3.83 1.91 2.04 2.17 4.32 3.94 8.61 4.32 4.67
+512 36.58 61.52 80.65 33.42 65.46 55.52 1.48 3.07 2.91 6.85 3.20 3.50 3.92 7.83 7.03 15.63 7.83 8.45
+ORF 64 37.58 60.36 80.78 34.17 66.90 55.96 0.53 0.88 0.91 1.68 0.89 0.98 0.85 1.69 1.62 3.35 1.68 1.84
+128 36.65 60.40 80.92 33.49 66.62 55.62 0.64 1.19 1.19 2.33 1.21 1.31 1.29 2.56 2.39 5.10 2.56 2.78
+256 36.32 61.06 80.82 33.13 63.93 55.05 0.91 1.80 1.78 3.84 1.92 2.05 2.17 4.32 3.94 8.61 4.32 4.67
+512 36.42 61.03 80.51 33.73 66.27 55.89 1.48 3.08 2.91 6.78 3.20 3.45 3.92 7.83 7.03 15.54 7.83 8.34
+QMC 64 37.48 60.84 80.70 34.43 66.93 56.08 0.53 0.88 0.91 1.69 0.89 0.98 0.85 1.69 1.62 3.35 1.68 1.84
+128 37.36 60.80 80.59 34.15 67.13 56.01 0.64 1.18 1.19 2.30 1.20 1.30 1.29 2.56 2.39 5.10 2.56 2.78
+256 37.40 60.57 80.68 34.44 63.52 55.32 0.91 1.79 1.78 3.84 1.92 2.05 2.17 4.32 3.94 8.61 4.32 4.67
+512 37.28 60.44 80.42 34.01 65.07 55.44 1.49 3.08 2.91 6.86 3.21 3.51 3.92 7.83 7.03 15.63 7.83 8.45
+SGQ 64 35.42 62.33 78.25 37.23 58.93 54.43 0.53 0.88 0.91 1.68 0.88 0.98 0.85 1.69 1.62 3.35 1.68 1.84
+128 34.37 62.08 78.74 37.49 62.84 55.10 0.64 1.19 1.19 2.32 1.21 1.31 1.29 2.56 2.39 5.10 2.56 2.78
+256 35.31 62.43 78.49 37.29 61.72 55.05 0.91 1.80 1.79 3.84 1.93 2.05 2.17 4.32 3.94 8.61 4.32 4.67
+512 35.27 62.29 78.60 37.41 61.47 54.69 1.48 3.08 2.94 6.85 3.20 3.52 3.92 7.83 7.12 15.63 7.83 8.48
+SORF 64 26.43 64.65 69.13 32.44 50.20 48.57 0.53 0.88 0.91 1.68 0.89 0.98 0.85 1.69 1.62 3.35 1.68 1.84
+128 28.60 63.90 72.90 34.29 59.11 51.76 0.64 1.19 1.19 2.30 1.20 1.30 1.29 2.56 2.39 5.10 2.56 2.78
+256 27.10 63.58 73.35 33.53 57.34 50.98 0.91 1.79 1.78 3.84 1.92 2.05 2.17 4.32 3.94 8.61 4.32 4.67
+512 25.82 63.71 73.03 33.63 59.19 51.08 1.49 3.08 2.91 6.85 3.20 3.51 3.92 7.83 7.03 15.63 7.83 8.45
+Figure 7.
+Efficiency of Training Time and Memory Usage w.r.t.
+Sequence Length for Different Number of Features.
+Only Text, Retrieval and Pathfinder data are used due to having the same number of labels (2) compared to ListOps and Image (10).